### v2

# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

Abstract
When prior information is lacking, the go-to strategy for probabilistic inference is to combine a “default prior” and the likelihood via Bayes’s theorem. Objective Bayes, (generalized) fiducial inference, etc. fall under this umbrella. This construction is natural, but the corresponding posterior distributions generally only offer limited, approximately valid uncertainty quantification. The present paper takes a reimagined approach offering posterior distributions with stronger reliability properties. The proposed construction starts with an inferential model (IM), one that takes the mathematical form of a data-driven possibility measure and features exactly valid uncertainty quantification, and then returns a so-called inner probabilistic approximation thereof. This inner probabilistic approximation inherits many of the original IM’s desirable properties, including credible sets with exact coverage and asymptotic efficiency. The approximation also agrees with the familiar Bayes/fiducial solution obtained in applications where the model has a group transformation structure. A Monte Carlo method for evaluating the probabilistic approximation is presented, along with numerical illustrations.
Keywords and phrases: Confidence distribution; credal set; fiducial inference; p-value; possibility theory; relative likelihood; validity.
1 Introduction
The program in which science advances by formulating what’s currently known, however vague or incomplete that might be, as prior beliefs and then applying a normative procedure to update those prior beliefs in light of new data is natural and deeply rooted in all of us who’ve studied probability and statistics. Despite its appeal, this program is challenging to implement for several reasons, not the least of which is that genuine prior information is rarely available, hence no particular formulation of prior beliefs is warranted. Brad Efron, in his presentation at the 2016 Joint Statistical Meetings in Chicago, said that “scientists like to work on new problems,” meaning that there’s often no history or experience to draw from in order to build a meaningful prior distribution. For similar reasons, Efron, (2013) stated more formally that “…the most important unresolved problem in statistical inference is the use of Bayes theorem in the absence of prior information.” The present paper offers new insights, theory, and methodology with the ambitious goal of resolving this most important unresolved problem.
Fisher, (1933); Fisher, 1935a ; Fisher, 1935b and his fiducial argument was the first proposed solution to the above problem, making “a bold attempt to make the Bayesian omelet without breaking the Bayesian egg” (Savage, 1961); see Zabell, (1992) for an excellent survey. The consensus is that Fisher’s solution failed, but even his “biggest blunder” (Efron, 1998) had a major impact, inspiring fundamental advances such as confidence limits (Neyman, 1941), imprecise probability (Dempster, 1966), and new proposed solutions, including generalized fiducial inference (Hannig et al., 2016; Murph et al., 2021), confidence distributions (Xie and Singh, 2013; Thornton and Xie, 2020; Cox, 2006; Schweder and Hjort, 2016), and objective/non-informative Bayes (Jeffreys, 1946; Berger, 2006; Berger et al., 2009; Kass and Wasserman, 1996). In what follows, to keep terminology simple, I’ll (inaccurately) refer to all of these approaches as no-prior Bayes solutions.
A difficulty with Bayesian-like probabilistic inference in this no-prior setting concerns the interpretation of the posterior probabilities themselves. When genuine prior information is available, then the Bayesian posterior probabilities are the unique coherent updates of prior beliefs given the observed data. But when there’s no prior beliefs to update, and a default prior is used in its place, of course the aforementioned “updating” interpretation disappears and, consequently, it’s unclear if the corresponding posterior distribution is meaningful at all. Indeed, “[Bayes’s formula] does not create real probabilities from hypothetical probabilities” (Fraser, 2014) and, put a bit more forcefully,
…any serious mathematician would surely ask how you could use [Bayes’s theorem] with one premise missing by making up an ingredient and thinking that the conclusions of the [theorem] were still available. (Fraser, 2011b )
Fortunately, coherent updating of prior beliefs isn’t necessary for a posterior to have belief-forming inferential weight. But making this case requires, first, defining what it takes for a framework to have belief-forming inferential weight and, second, demonstrating that a proposed framework meets this requirement. To me, the only way forward is to demonstrate that the posterior probabilities are reliable, i.e., it’s a demonstrably rare event that the posterior assigns high (resp. low) probabilities to assertions about the unknown that are false (resp. true). The advantage of such a requirement is that the belief-forming inferential weight is inherited from Fisher’s inductive logic: if, e.g., assigning high probability to a false assertion is a rare event and, in the application at hand, the assertion is assigned high probability, then it’s safe to infer that the assertion is true because rare events effectively don’t happen. No-prior Bayes solutions often have some form of reliability, but the false confidence theorem (Balch et al., 2019) says that none have the strong reliability property I just described; see, also, Martin, (2019); Martin, 2024b and Section 2 below. Since no-prior Bayes solutions can’t offer reliable belief assignments, there’s motivation to look beyond these existing approaches.
An alternative to these probabilistic approaches is the inferential model (IM) framework (e.g., Martin and Liu, 2013; Martin and Liu, 2015a ; Martin, 2022b ; Martin, 2024a ), which returns a data-dependent possibility measure offering reliable uncertainty quantification about unknown parameters. Further details and references about possibility measures and the specific IM construction are given in Section 2 below. Importantly, my proposed shift from probabilistic to possibilistic uncertainty quantification facilitates strong reliability and, among other things, this strong reliability implies the usual frequentist error rate control without giving up on the fully conditional Bayesian-like reasoning.
Despite the obvious benefits, I’m under no delusion that statisticians will abandon probabilistic inference in favor of possibilsitic IMs in the near future. But that doesn’t mean IMs have to wait for the distant future to make a contribution. The “reIMagined” approach I’m proposing here starts with the provably reliable possibilistic IM and then reads off and returns an inner probabilistic approximation as a novel no-prior Bayes solution. As I’ll demonstrate below, the proposed inner probabilistic approximation inherits some—but necessarily not all—of the original possibilistic IM’s strong reliability properties: in particular, the corresponding credible sets for the full parameter are exact confidence sets. There’s just only so much that can be achieved by constructing default priors and checking if the corresponding posterior distributions are reliable. This motivates a new approach, and my suggestion is to prioritize the reliability properties and directly construct a data-dependent probability that has those desired properties.
The remainder of the paper is organized as follows. Some brief background on possibility theory and possibilistic IMs is presented in Section 2. A characterization of the possibilistic IM’s credal set is presented in Section 3, which is then used to define the corresponding inner probabilistic approximation. Various properties of the proposed inner probabilistic approximation are developed in Section 4, including a correspondence with the familiar no-prior Bayes solutions in models that feature a group-theoretic invariance property and a version of the celebrated Bernstein–von Mises theorem that establishes its asymptotic efficiency. A strategy for (approximately) computing the IM’s inner probabilistic approximation is presented in Section 5 and a new, valid, and efficient solution to the technically challenging and practically relevant Behrens–Fisher problem is presented in Section 6. Some concluding remarks are given in Section 7 and the appendices contain some additional technical details and illustrations.
2 Background
2.1 Possibility theory
Possibility measures (e.g., Dubois and Prade, 1988) are among the simplest forms of imprecise probability, closely linked to fuzzy set theory (e.g., Zadeh, 1978) and Dempster–Shafer theory (e.g., Shafer, 1976, 1987). Some applications in statistics are described in Dubois, (2006), and below I’ll mention one possibility-theoretic approach to statistical inference that has a long history. Probability and possibility theory differ philosophically and mathematically, but here I’ll only briefly discuss the latter; for a discussion of the former, the presentation in Shackle, (1961) is clear and compelling.
The mathematical differences between probability and possibility theory can be succinctly summarized as follows: optimization is to possibility theory what integration is to probability theory. That is, a possibility measure $\overline{\Pi}$ defined on a space $\mathbb{Z}$ is determined by a function $\pi:\mathbb{Z}\to[0,1]$ with the property that $\sup_{z\in\mathbb{Z}}\pi(z)=1$. This function is called the possibility contour and the supremum-equals-1 property is a normalization condition analogous to the familiar integral-equals-1 property of probability densities. Then the possibility measure is determined by optimizing its contour, i.e., $\overline{\Pi}(A)=\sup_{z\in A}\pi(z)$, for any $A\subseteq\mathbb{Z}$, just like a probability measure is determined by integrating its density.
The different calculus has a number of implications. Of particular relevance for the developments in this paper is that the aforementioned supremum-equals-1 normalization condition ensures that $\overline{\Pi}$ is a coherent upper probability (e.g., De Cooman, 1997; De Cooman and Aeyels, 1999; Walley, 1997) in the spirit of Walley, (1991) and others. Among other things, this means that $\overline{\Pi}$ determines a non-empty (closed and convex) set of ordinary probabilities that it dominates:
$\mathscr{C}(\overline{\Pi})=\{\mathsf{Q}\in\text{probs}(\mathbb{Z}):\mathsf{Q}(H)\leq\overline{\Pi}(H)\text{ for all measurable $H$}\},$ (1)
where $\text{probs}(\mathbb{Z})$ is the set of probabilities supported on (the Borel $\sigma$-algebra of measurable subsets of) $\mathbb{Z}$. The set $\mathscr{C}(\overline{\Pi})$ is called the credal set and all (coherent) upper probabilities have one. Aside from being relatively simple, a key advantage to possibilistic uncertainty quantification is that the associated credal set has a statistically oriented characterization, which is important for the statistical developments below.
2.2 Possibilistic IMs
Consider a model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ consisting of probability distributions supported on a sample space $\mathbb{X}$, with parameter space $\mathbb{T}$. Suppose that the observable data $X$, taking values in $\mathbb{X}$, is a sample from the distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the unknown/uncertain “true value.” Prior information about $\Theta$ is assumed to be vacuous. The model and observed data $X=x$ determine a relative likelihood function
$R(x,\theta)=\frac{p_{\theta}(x)}{\sup_{\vartheta}p_{\vartheta}(x)},\quad\theta\in\mathbb{T}.$
I will assume throughout that the denominator is finite for almost all $x$.
The relative likelihood defines a possibility contour, i.e., a non-negative function such that $\sup_{\theta}R(x,\theta)=1$ for almost all $x$. This contour determines a possibility measure that can be used for data-driven uncertainty quantification about $\Theta$, which has been extensively studied in (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014). This likelihood-driven possibility has a number of desirable properties. What it lacks, however, is a justification for why the “possibilities” assigned to hypotheses about $\Theta$ have belief-forming inferential weight. Fortunately, it’s straightforward to achieve the desired goal by “validifying” (Martin, 2022a ; Martin, 2022b ) the relative likelihood; see, also, Martin, (2015, 2018). This amounts to applying a version of the probability-to-possibility transform (e.g., Dubois et al., 2004; Hose, 2022), and the result is the possibilistic IM’s contour function:
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\bigl{\{}R(X,\theta)\leq R(x,\theta)\bigr{\}},\quad\theta\in\mathbb{T}.$ (2)
The corresponding possibility measure is defined via optimization as
$\overline{\Pi}_{x}(H)=\sup_{\theta\in H}\pi_{x}(\theta),\quad H\subseteq\mathbb{T}.$ (3)
An illustration of what a typical contour $\pi_{x}$ looks like is shown in Figure 1. Some further details about this simple example are presented at the end of Section 3 below.
Figure 1: Illustration of a typical possibility contour $\theta\mapsto\pi_{x}(\theta)$ and how the possibility measure $\overline{\Pi}_{x}$ is determined from it. Here, the hypothesis $H$ is the interval $[3,5]$ and the maximum value $\overline{\Pi}_{x}(H)$ on contour is highlighted. Horizontal dashed line at $\alpha=0.1$ determines the 90% confidence interval $C_{0.1}(x)$ in (5).
As in Section 1, an essential feature of this IM construction is its validity property:
$\sup_{\theta\in\mathbb{T}}\mathsf{P}_{\theta}\bigl{\{}\pi_{X}(\theta)\leq\alpha\bigr{\}}\leq\alpha,\quad\text{$\alpha\in[0,1]$}.$ (4)
This has a number of important consequences. First, (4) immediately implies that the upper $\alpha$ level set of the possibility contour,
$C_{\alpha}(x)=\{\theta\in\mathbb{T}:\pi_{x}(\theta)\geq\alpha\},\quad\alpha\in[0,1],$ (5)
is a $100(1-\alpha)$confidence set, i.e., $\sup_{\theta\in\mathbb{T}}\mathsf{P}_{\theta}\bigl{\{}C_{\alpha}(X)\not\ni\theta\bigr{\}}\leq\alpha$. Figure 1 gives a depiction of $C_{\alpha}(x)$, with $\alpha=0.1$. Second, from (3) and (4),
$\sup_{\theta\in H}\mathsf{P}_{\theta}\bigl{\{}\overline{\Pi}_{X}(H)\leq\alpha\bigr{\}}\leq\alpha,\quad\text{$\alpha\in[0,1]$, $H\subseteq\mathbb{T}$}.$ (6)
In words, a valid IM assigns possibility $\leq\alpha$ to true hypotheses at rate $\leq\alpha$ as a function of data $Z$. This gives the IM its “inferential weight”—(6) implies that $\overline{\Pi}_{x}(H)$ is not expected to be small when $H$ is true, so one is inclined to doubt the truthfulness of a hypothesis $H$ if $\overline{\Pi}_{x}(H)$ is small, and the strength of this inclination is determined by the magnitude of $\overline{\Pi}_{x}(H)$. This of course can be turned into a formal test procedure with frequentist error rate guarantees: by (6), the test “reject $H$ if $\overline{\Pi}_{x}(H)\leq\alpha$” controls the Type I error probability at level $\alpha$. Third, the above property ensures that the possibilistic IM is safe from false confidence (Balch et al., 2019; Martin, 2019; Martin, 2024b ), unlike default-prior Bayes and fiducial solutions. Further insights about IMs are presented in Appendix A.
The IM output is a coherent imprecise probability and, consequently, there is an associated (non-empty) credal set, $\mathscr{C}(\overline{\Pi}_{x})$ as in (1). The members of $\mathscr{C}(\overline{\Pi}_{x})$, which I’ll denote by $\mathsf{Q}_{x}$, depend on $x$ because $\overline{\Pi}_{x}$ does, but need not correspond to Bayesian posterior distributions under any prior. Fortunately, an interpretation can be given to the members of $\mathscr{C}(\overline{\Pi}_{x})$, and this is key to my developments in Section 3.
3 Inner probabilistic approximations
A well-known characterization (e.g., Couso et al., 2001; Destercke and Dubois, 2014) of the credal set for general possibility measures applied to the present case says
$\mathsf{Q}_{x}\in\mathscr{C}(\overline{\Pi}_{x})\iff\mathsf{Q}_{x}\{C_{\alpha}(x)\}\geq 1-\alpha\;\;\text{for all $\alpha\in[0,1]$},$ (7)
where $C_{\alpha}(x)$ is as defined in (5) with $\pi_{x}$ the contour corresponding to $\overline{\Pi}_{x}$. Validity of the IM implies that $C_{\alpha}(x)$ is a $100(1-\alpha)$confidence set, so, since the elements $\mathsf{Q}_{x}$ in the credal set assign at least probability $1-\alpha$ to the $100(1-\alpha)$confidence seet $C_{\alpha}(x)$, there’s good reason to call these elements confidence distributions. This definition of confidence distributions agrees with that given in Taraldsen, (2021) and generalizes those commonly found in the statistics literature (Xie and Singh, 2013; Schweder and Hjort, 2013; Cunen et al., 2020) that—for reasons discussed in Section 4.3—focus mainly on scalar parameters and the corresponding univariate confidence distributions. This connection between confidence distributions and the credal set associated with the possibilistic IM was first made in Martin, 2023a , and I’ll revisit the specific results in Section 4.1.
Refinements to the characterization in (7) can be given. For example, Wasserman, 1990b (, Theorem 2.1) characterizes the credal set associated with a belief function as a collection of suitable mixture distributions. The result below gives a similar mixture characterization but for general possibility measures, including for the IM output $\overline{\Pi}_{x}$.
Theorem 1 (Martin, 2025).
Let $\pi_{x}$ be the contour associated with the possibility measure $\overline{\Pi}_{x}$, and set $C_{\alpha}(x)=\{\theta:\pi_{x}(\theta)\geq\alpha\}$ as in (5). Then $\mathsf{Q}_{x}\in\mathscr{C}(\overline{\Pi}_{x})$ if and only if
$\mathsf{Q}_{x}(\cdot)=\int_{0}^{1}\mathsf{K}_{x}^{\beta}(\cdot)\,\mathsf{M}_{x}(d\beta),$ (8)
for some Markov kernel $\mathsf{K}_{x}^{\beta}$, fully supported on $C_{\beta}(x)$, i.e., $\mathsf{K}_{x}^{\beta}\{C_{\beta}(x)\}=1$ for each $\beta\in[0,1]$ and some probability measure $\mathsf{M}_{x}$ on $[0,1]$ such that a random variable with distribution $\mathsf{M}_{x}$ is stochastically no smaller than ${\sf Unif}(0,1)$.
A relevant question is if there exists a distinguished member $\mathsf{Q}_{x}^{\star}$ of $\mathscr{C}(\overline{\Pi}_{x})$ such that equality on the right-hand side of (7) is achieved for each $\alpha$, i.e.,
$\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}=1-\alpha,\quad\text{for all $\alpha\in[0,1]$}.$ (9)
There are cases in which attainment of the equality in (9) is out of reach. For instance, suppose that $\theta\mapsto\pi_{x}(\theta)$ is bounded away from 0 for some $x$. This is rare in applications, but not impossible. It may happen that the data is not particularly informative about a feature $\Phi=m(\Theta)$ of $\Theta$, and then the corresponding marginal IM (see Section 4.3) has a non-vanishing contour, i.e., the data is insufficiently informative to exclude any values of $\Phi$ with high degree of confidence. In any case, a probability distribution $\mathsf{Q}_{x}^{\star}$ satisfying (9), if one exists, will be called the inner probabilistic approximation of the possibility measure $\overline{\Pi}_{x}$ (Dubois et al., 2004).
Note that $C_{\alpha}(x)$ in (5) is a $100(1-\alpha)$credible set associated with an inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ and, since the IM’s validity property implies that it’s also a $100(1-\alpha)$confidence set, it follows that $\mathsf{Q}_{x}^{\star}$ exactly—not just asymptotically—achieves the “probability matching” property (e.g., Staicu and Reid, 2008; Datta and Ghosh, 1995) desired of no-prior Bayes solutions.
Theorem 2.
$\mathsf{Q}_{x}^{\star}$-credible sets are confidence sets.
What does it take to satisfy (9)? From the characterization (8) in Theorem 1, it is relatively easy to see that equality in (9) amounts to a choice of kernel $\mathsf{K}_{x}^{\beta}$ that is fully supported on $\partial C_{\beta}(x)$, the boundary of $C_{\beta}(x)$, and marginal $\mathsf{M}_{x}={\sf Unif}(0,1)$. How the kernel allocates its mass to $\partial C_{\beta}(x)$ isn’t determined by the equality in (9), hence the inner probabilistic approximation of $\overline{\Pi}_{x}$ isn’t unique. But the sets $C_{\beta}(x)$ are typically bounded, i.e., contained in a compact subset of $\mathbb{T}$, and in fact, when the sample size is moderate to large, they’re approximately ellipsoids (Section 4.2). So, in such cases, choosing the kernel $\mathsf{K}_{x}^{\beta}$ to be a uniform distribution on $\partial C_{\beta}(x)$, for each $\beta\in(0,1)$, is quite natural, and consistent with both the maximum entropy (e.g., Cover and Thomas, 2006; Bernardo and Smith, 1994) and indifference principles (e.g., Keynes, 1921; Jaynes, 2003).
To summarize, my proposed inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of the possibilistic IM is a uniform mixture of uniform distributions. More specifically, a sample $(\Theta\mid x)\sim\mathsf{Q}_{x}^{\star}$ can be drawn via the following two steps:
  1. 1.
    sample $A\sim{\sf Unif}(0,1)$, and
  2. 2.
    sample $(\Theta\mid x,A=\alpha)\sim\mathsf{K}_{x}^{\alpha}:={\sf Unif}\{\partial C_{\alpha}(x)\}$.
Step 2 is far simpler to state than it is to carry out, and I’ll discuss how this computational challenge can be addressed—at least approximately—in Section 5.
As a quick illustration, consider a gamma model $\mathsf{P}_{\theta}={\sf Gamma}(n,\theta)$ for a single observation $X$, with $n$ a fixed constant; the more practical case of $n$ iid exponential samples where $X$ is the sum, the complete sufficient statistic, reduces to this one. Then the likelihood-based possibilistic IM has contour function given by
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\bigl{\{}(X/\theta)^{n}e^{-X/\theta}\leq(x/\theta)^{n}e^{-x/\theta}\bigr{\}},\quad\theta>0.$
The right-hand side has no simple closed-form expression but, since $X/\theta$ is a pivot under $\mathsf{P}_{\theta}$, the computations are straightforward using Monte Carlo; a complicated exact expression, whose evaluation doesn’t require Monte Carlo, is given in Proposition 1 of Hector et al., (2025). The contour shown in Figure 1 is actually the one above, with $n=7$ and $x=14$. Clearly, the level sets $C_{\alpha}(x)$ are intervals, say, $[a(x,\alpha),b(x,\alpha)]$, where $b\geq a$, but with $a(x,1)=b(x,1)=x/n$, the maximum likelihood estimator. Then the kernel $\mathsf{K}_{x}^{\alpha}$ in the inner probabilistic approximation is a probability distribution supported on the endpoints $a(x,\alpha)$ and $b(x,\alpha)$, i.e.,
$\mathsf{K}_{x}^{\alpha}=w_{x,\alpha}\,\delta_{a(x,\alpha)}+(1-w_{x,\alpha})\,\delta_{b(x,\alpha)},$
where $w_{x,\alpha}\in[0,1]$ and $\delta_{z}$ denotes a probability measure that puts all its mass at $z$. The question I want to address here is: what kind of distribution is an inner probabilistic approximation, what does it look like? In the case where $w_{x,\alpha}\equiv w$ is a fixed value that doesn’t depend on $\alpha$, the distribution function of $\Theta\sim\mathsf{Q}_{x}^{\star}$ can easily be derived:
$\mathsf{Q}_{x}^{\star}(\Theta\leq t)=\begin{cases}w\,\pi_{x}(t)&\text{if $t\leq x/n$}\\ 1-(1-w)\,\pi_{x}(t)&\text{if $t>x/n$}.\end{cases}$
Two such distribution functions are shown in Figure 2(a), one based on $w=0.5$ (solid) and another based on $w\approx 0.45$ (dashed). A third distribution function (dashed) is also shown, namely, the Bayesian posterior distribution function based on a default (right Haar) prior with density $\theta\mapsto\theta^{-1}$, which is an inverse gamma distribution. This Bayes posterior doesn’t obviously have a form like that above, [^2] I wasn’t able to find a constant $w$ that would recover the inverse gamma distribution—the choice $w\approx 0.45$ was my best effort. Consequently, this particular instantiation of the general form of $\mathsf{Q}_{x}^{\star}$ must involve an $\alpha$-dependent weight and it’d be interesting to determine what that $w_{x,\alpha}$ is. but that it too is an inner probabilistic approximation of the IM is a general fact established in Section 4.1 below. The three distribution functions are similar, but not exactly the same. To confirm that all three are, indeed, inner probabilistic approximations, I plot, in Figure 2(b), the non-coverage probability functions $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$. According to (9), the non-coverage probability for each inner probabilistic approximation should be equal to $\alpha$, which is confirmed in the plot.
(a) Distribution function of $\mathsf{Q}_{x}^{\star}$
(b) Non-coverage probability of $\mathsf{Q}_{x}^{\star}$
Figure 2: Plots related to three inner probabilistic approximations for the gamma illustration with $n=7$ and $x=14$: uniform kernel (solid), non-uniform kernel with $w\approx 0.45$ (dashed), and the Bayesian posterior distribution (dotted) based on the default prior $\theta\mapsto\theta^{-1}$. Panel (a) shows the cumulative distribution function $t\mapsto\mathsf{Q}_{x}^{\star}(\Theta\leq t)$ and Panel (b) shows the non-coverage probability function $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$.
4 Properties
4.1 Agrees with Bayes in group invariant models
Let $\mathscr{G}$ denote a group of transformations from $\mathbb{X}$ to $\mathbb{X}$, with function composition $\circ$ as the binary operation. As is customary, I’ll write $gx$ for the image of $x\in\mathbb{X}$ under transformation $g\in\mathscr{G}$; and if $g_{1}$ and $g_{2}$ are two group elements, then $g_{1}\circ g_{2}$ denotes their composition. Since $\mathscr{G}$ is a group, it’s associative, i.e., $g_{1}\circ(g_{2}\circ g_{3})=(g_{1}\circ g_{2})\circ g_{3}$ for all $g_{1},g_{2},g_{3}\in\mathscr{G}$, it contains the identity transformation $e$, and for every $g\in\mathscr{G}$, there exists an inverse $g^{-1}\in\mathscr{G}$ such that $g\circ g^{-1}=g^{-1}\circ g=\text{identity}$. Examples of transformation groups include translation, scaling, rotation, permutation, etc.
The group $\mathscr{G}$ connects to the statistical model as follows. Suppose that, for each $g\in\mathscr{G}$ and each $\theta\in\mathbb{T}$, there exists a corresponding $\bar{g}\theta\in\mathbb{T}$ such that
$\mathsf{P}_{\theta}(gX\in\cdot)=\mathsf{P}_{\bar{g}\theta}(X\in\cdot),\quad(\theta,g)\in\mathbb{T}\times\mathscr{G}.$ (10)
For example, if the distribution of $X$ depends on a location parameter $\theta$, then the distribution of $X+\tau$ depends on parameter $\theta+\tau$. When the statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ satisfies (10), it’s called an invariant statistical model. Let $\overline{\mathscr{G}}$ denote the collection of transformations $\bar{g}$ on $\mathbb{T}$, corresponding to the mappings $g\in\mathscr{G}$ and parameters $\theta\in\mathbb{T}$; it’s easy to check that $\overline{\mathscr{G}}$ is a group too. More details on group invariant models can be found in Fraser, (1968), Eaton, (1989), and Schervish, (1995, Ch. 6), among others. The necessary technical background is given in Appendix B.
When the model is invariant in the sense described above, there are standard and broadly agreed upon no-prior Bayes solutions, namely, those based on combining the likelihood with the so-called right Haar prior using Bayes’s formula (e.g., Kass and Wasserman, 1996). This also agrees with the solution obtained via Fisher’s fiducial argument, among others. In Martin, 2023a , I showed that, under the above conditions, there is a close connection between the possibilistic IM and this standard no-prior Bayes solution. In the present context, the natural interpetation of the result established there is that the standard no-prior Bayes solution—namely, the Bayes posterior with respect to the invariant right Haar prior—is an inner probabilistic approximation of the possibilistic IM. Further details, including a new proof, are provided in Appendix B.
Theorem 3 (Martin, 2023a ).
For an invariant statistical model, under the conditions laid out in Appendix B, the Bayesian posterior distribution relative to the right Haar prior is an inner probabilistic approximation of the possibilistic IM.
That is, under an invariant statistical model, the familiar Bayes/fiducial distribution is an inner probabilistic approximation of the possibilistic IM. This justifies my previous claim that the Bayes posterior in the gamma illustration, as shown in Figure 2 above, is an inner probabilistic approximation. This connection also serves as a strong proof of concept for the no-prior Bayes solution that I’m proposing here: that is, the right Haar prior Bayesian solutions in such problems are reasonable and widely used, so what I’m proposing must be equally reasonable and usable. Finally, it also helps to explain certain phenomena from the Bayesian perspective, namely, why the Bayesian/fiducial solutions in these problems are exactly probability matching.
A point made above that deserves to be mentioned again is that the inner probabilistic approximations are not unique. So, the Bayes solution here might not exactly correspond to the generically recommended strategy that defines $\mathsf{K}_{x}^{\beta}$ to be uniform on $\partial C_{\beta}(x)$; the illustration in Figure 2 is such a case. In fact, when a direct connection between a Bayesian solution and the IM’s inner probabilistic approximation can be made, then I might recommend the former since its computation is more familiar than the latter’s. But, like in Figure 2, the difference between these inner probabilistic approximations will be negligible and will vanish as the information in data increases; see Section 4.2.
As an illustration, consider a case involving directional data, where the observations correspond to directions—points on a unit circle in the plane—or simply angles with respect to a reference direction. Real-world applications involving these kind of data include wind and animal movement studies; see Mardia and Jupp, (2000) for a comprehensive treatment. More generally, direction measurements can be represented by points on the surface of a hypersphere. This kind of data is common in astronomy, where the position of planets, stars, etc. can be described by a point on the celestial sphere. The present example involves $n=25$ such planar directional observations, plotted in Figure 3. These data are simulated from a model very similar ot that used to analyze a very small data set on roulette wheel spins from Example 1.1 in Mardia and Jupp, (2000). The posited model is the so-called von Mises distribution, parametrized by a mean direction/angle $\theta$ and a concentration parameter $\kappa$. Further details about this model and the corresponding IM/Bayes/fiducial solutions are given in Appendix B.2. Here I’ll take the concentration parameter $\kappa=4$ to be fixed and known. I’ll also interpret the data as angles (radians) so that the parameter $\theta$ acts like a location paramter. With this simplification, the $\theta$-only von Mises model is invariant with respect to addition modulo $2\pi$. Therefore, Theorem 3 implies that the Bayes/fiducial distribution is an inner probabilistic approximation of the possibilistic IM. Figure 4 shows, first, a plot of the possibilistic IM’s contour for these data, with its peak at the maximum likelihood estimator, $\hat{\theta}_{x}=1.35$ and, second, a histogram of samples drawn from the inner probabilistic approximation (with $\mathsf{K}_{x}^{\beta}$ a uniform distribution) and the no-prior Bayesian posterior density overlaid. The similarities provided a visual confirmation of the claim in Theorem 3.
Figure 3: Plot of the directional data analyzed in Section 4.1. Points on the circle determine the direction. The small closed circle denotes the sample mean in Cartesian coordinates and the corresponding angle determined by that point, $\hat{\theta}_{x}=1.35$, is the maximum likelihood estimator of the mean angle. The length of that vector, 0.87, is a to-be-conditioned-on ancillary statistic; see Appendix B.2.
(a) Possibilistic IM contour
(b) Inner probabilistic approximation
Figure 4: Analysis of the directional data in Section 4.1. Panel (a) shows the possibilistic IM contour and Panel (b) shows a histogram of samples from the corresponding inner probabilistic approximation with the flat-prior Bayes posterior density overlaid.
4.2 Asymptotic normality and efficiency
This section presents a version of the celebrated Bernstein–von Mises theorem for the possibilistic IM, which establishes its asymptotic normality and efficiency. That is, for large $n$, the possibilistic IM’s contour resembles a Gaussian possibility contour, defined below, with covariance matrix matches the familiar Cramér–Rao lower bound. This implies a traditional Bernstein–von Mises theorem for the proposed inner probabilistic approximation and, therefore, it agrees with any other reasonable no-prior Bayesian solution asymptotically. Moreover, this near-Gaussian form of the IM output offers some valuable insights and simplifications relevant to computing the inner probabilistic approximation; see Section 5. The first order of business is to define Gaussian possibility.
Let $g_{m,v}$ denote the $D$-dimensional Gaussian probability density function, parametrized by a mean vector $m\in\mathbb{R}^{D}$ and a covariance matrix $v\in\mathbb{R}_{+}^{D\times D}$. Define the corresponding Gaussian possibility measure as the outer possibilistic approximation (e.g., Dubois and Prade, 1990) of ${\sf N}_{D}(m,v)$, i.e., the possibility measure with the smallest credal set that contains ${\sf N}_{D}(m,v)$. Analogous to the formula (2) presented in Section 2, this is just the probability-to-possibility transform applied to the aforementioned Gaussian distribution, and the corresponding possibility contour is
$\gamma_{m,v}(y)=\mathsf{P}\{g_{m,v}(Y)\leq g_{m,v}(y)\},\quad y\in\mathbb{R}^{D},\quad Y\sim{\sf N}_{D}(m,v).$
This can easily be rewritten in a beneficial and effectively closed-form expression,
$\gamma_{m,v}(y)$ $=1-F_{D}\bigl{(}(y-m)^{\top}v^{-1}(y-m)\bigr{)},$
where $F_{D}$ is the ${\sf ChiSq}(D)$ distribution function. Write $\overline{\Gamma}_{m,v}$ for the corresponding possibility measure obtained by maximizing the contour as described above. Other authors have used a different form of Gaussian possibility (e.g., Denœux, 2023b ; Denœux, 2023a ), but I find the one above appealing and convenient because of the following duality between ${\sf N}_{D}(m,v)$ and $\overline{\Gamma}_{m,v}$: the former is the inner probabilistic approximation of the latter, and the latter is the outer possibilistic approximation of the former.
For some intuition about how this is related to the possibilistic IMs, consider a simple Gaussian location problem, $X\sim\mathsf{P}_{\Theta}={\sf N}_{D}(\Theta,v)$, with known covariance matrix $v$. Then the relative likelihood is $R(x,\theta)=\exp\{-\frac{1}{2}(x-\theta)^{\top}v^{-1}(x-\theta)\}$, quadratic on the log scale, and, since $\|X-\theta\|^{2}\sim{\sf ChiSq}(D)$ under $\mathsf{P}_{\theta}$, it follows that possibilistic IM’s contour is
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}=1-F_{D}\bigl{(}(x-\theta)^{\top}v^{-1}(x-\theta)\bigr{)},\quad\theta\in\mathbb{R}^{D}.$
This is exactly the Gaussian possibility contour with mean $x$ and covariance matrix $v$. More generally, by the famous theorem of Wilks, (1938), the relative likelihood $R(X^{n},\theta)$, as a function of iid samples $X^{n}=(X_{1},\ldots,X_{n})$ from $\mathsf{P}_{\theta}$, will be asymptotically distributed as $\exp\{-\frac{1}{2}{\sf ChiSq}(D)\}$. Then the following approximation is expected when $n$ is large,
$\pi_{x^{n}}(\theta)\approx 1-F_{D}\bigl{(}(\hat{\theta}_{x^{n}}-\theta)^{\top}J_{x^{n}}(\hat{\theta}_{x^{n}}-\theta)\bigr{)},\quad\theta\in\mathbb{R}^{D},$
where $\hat{\theta}_{x^{n}}$ is the maximum likelihood estimator and $J_{x^{n}}$ is the observed Fisher information matrix. The reader will notice that the right-hand side above is exactly the Gaussian possibility contour with mean vector $\hat{\theta}_{x^{n}}$ and covariance matrix $J_{x^{n}}^{-1}$, hence the connection between the possibilistic IM and the Gaussian possibility is very natural and holds quite generally, at least approximately. Making this connection rigorous, however, requires considerable care, both in formulating the proper regularity conditions and in suitably bounding all the error terms. Along these lines, under regularity conditions comparable to those advanced in Le Cam, (1956, 1960, 1970)—see, also, van der Vaart, (1998) and Bickel et al., (1998)Martin and Williams, (2025) proved the following novel Bernstein–von Mises theorem for the possibilistic IM.
Theorem 4 (Martin and Williams, 2025).
Under the regularity conditions presented in Appendix C, the possibilistic IM’s contour $\pi_{X^{n}}$ satisfies
$\sup_{\theta\in\mathcal{T}}\bigl{|}\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\bigr{|}\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$},$
where $\mathcal{T}$ is a fixed but arbitrary compact subset of $\mathbb{T}$, $\gamma_{X^{n}}(\theta)=\gamma_{\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1}}(\theta)$ is the centered and scaled Gaussian possibility contour, $I_{\Theta}$ the Fisher information, and $\Delta_{\Theta}(X^{n})$ a sequence of random variables that converges in distribution to ${\sf N}_{D}(0,I_{\Theta}^{-1})$. If, in addition, the likelihood function is twice differentiable, then
$\sup_{u\in\mathcal{U}}\bigl{|}\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}u)-\gamma(u)\bigr{|}\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$},$
where $\mathcal{U}$ is a fixed but arbitrary compact subset of $\mathbb{R}^{D}$, and $\gamma$ is the standard Gaussian possibility with zero mean and identity covariance matrix.
Note that the covariance matrix in the Gaussian limit agrees with the Cramér–Rao lower bound. It’s in this sense that the possibilistic IM is asymptotically efficient and, hence, there’s no cost in terms of efficiency for the IM’s exact validity (and imprecision).
A direct consequence of Theorem 4 is that the IM’s inner probabilistic approximation converges to a Gaussian probability distribution just like in the classical, Bayesian Bernstein–von Mises theorem. The proof is in Appendix C.
Theorem 5.
Under the conditions of Theorem 4, the approximation $\mathsf{Q}_{X^{n}}^{\star}$ of the possibilistic IM satisfies a probabilistic Bernstein–von Mises theorem, i.e.,
$d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}}):=\sup_{H}\bigl{|}\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr{|}\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$},$
where $\mathsf{G}_{X^{n}}={\sf N}_{D}(\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1})$ is the approximating Gaussian distribution and the supremum is over all Borel measurable subsets $H$ of $\mathbb{T}$.
I’ll give a quick sketch of the proof for a scalar parameter $\Theta$. For a generic value $\theta$, if the level set $C_{\alpha}(x^{n})$ is the interval $[a_{\alpha}(x^{n}),b_{\alpha}(x^{n})]$, then it’s easy to verify that
$\{a_{\alpha}(x^{n})\geq\theta\}\;\text{or}\;\{b_{\alpha}(x^{n})\leq\theta\}\iff\pi_{x^{n}}(\theta)\leq\alpha.$
Using the mixture characterization of the inner probabilistic approximation, one can verify that the cumulative distribution function of $\Theta$, relative to $\mathsf{Q}_{x^{n}}^{\star}$, at $\theta$ is
[…]
=$=\begin{cases}\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta\leq\hat{\theta}_{x^{n}}$}\\ 1-\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta>\hat{\theta}_{x^{n}}$}.\end{cases}$
Switching to local coordinates, i.e., $\theta=\hat{\theta}_{x^{n}}+J_{x^{n}}^{-1/2}z$ for $z$ fixed (or bounded in $n$), the case $\theta\leq\hat{\theta}_{x^{n}}$ corresponds to $z\leq 0$ and Theorem 4 implies
$\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)$ $\to\tfrac{1}{2}\,\gamma(z)$
$=\tfrac{1}{2}\,\mathsf{P}(Z^{2}\geq z^{2}),\quad\text{for $Z\sim{\sf N}(0,1)$}$
$=\mathsf{P}(Z\leq z).$
Similarly, if $z>0$, then $1-\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)\to\mathsf{P}(Z\leq z)$. Therefore, with probability tending to 1, the distribution of $J_{X^{n}}^{1/2}(\Theta-\hat{\theta}_{X^{n}})$, under $\Theta\sim\mathsf{Q}_{X^{n}}^{\star}$, is converging to a standard normal, identical to the classical Bernstein–von Mises theorem.
A point that follows immediately from the above sketch is that assigning probability $\frac{1}{2}$ to the boundary points in the inner probabilistic approximation is crucial; of course, any choice that assigns probability to the boundary points that converges to $\frac{1}{2}$ works fine too. More generally, since $\mathsf{G}_{X^{n}}$ is an inner probabilistic approximation to the possibility measure with contour $\gamma_{X^{n}}$, and corresponds to taking a uniform kernel on the level sets of $\gamma_{X^{n}}$, it’s necessary that the choice of kernel defining $\mathsf{Q}_{X^{n}}^{\star}$ at least be asymptotically uniform. This implies that the inner probabilistic approximation’s non-uniqueness vanishes as $n\to\infty$. This observation adds further support for my recommendation above to adopt the uniform kernel in the inner probabilistic approximation.
4.3 Marginalization risks, or lack thereof
Except for the impossibly rare cases where a genuine prior distribution is available, reliable statistical inference is inherently imprecise—those familiar test and confidence procedures that control error rates all have an imprecise probabilistic characterization (Martin, 2021a ). The point is that there’s no single probability distribution that can reliably quantify uncertainty about the unknowns in a statistical model. So, insisting that uncertainty quantification be probabilistic has risks:
[Xie and Singh, (2013)] are thus recommending that we ignore the restriction to confidence sets or equivalent, and free confidence to allow the production of parameter distributions. Certainly distributions are easier to think about, are largely in accord with Fisher’s original proposal, and are more in the freedom of the Bayes approach, but they do overlook inherent risks… (Fraser, 2013)
Fraser’s “risks” primarily concern marginalization (e.g., Dawid et al., 1973; Fraser, 2011a ; Balch et al., 2019). Roughly, whatever desirable properties a no-prior Bayes solution for $\Theta$ has, the corresponding no-prior Bayes solution for $\Phi=m(\Theta)$ obtained by probabilistic marginalization generally doesn’t share those desirable properties. When Schweder and Hjort, (2013) warn that “joint [confidence distributions] should not be sought, we think, since they might easily lead the statistician astray,” they’re concerned about users being tempted to do these familiar probabilistic operations, thus creating a risk of unreliability. Fraser, 2011a and Fraser et al., (2016) highlight non-linearity in $m$ as a major contributor to this risk; see, also, Martin, 2023a ; Martin, 2024b and the discussion below.
There are different ways to transform a quantification of uncertainty about $\Theta$ to a one about $\Phi=m(\Theta)$. Start with a confidence set $C_{\alpha}(x)$. The image $C_{\alpha}^{m}(x)=m\{C_{\alpha}(x)\}$ is a confidence set for $\Phi$ as a result of the following implication:
$\Theta\in C_{\alpha}(x)\implies m(\Theta)\in C_{\alpha}^{m}(x).$ (11)
The opposite implication generally doesn’t hold, which is the origin of the aforementioned risk. So, while the transformed confidence set for $\Phi$ maintains the nominal coverage probability enjoyed by the original confidence for $\Theta$, it can be conservative.
Next, take a no-prior Bayes posterior $\mathsf{Q}_{x}$ for $\Theta$ and map it to the corresponding marginal posterior $\mathsf{Q}_{x}^{m}$ for $\Phi=m(\Theta)$ via the probability calculus. An advantage of this approach is its convenience: draw samples of $\Phi$ from $\mathsf{Q}_{x}^{m}$ by first sampling $\Theta$ from $\mathsf{Q}_{x}$ and then transforming them via $m$. But, by (11):
$\mathsf{Q}_{x}^{m}\{C_{\alpha}^{m}(x)\}\geq\mathsf{Q}_{x}\{C_{\alpha}(x)\},\quad\alpha\in[0,1].$ (12)
That is, $\mathsf{Q}_{x}^{m}$ is inclined to assign more of its mass to the marginal confidence sets than $\mathsf{Q}_{x}$ assigned to the original confidence sets. So, if $\mathsf{Q}_{x}$ was a confidence distribution (CD), i.e., it assigns probability $1-\alpha$ to $C_{\alpha}(x)$ for each $\alpha\in[0,1]$, then (12) suggests calling $\mathsf{Q}_{x}^{m}$ an over-confidence distribution (oCD). Illustrations of this over-confidence phenomenon are presented in Appendix D. This can be problematic because surely the user of $\mathsf{Q}_{x}^{m}$ will disregard the safe confidence set $C_{\alpha}^{m}(x)$ and, instead, construct his/her own quantification of uncertainty directly based on the over-confident $\mathsf{Q}_{x}^{m}$, risking unreliability. This begs the question: when does $\mathsf{Q}_{x}^{m}$’s credible set agree with the safe $C_{\alpha}^{m}(x)$?
Let $\overline{\Pi}_{x}$ be the possibilistic IM as described above, quantifying uncertainty about $\Theta$, given $X=x$. One way—but not the only way, see Section 6—to carry out possibilistic marginalization is by applying the general extension operation (e.g., Zadeh, 1975a ; Zadeh, 1975b ; Zadeh, 1978). Extension yields a new marginal IM $\overline{\Pi}_{x}^{m}$ for $\Phi=m(\Theta)$ with possibility contour
$\pi_{x}^{m}(\phi)=\sup_{\theta:m(\theta)=\phi}\pi_{x}(\theta),\quad\phi\in m(\mathbb{T}).$ (13)
If the confidence set $C_{\alpha}(x)$ in question is determined by the contour $\pi_{x}$ as in (5), then its image $C_{\alpha}^{m}(x)$ under $m$ can be expressed as
$C_{\alpha}^{m}(x)=\{m(\theta):\pi_{x}(\theta)>\alpha\}=\{\phi:\pi_{x}^{m}(\phi)>\alpha\}.$
That the possibilistic marginalization strategy appears organically in the common-sense treatment of confidence sets is no coincidence; see Martin, 2021a . The inner probabilistic approximations of $\overline{\Pi}_{x}$ are confidence distributions, and these are characterized as mixtures of kernels supported on the boundaries of the confidence sets. Then the derived marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ clearly also can be characterized by a similar mixture, but $\mathsf{Q}_{x}^{m}$ would only be the inner probabilistic approximation of $\overline{\Pi}_{x}^{m}$ if the boundaries $\partial C_{\alpha}(x)$ of the original confidence sets get mapped via $m$ to the boundaries $\partial C_{\alpha}^{m}(x)$ of the marginal confidence sets. Define $m$ to be boundary-preserving if $m\{\partial C_{\alpha}(x)\}=\partial C_{\alpha}^{m}(x)$ for each $\alpha\in[0,1]$. Among other forms of mappings, linear transformations are boundary preserving because they attain their extrema on the boundary of their domain.
Theorem 6.
If $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\overline{\Pi}_{x}$, and $m$ is a boundary-preserving mapping, then the marginal distribution $\mathsf{Q}_{x}^{m}$ for $\Phi=m(\Theta)$ derived from $\mathsf{Q}_{x}$ is the inner probabilistic approximation of $\overline{\Pi}_{x}^{m}$.
That the probability calculus leads to reliable marginalization effectively only when the transformation is linear doesn’t imply that reliable marginalization is out of reach in the non-linear case. It just means non-probabilistic marginalization strategies are needed. This isn’t unfamiliar: Bayesians choose default prior distributions that are tailored to the particular inferential target $\Phi$ (e.g., Tibshirani, 1989; Datta and Ghosh, 1995; Bernardo, 1979; Berger et al., 2009), and confidence and fiducial distribution users carry out various manipulations akin to those used to construct approximate pivotal quantities. The above insights motivate the following suggestion: first carry out the appropriate possibilistic marginalization to get $\pi_{x}^{m}$ and $\overline{\Pi}_{x}^{m}$, and then construct the inner probabilistic approximation of this. An example of this using the extension-based marginalization is shown below. Extension-based marginalization is a simple and very general, but is not the most efficient. Typically, sharper inferences can be drawn when marginalization is carried out by “validifying” a relative profile likelihood. Details of this profile-based approach will be presented in the context of a concrete example in Section 6 below.
Here I’ll revisit the directional data example presented in Section 4.1 above, where the quantity of interest is $\Phi=\cos(a\Theta)$, for two different values of $a$. Start with the case of $a=1$. Since the Bayes/fiducial distribution—which is also an inner probabilistic approximation of the possibilistic IM by Theorem 3—is almost fully supported on the interval $(0,\pi)$, where $\theta\mapsto\cos(\theta)$ is one-to-one, this mapping is effectively boundary-preserving. Then Theorem 6 implies that the marginal Bayes/fiducial distribution for $\Phi$ is an inner probabilistic approximation of the marginal possibilistic IM for $\Phi$. Figure 5 shows, for $a=1$, the extension-based marginal IM contour for $\Phi$, samples of the marginal Bayes/fiducial with the corresponding density function overlaid, and the non-credibility function $\alpha\mapsto 1-\mathsf{Q}_{x}^{m}\{C_{\alpha}^{m}(x)\}$, for $\mathsf{Q}_{x}^{m}$ equal to the Bayes/fiducial marginal distribution and to the inner probabilistic approximation of the extension-based marginal IM. The key observation is that the contour and the posterior distribution have similar shapes and, as Theorem 6 predicts, the non-credibility function behaves exactly as it should for an inner probabilistic approximation, according to the definition (9). Again, this all happens because $\theta\mapsto\cos(a\theta)$ with $a=1$ is approximately boundary-preserving on the interval where the relevant posterior distributions are supported.
(a) Marginal IM contour
(b) Bayes marginal
(c) Non-credibility
Figure 5: Marginal inference results for $\Phi=\cos(a\Theta)$ with $a=1$ in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
Next, take $a=1.5$. Then the aforementioned “effectively boundary-preserving” property no longer holds and, surprisingly, this seemingly small change in the quantity of interest has non-negligible impact on the results. Figure 6 shows the same plots as in Figure 5, this time for $a=1.5$. The key observation is that, while the general shape of the marginal IM contour and the Bayes posterior distribution were similar in Figure 5, they are quite different in Figure 6; this is a consequence of the entirely different style of marginalization, from $\Theta$ to $\Phi$, with the IM’s possibilistic marginalization preserving the validity properties enjoyed by both approaches’ inference on $\Theta$ whereas the Bayesians’ probabilistic marginalization doesn’t. The IM’s inner probabilistic approximation (not shown) looks similar to the Bayesian posterior, but it’s not the same. Indeed, Figure 6(c) shows the non-credibility for the Bayes posterior distribution and the IM’s inner probabilistic approximation, and there are clear signs of the Bayesians’ over-confidence. The IM’s inner approximation also shows (milder) signs of over-confidence, but there’s a perfectly reasonable explanation for this. Recall from Section 3 that exact inner probabilistic approximations need not exist, e.g., if the possibility contour is bounded away from 0. In Figure 6(a), note that the contour is in fact bounded away from 0 and, consequently, the sets $C_{\alpha}^{m}(x)$ are equal to the full space $[-1,1]$ for sufficiently small $\alpha$. Since any probability distribution $\mathsf{Q}_{x}^{m}$ supported on $[-1,1]$ must assign probability $1>1-\alpha$ to $C_{\alpha}^{m}(x)=[-1,1]$ for all sufficiently small $\alpha$, there is no $\mathsf{Q}_{x}^{\star}$ that satisfies (9) for all $\alpha$. But the IM’s pseudo inner probabilistic approximation—“pseudo” because the inner probabilistic approximation strictly doesn’t exist—quickly corrects for this over-confidence at small $\alpha$ and achieves the objective (9) for all $\alpha$ that are not too small.
(a) Marginal IM contour
(b) Bayes marginal
(c) Non-credibility
Figure 6: Marginal inference results for $\Phi=\cos(a\Theta)$ with $a=1.5$ in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
5 Computation
Until recently, only naive and relatively inefficient strategies for computing the IM contour were available. In particular, the go-to strategy was to approximate $\pi_{x}$ via
$\pi_{x}(\theta)\approx\frac{1}{M}\sum_{m=1}^{M}1\{R(X_{m,\theta},\theta)\leq R(x,\theta)\},\quad\theta\in\mathbb{T},$ (14)
where $X_{m,\theta}$ are independent copies of the data $X$, drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at a few different $\theta$ values, but all the practically relevant calculations—e.g., identifying the confidence set in (5)—require evaluation over a sufficiently fine grid covering the relevant portions of $\mathbb{T}$, and this is expensive. Recently, Martin, (2025) developed a new and more efficient strategy for doing the relevant IM computations that replaces most of the naive contour evaluations as in (14) with Monte Carlo sampling from a sort of “posterior distribution.” That distribution is, in fact, the inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\overline{\Pi}_{x}$ defined in Section 3 above. Therefore, the same computational strategy for (approximately) sampling from $\mathsf{Q}_{x}^{\star}$ advanced in Martin, (2025) can be reIMagined here as an implementation of the new no-prior Bayes solution.
The strategy in Martin, (2025) is, roughly, to stitch together a suitable collection of Gaussian possibilistic IMs. While the “stitched” collection of Gaussian possibilities isn’t a Gaussian possibility, the use of a basic Gaussian building block can be justified by the asymptotic normality result in Section 4.2. Interest here is solely in the inner probabilistic approximation, so it’s not necessary for me to describe all the details involved in the “stitched IM.” For $\xi$ a generic $D$-vector of positive numbers, consider a collection of Gaussian confidence/level sets
$C_{\alpha}^{\xi}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}\,J_{x}^{\xi}\,(\theta-\hat{\theta})\leq\chi_{D,1-\alpha}^{2}\},$
where $\chi_{D,1-\alpha}^{2}$ is the $1-\alpha$ quantile of the ${\sf ChiSq}(D)$ distribution, and $J_{x}^{\xi}$ is an embellished version of the observed Fisher information matrix with spread at least partially controlled by the parameters in $\xi$. Specifically, if $J_{x}$ has spectral decomposition $J_{x}=E\Lambda E^{\top}$, then the embellished version $J_{x}^{\xi}$ is defined as
$J_{x}^{\xi}=E\,\text{diag}(\xi^{-1})\,\Lambda\,\text{diag}(\xi^{-1})\,E^{\top},$ (15)
where $\text{diag}(\xi^{-1})$ is the diagonal matrix with elementwise reciprocal of the $\xi$ vector on the diagonal. Then $\xi$ controls confidence set’s size by manipulating the eigenvalues of the underlying observed Fisher information matrix, e.g., a large (resp. small) entry in $\xi$ stretches (resp. contracts) the level set in the direction of the corresponding eigenvector. Theorem 4 says that the original $C_{\alpha}(x)$ will be roughly elliptical, so it’s safe to assume that there exists a vector $\xi$ such that the following inclusion holds snugly:
$C_{\alpha}(x)\subseteq C_{\alpha}^{\xi}(x).$ (16)
Of course, this vector would generally depend on $(x,\alpha)$, so I’ll denote this as $\xi(x,\alpha)$. A cartoon depiction of the proposed strategy is shown in Figure 7. The motivation for bounding $C_{\alpha}(x)$ inside $C_{\alpha}^{\xi}(x)$ is that it’s not possible to describe the former exactly but the latter is very simple, at least for fixed $\xi$. The data-driven algorithm recommended in Martin, (2025), relying on the foundations laid in Cella and Martin, (2024), for selecting $\xi(x,\alpha)$ is presented in Appendix E. Once this $\xi(x,\alpha)$ has been determined, the inner probabilistic approximation uses the kernel
$\mathsf{K}_{x}^{\alpha}={\sf Unif}\{\partial C_{\alpha}^{\xi(x,\alpha)}(x)\},\quad\alpha\in(0,1).$
Thanks to the direction of the inclusion in (16), the proposed implementation of the inner probabilistic approximation is conservative. So, the relevant properties enjoyed in theory by the inner probabilistic approximation of the IM should also be at least approximately enjoyed by the practical proposal above; see Section 6.
[…]
Figure 7: Approximating the IM contour’s $\alpha$-level set $C_{\alpha}(x)$ (gray) by an ellipse $C_{\alpha}^{\xi}(x)$ with a “good” choice of $\xi$ (solid) and with a “bad” choice (dashed).
6 Example: Behrens–Fisher problem
For brevity, I only give one detailed illustration of the inner probabilistic approximation here, but two additional examples are presented in Appendix F.
Few examples have deeper practical and historical relevance than the Behrens–Fisher problem, whose name comes from Fisher’s reference to a technical result of Walter Behrens in his proposed fiducial solution to the problem (Fisher, 1935a ; Fisher, 1939). The original problem is simple to state: independent samples of size $n_{1}$ and $n_{2}$, respectively, are available from two distinct normal populations, ${\sf N}(\Theta_{11},\Theta_{12}^{2})$ and ${\sf N}(\Theta_{21},\Theta_{22}^{2})$, with $\Theta=(\Theta_{11},\Theta_{12},\Theta_{21},\Theta_{22})$ unknown, and the goal is marginal inference on the difference $\Phi=m(\Theta)=\Theta_{21}-\Theta_{11}$ of the two means. The problem is straightforward if the two variances are known or if their ratio is known, but the fully unknown variance case has remained elusive, i.e., lots of candidate solutions are available but, despite its apparent simplicity and practical relevance, there is no consensus on which solution is “right” or “best.” For a review of the problem and various solutions, see Kim and Cohen, (1998).
By far the most widely used solution to the Behrens–Fisher problem is the modification of the basic Student-t pivot with the degrees of freedom approximations due to Welch, (1938, 1947); this is implemented in R’s t.test function. Other standard approaches include the simple-but-conservative solution proposed by Hsu, (1938) and Scheffé, (1970), and the Bayesian solution proposed by Jeffreys, (1940), based on the right Haar prior for $\Theta$, which is mathematically equivalent to Fisher’s fiducial solution. Ironically, the solution proposed by Jeffreys is different from the Bayesian solution based on the Jeffreys prior, both in its construction and its performance. A conservative IM solution was advanced in Martin and Liu, 2015b , but here I take a different approach.
To set the scene, start by considering inference on the full parameter $\Theta$. This model has lots of structure so the relative likelihood-based, possibilistic IM construction for inference on $\Theta$ is conceptually and computationally straightforward. Thanks to the model’s underlying affine group invariance, it follows from Theorem 3 that the corresponding inner probabilistic approximation for $\Theta$ is exactly the Bayesian posterior distribution for $\Theta$ based on the right Haar prior, which is also Fisher’s fiducial distribution for $\Theta$. Since the mapping from $\Theta$ to $\Phi$ is linear, it follows from Theorem 6 that the (mathematically equivalent) Bayes and fiducial solutions proposed by Jeffreys and Fisher, respectively, correspond to the marginal distribution for $\Phi$ derived from the IM’s inner probabilistic approximation; moreover, these are also the inner probabilistic approximation to the marginal IM based on following the extension principle.
The new approach I’m proposing here first carries out a different—and generally more efficient—IM marginalization strategy, based on validifying the profile relative likelihood. This idea was first developed in Martin, 2023b and an illustration of its use in the Behrens–Fisher problem is considered in Example 5 there. Roughly speaking, the difference between this profile-based marginal IM construction and the aforementioned extension-based construction is where the marginalization is carried out: the former first eliminates nuisance parameters in the relative likelihood and then constructs the IM contour directly for $\Phi$, whereas the latter first constructs the IM contour for $\Theta$ and then marginalizes to $\Phi$. It had been conjectured that the profile-based marginal IM construction was more efficient than the extension-based construction, and Martin and Williams, (2025) recently confirmed this, at least asymptotically. My proposal here is to extract the inner probabilistic approximation of this profile-based marginal possibilistic IM for $\Phi$. This requires no change in the computational approach outlined in Section 5, and the exact validity of the marginal IM implies, e.g., that credible intervals for $\Phi$ based on this “posterior” are exact confidence intervals. The only wrinkle is that the relative profile likelihood doesn’t have a closed-form expression in the Behrens–Fisher problem, and its distribution depends on a nuisance parameter; this is precisely what makes the Behrens–Fisher problem challenging. This increases the computational cost of the inner probabilistic approximation compared to other examples (see Appendix F), but this pays off in terms of efficiency gains as I demonstrate below.
The go-to real-data illustration of the Behrens–Fisher problem is the example in Lehmann, (1975, p. 83) on travel times to work via two different routes. The relevant summary statistics—sample sizes, sample means, and sample standard deviations—are as follows: $n_{1}=5$, $\hat{\theta}_{11}=7.580$, and $\hat{\theta}_{12}=2.237$; $n_{2}=11$, $\hat{\theta}_{21}=6.136$, and $\hat{\theta}_{22}=0.073$. The wide discrepancy between the two standard deviations, $\hat{\theta}_{12}$ and $\hat{\theta}_{22}$, would make it difficult to justify assuming the two variances are equal. Figure 8 shows a histogram of the samples of $\Phi$ from the inner probabilistic approximation of the marginal IM, with (kernel estimates) of the density functions from the right Haar and Jeffreys prior Bayes solutions overlaid; the former also agrees with Fisher’s fiducial distribution. The point is that all three distributions are similar, with the new inner probabilistic approximation and the right Haar prior posterior being a bit more diffuse than the Jeffreys prior posterior.
Figure 8: Histogram summarizing the distribution of $\Phi$ in the Behrens–Fisher problem based on the marginal IM’s inner probabilistic approximation. The solid and dashed lines are (kernel estimates of) the Bayesian posterior distributions for $\Phi$ based on the right Haar and Jeffreys priors, respectively.
For further comparison, I run a small simulation study. I focus on a rather severely imbalanced case—with $n_{1}=3$ and $n_{2}=20$—to ensure that the differences in performance are apparent. Otherwise, the simulation settings are standard: $\Theta_{11}=2$, $\Theta_{21}=0$, $\Theta_{12}^{2}=1$, and $\Theta_{22}^{2}=2$. I generated 10000 samples under this setup, and Table 1 presents the coverage probability and expected length of the various 90% confidence intervals for $\Phi$. Notably, only the marginal possibilistic IM’s inner probabilistic approximation is able to hit the target coverage probability nearly on the nose, and, as desired, it’s more efficient than the valid-but-conservative Bayes/fiducial solution based on the right Haar prior.
Method Coverage probability Mean length
Bayes with Jeffreys prior 0.861 2.31
Bayes with right Haar prior 0.929 3.26
t.test with Welch df 0.881 2.79
Marginal IM’s inner approximation 0.904 2.71
Table 1: Coverage probabilities and mean lengths for the 90% confidence intervals returned by the various methods. Standard errors for all of the coverage probabilities and mean lengths are roughly 0.003 and 0.01, respectively.
7 Conclusion
This paper offers a new perspective on no-prior Bayesian inference, one that has certain close ties to Bayesian inference but is not actually Bayesian at all. My approach starts with an inferential model (IM) framework for data-driven uncertainty quantification that prioritizes reliability, insisting on calibration of its data-dependent degrees of belief. The kind of calibration IMs require makes them incompatible with probabilistic Bayesian inference, and it’s for this reason that the IM’s output possibilistic, i.e., takes the mathematical form of a possibility measure. However, if one desires probabilistic uncertainty quantification, then the IM can accommodate this with an inner probabilistic approximation of its possibilistic output. Aside from being exactly probability matching, the proposed solution agrees with existing no-prior Bayes/fiducial solutions in group invariant problems where there’s an overall consensus, and is also asymptotically efficient by a version of the celebrated Bernstein–von Mises theorem. And there’s literally no choice of prior distribution required: for the given model and data, the possibilistic IM is what it is, so it’s just a matter of extracting the inner probabilistic approximation.
The characterization of the contents of the IM’s credal set in Theorem 1 suggests a strategy for evaluating, at least approximately/conservatively, the inner probabilistic approximation via Monte Carlo. I explored this thoroughly in Martin, (2025), but mostly in the context of possibilistic IMs. Here I use this same strategy for my proposed reIMagined no-prior Bayesian inference, and I show that this leads to a new and broadly reliable solution to the technically challenging and practically relevant Behrens–Fisher problem. Nothing about this approach was tailored to the Behrens–Fisher problem specifically, so I fully expect similarly good performance in many other important applications. Surely, the computational efficiency of my proposed solution can be improved, so I welcome advancements along these lines from computationally savvy readers.
The focus here has been on no-prior solutions, but there are cases in which incomplete or partial prior information is available, including high-dimensional problems in which structural assumptions like sparsity are common. Bayesian solutions can’t work with incomplete prior information—a prior distribution, perhaps diffuse, must be specified for every aspect of the unknown $\Theta$. Genuine partial-prior possibilistic IMs have recently been developed (e.g., Martin, 2022b ; Martin, 2023b ), and an interesting idea would be to apply the same inner probabilistic approximation proposed here in the no-prior case to the partial-prior case in the above references. Just as in this paper, these inner probabilistic approximations would enjoy some of the inherent reliability properties satisfied by the partial-prior IM, but the details remain to be investigated.
Acknowledgments
This work is partially supported by the U.S. National Science Foundation, under grants SES–2051225 and DMS–2412628.
Appendix A Further IM insights, intuition, etc.
In the main text, I gave only the mathematical definition and briefly explanation of the possibilistic inferential model (IM) and its basic properties. But these developments are deeper and more insightful than can be communicated in such a brief explanation, so I’ll share a bit more here about the IM intuition and logic. I’ll organize this presentation as a list of answers to some common/frequently-asked questions.
Question 1.
Why is imprecision needed? The false confidence theorem (Balch et al., 2019) says that for any precise data-dependent probability distribution $\mathsf{Q}_{X}$, e.g., a Bayesian or fiducial posterior, and for any thresholds $\rho$ and $\tau$, there exists a hypothesis $H$ such that
$H\not\ni\Theta\quad\text{and}\quad\mathsf{P}_{\Theta}\{\mathsf{Q}_{X}(H)>\tau\}>\rho.$
This is called false confidence because, with $\tau$ large and $\rho$ not small, the posterior tends (at non-negligible rate $\rho$) to be confident (high probability $\tau$) in the truthfulness of $H$ even though $H$ is actually false. There’s a sense in which this result is trivial: take $H$ equal to the complement of $\Theta$, so that $H$ is false and, at least in the absolutely continuous case, $\mathsf{Q}_{X}(H)=1$ almost surely. But there’s more nuance here because it’s not just for these extreme/trivial $H$ that false confidence manifests; it’s not even just for “big” sets that happen to be false. The shape of $H$ and the effects of non-linearity, just like in Section 4.3 in the present paper, play an important albeit not yet well understood role.
In addition, there’s no reason why the size of a hypothesis should matter at all—data are compatible with a hypothesis as soon as they’re compatible with any single point in the hypothesis. Size/measure only becomes relevant when integration enters the picture, but there’s no justification for making this step except that statisticians can’t help it, the theory of integration is burned into our minds. Ideally, integration could be replaced by some other calculus that doesn’t rely on notions of size, no differential element.
Fortunately, both of the issues identified above can be resolved at once if ordinary probability can be replaced by a suitable imprecise probability. The proof of the false confidence theorem leverages the additivity of the posterior, so, if the additivity assumption is dropped, then the proof breaks down. But, also, certain data-dependent imprecise probabilities, namely, (possibilistic) IMs, have been shown to be immune to false confidence. Moreover, possibility theory specifically operates under a different calculus than integration: the possibilistic calculus is optimization, so there is no differential element and the size of hypotheses is irrelevant when it comes to if and to what extent they are compatible with data; see Martin, 2024a .
Question 2.
Why the possibilistic form specifically? It helps to start with being clear about what the objectives are when it comes to the formulation of a new and improved framework for statistical uncertainty quantification, so I list here three desiderata:
  1. 1.
    Mathematically rigorous. Statistical reasoning shouldn’t be based on a patchwork disparate pieces but, rather, a coherent whole from which an answer to any question about the relevant unknowns can be obtained.
  2. 2.
    As simple and “familiar-looking” as possible. Much of the theory that is currently being used for statistical inference (e.g., p-values for tests) generally works. So, a new framework doesn’t need to be completely different and, in fact, it shouldn’t give altogether different solutions to classical inference problems.
  3. 3.
    Reliable. Not only should procedures (tests, confidence sets, etc) derived from ones uncertainty quantification be reliable in the sense that they control frequentist error rates, but the (lower and upper) probabilities and expectations that it assigns to relevant hypotheses or assessments should be calibrated too. In fact, it’s this reliability property that gives the uncertainty quantification its meaning.
To my knowledge, the only options for a mathematically rigorous framework of uncertainty quantification are those that can be described by an imprecise probability of some sort. Possibility theory is arguably one of the simplest imprecise probability theories and, since the possibilistic IM framework that I’ve been advancing has close connections to p-values for likelihood ratio tests, confidence distributions, and now no-prior Bayes solutions, it clearly satisfies the first two desiderata.
The main text highlighted some of the possibilistic IM framework’s reliability properties, but there is a more to say as it concerns the topic of this section: why the possibilistic form? The property I referred to as validity in the main paper—Equation (4)—is actually quite strong. The frequentist error rate control guarantees it implies for test and confidence set procedures derived from the IM output are important, but doesn’t do justice to the strength of this property. As Martin, 2022a and Cella and Martin, (2023) have shown, the validity property in Equation (4) of the main paper is equivalent to
$\sup_{\Theta\in\mathbb{T}}\mathsf{P}_{\Theta}\{\overline{\Pi}_{X}(H)\leq\alpha\text{ for some $H$ with $H\ni\Theta$}\}\leq\alpha,\quad\alpha\in[0,1].$ (17)
This is similar to the property stated in Equation (6) of the main text, but with one key difference: the property in (6) establishes that the IM is calibrated “pointwise” in $H$, whereas (17) shows that, in fact, the IM is calibrated “uniformly” in $H$. This is relevant because modern statisticians and data scientists aren’t quantifying uncertainty for themselves, they’re developing general tools and software to assist other users in quantifying their statistical uncertainty. Since I can’t possibly know how a user might want to do with the uncertainty quantification framework that I offer them, if I want to guarantee that it’s safe—which is important both for science/society and for me as the developer of such a tool—then I must ensure that calibration is maintained even if the user is, e.g., peeking at the data first to decide what hypotheses to consider. The pointwise validity result in (6) of the main paper doesn’t guarantee this, but the uniform validity result in (17) does.
There are mathematical obstacles that would make it difficult for any non-possibilistic quantification of uncertainty to satisfy the strong validity property in Equation (4) of the main paper, and hence the uniform calibration property in (17) above. Indeed, if $(\underline{\Omega}_{x},\overline{\Omega}_{x})$ was some other imprecise-probabilistic quantification of uncertainty about $\Theta$, given $X=x$, then one could define the analogue of its “contour,” i.e., $\theta\mapsto\overline{\Omega}_{x}(\{\theta\})$, but this may not meet the requirement of a possibility contour, that is, generally,
$\sup_{\theta\in\mathbb{T}}\overline{\Omega}_{x}(\{\theta\})<1,\quad\text{and often $\ll 1$}.$
Nothing in imprecise probability theory says that this “contour” must satisfy this supremum-equals-1 property and, in fact, in many (non-statistical) cases there are good reasons why such a property would be undesirable. But, as Shafer, (1976, p. 226) explains, statistical inference problems are special and well-suited for this supremum-equals-1 property. What I can add to Shafer’s justification is that it is virtually impossible for anything other than a possibilistic quantification of uncertainty, one for which the supremum-equals-1 property holds, to satisfy the validity property as stated in Equation (4) of the main paper and, therefore, the equivalent uniform calibration property in (17) above. The point is that, if $\overline{\Omega}_{X}(\{\Theta\})$ is bounded away from 1 for almost all $X$, then it can’t be stochastically no less than ${\sf Unif}(0,1)$. Moreover, if one was able to find a non-possibilistic IM $(\underline{\Omega}_{x},\overline{\Omega}_{x})$ that does satisfy the validity property in Equation (4) in the main text, then there is a possibilistic IM that is valid and no less efficient. So, there is at least no loss of generality in focusing on IMs that take on the possibilistic form.
Question 3.
Beyond assigning beliefs to hypotheses about model parameters? Classical inference on the parameters is a fully-specified statistical model is the right context to describe this paper’s contribution, namely, a new perspective on no-prior Bayesian inference, but there are many other kinds of problems and statistical tasks that scientists are interested in. Fortunately, IMs can contribute along these lines too, and they already have. First, it’s often the case that the quantity of interest is not best described as the parameter of a statistical model, but as some feature (or functional) of the underlying distribution $\mathsf{P}$, which doesn’t have a familiar parametric form. The construction of approximately valid IMs for such cases is presented in Cella and Martin, 2022a ; work on removing the “approximately” adjective is ongoing. Prediction is also an important problem and usually requires a nonparametric touch. IM advances along these lines can be found in Cella and Martin, 2022b , with a more possibility-centric perspective offered in Martin, 2023b (, Sec. 6.3). Sometimes inference isn’t the objective, the focus is on decision-making, i.e., choosing an appropriate (data-driven) action. In Martin, 2021b , I propose the use of Choquet integration to use the IM output to evaluate upper (and lower) expected losses and recommend choosing actions that minimize the upper expected loss. I also show that the IM’s inherent reliability in inference carries over to the decision problem as well: it’s a provably rare event that the IM’s upper expected loss assessment is substantially more optimistic than an oracle’s, uniformly over actions. Finally, novel and provably reliable IM solutions for model assessment and selection, information fusion (e.g., meta-analysis), etc. are currently in development.
Appendix B Details from Section 4.1
B.1 Theory
This section provides some additional background and details about the invariant model setting and the result in Theorem 3.
Recall that $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ is a family of probability distributions supported on $\mathbb{X}$ that is also invariant with respect to a group of transformations $\mathscr{G}$ from $\mathbb{X}$ to itself. If the distributions $\mathsf{P}_{\theta}$ have densities with respect to some underlying dominating measure, which I’ll assume here, then invariance means
$p_{\theta}(x)=p_{\bar{g}\theta}(gx)\,\chi(g),\quad x\in\mathbb{X},\;\theta\in\mathbb{T},\;g\in\mathscr{G},$ (18)
where $\chi(g)$ is the “multiplier,” a change-of-variables Jacobian term; see Schervish, (1995, Ch. 6) and Eaton, (1989) for details. A further simplification that I’ll adopt here, common in the literature (e.g., Schervish, 1995, p. 371), is to assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$.
The set $\mathscr{G}x=\{gx:g\in\mathscr{G}\}\subseteq\mathbb{X}$ is called the orbit of $\mathscr{G}$ corresponding to $x$. The orbits partition $\mathbb{X}$ into equivalence classes, so every point $x\in\mathbb{X}$ falls on exactly one orbit. This partition can be used to construct a new coordinate system on $\mathbb{X}$ which will be useful for us in what follows. Identify $x\in\mathbb{X}$ with $(g_{x},u_{x})$, where $u_{x}\in\mathbb{U}$ denotes the label of orbit $\mathscr{G}x$ and $g_{x}\in\mathscr{G}$ denotes the position of $x$ on the orbit $\mathscr{G}x$.
Assumptions.
Let $\{p_{\theta}:\theta\in\mathbb{T}\}$ be a family of densities invariant with respect to a locally compact topological group $\mathscr{G}$ in the sense of (18) and, for simplicity, assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$. In addition, assume that the following hold:
  • A1.
    The left Haar measure $\lambda$ and the corresponding right Haar measure $\rho$ on (the Borel $\sigma$-algebra of) $\mathscr{G}$ exist and are unique up to scalar multiples.
  • A2.
    There exists a bijection $t:\mathbb{X}\to\mathscr{G}\times\mathbb{U}$, with both $t$ and $t^{-1}$ measurable, that maps $x\in\mathbb{X}$ to its position–orbit coordinates $(g_{x},u_{x})\in\mathscr{G}\times\mathbb{U}$.
  • A3.
    The distribution of $t(X)=(G,U)\in\mathscr{G}\times\mathbb{U}$ induced by the distribution of $X\sim\mathsf{P}_{\theta}$ has a density with respect to $\lambda\times\mu$ for some measure $\mu$ on $\mathbb{U}$.
A couple quick remarks are in order. First, for A1, existence and uniqueness of the left and associated right Haar measures is a classical result; see Halmos, (1950) and Nachbin, (1965). For A3, existence of a joint density with respect to a product measure ensures that there’s no difficulty in defining a conditional distribution for $G$, given $U=u$, which is needed below. Finally, $U=U_{X}$, as a function of $X$, is an ancillary statistic.
The simplest example is a location model where $\mathscr{G}=\overline{\mathscr{G}}=(\mathbb{R},+)$. This is an abelian group, so the left and right Haar measures are the same and both equal to Lebesgue measure. The function $x\mapsto t(x)$ in A2 consists of two components: in its “$g_{x}$” coordinate an equivariant function of $x$ that estimates the location and, in its “$u_{x}$” component, an invariant function of $x$, such as residuals. For example, $g_{x}=\bar{x}$ the arithmetic mean of $x=(x_{1},\ldots,x_{n})$ and $u_{x}=\{x_{i}-\bar{x}:i=1,\ldots,n\}$. Note that the $u_{x}$ coordinate satisfies a constraint, so, after it’s represented in a suitable lower-dimensional space $\mathbb{U}$, $\mu$ can be taken as Lebesgue measure there. There are many other problems that fit this general form; see, e.g., Chapters 1–2 of Fraser, (1968), including the exercises.
Since I’m assuming $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$, generic $\theta\in\mathbb{T}$ can be identified as transformations in $\mathscr{G}$; the same goes for the uncertain $\Theta$. That means $\theta$ can be inverted to $\theta^{-1}$ and can be composed with other transformations in $\mathscr{G}$. A key result (e.g., Schervish, 1995, Cor. 6.64) is that the density of $X$ or, equivalently, of $t(X)=(G,U)$, under $\mathsf{P}_{\theta}$, is given by
$p_{\theta}(g,u)=f(\theta^{-1}\circ g,u),$ (19)
where the function $f:\mathscr{G}\times\mathbb{U}\to\mathbb{R}$ doesn’t directly depend on $\theta$. The particular form of $f$ isn’t important for what follows.
  • Lemma 6.65 in Schervish, (1995) gives a Bayesian posterior distribution $\mathsf{Q}_{x}^{\text{\sc b}}$ for $\Theta$ that has the following density with respect to the right Haar prior measure:
    $q_{x}^{\text{\sc b}}(\theta)\propto f(\theta^{-1}\circ g,u),\quad\theta\in\mathbb{T},\quad x\equiv(g,u).$ (20)
    This also agrees with the accepted fiducial distribution for $\Theta$, given $x$.
  • Proposition 2 in Martin, 2023a shows that the relative likelihood is given by
    $R(x,\theta)=d_{u}\,f(\theta^{-1}\circ g,u),\quad\theta\in\mathbb{T},\quad x\equiv(g,u),$
    where $d_{u}$ depends only on the $u$-component of $x$.
These points are crucial to the following calculation. Indeed, since the $U$ component of $X$ is ancillary, the minimum complexity principle in Martin, 2022b says the IM construction should condition on the observation $U=u$. Write $\pi_{g|u}(\theta)$ for the possibilistic IM’s contour, where the subscript is meant to emphasize that the conditional distribution of $X\equiv(G,U)$, given the observed value of $u$, is used in the validification step (2). Then the IM contour is defined as
$\pi_{g|u}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\mid U=u\},\quad\theta\in\mathbb{T}.$
Then the following sequence of identities are consequences of the two results above that link relative likelihood and the posterior density:
$\pi_{g|u}(\theta)$ $=\mathsf{P}_{\theta}\{f(\theta^{-1}\circ G,u)\leq f(\theta^{-1}\circ g,u)\mid U=u\}$
$=\mathsf{P}\{f(H,u)\leq f(\theta^{-1}\circ g,u)\mid U=u\}$
$=\mathsf{Q}_{x}^{\text{\sc b}}\{f(H,u)\leq f(\theta^{-1}\circ g,u)\}$
$=\mathsf{Q}_{x}^{\text{\sc b}}\{q_{x}^{\text{\sc b}}(\Theta)\leq q_{x}^{\text{\sc b}}(\theta)\},$
where the first line is by simplification of the definition of $\pi_{g|u}$; the second line is by the fact that $H:=\theta^{-1}\circ G$ is a pivot with respect to the conditional distribution of $G$, given $U=u$, under $\mathsf{P}_{\theta}$ (e.g., Schervish, 1995, Cor. 6.67); the third line is by the fact that $H:=\Theta^{-1}\circ g$ is also a pivot with respect to the Bayes posterior $\Theta\sim\mathsf{Q}_{x}^{\text{\sc b}}$ and has the same distribution as $H$ in previous line; and the last line is by simplification via (20).
The last line in the above display can be recognized as the probability-to-possibility transform applied to the Bayesian/fiducial posterior distribution. This identifies the posterior distribution $\mathsf{Q}_{x}^{\text{\sc b}}$ as the inner probabilistic approximation of the possibilistic IM $\overline{\Pi}_{x}=\overline{\Pi}_{g|u}$, and hence a new/different proof of the main result in Martin, 2023a , which is also Theorem 3 above.
B.2 Directional data example
Let $X=(X_{1},\ldots,X_{n})$ be the observable angles for $n$ independent spins of a roulette wheel. The model assumed here is the von Mises distribution, with density
$p_{\theta}(x)=\{2\pi I_{0}(\kappa)\}^{-1}\exp\{\kappa\cos(x-\theta)\},\quad x,\theta\in[0,2\pi).$
Here $\kappa>0$ is a known concentration parameter, and $I_{0}$ denotes the Bessel function of order 0. The unknown mean angle $\Theta$ is the parameter to be inferred from the data.
Digging into this model further, using the cosine formula
$\cos(x-\theta)=\cos x\cos\theta+\sin x\sin\theta,$
it’s easy to see that the minimal sufficient statistic for this model is the pair
$Y=(\bar{C},\bar{S})=\Bigl{(}\frac{1}{n}\sum_{i=1}^{n}\cos X_{i},\,\frac{1}{n}\sum_{i=1}^{n}\sin X_{i}\Bigr{)}.$
Relative to the posited von Mises model, $Y$ can be treated as “the data.” Since $(C_{i},S_{i})=(\cos X_{i},\sin X_{i})$ denotes the Cartesian coordinates of the point on the unit circle corresponding to the angle $X_{i}$, it’s clear that $Y=(\bar{C},\bar{S})$ is generally inside the unit circle. The more “concentrated” the original data points are around a particular angle, the closer $Y$ will be to the boundary of the circle. Convert this average position to polar coordinates:
$G=\arctan(\bar{S}/\bar{C})\quad\text{and}\quad U=(\bar{C}^{2}+\bar{S}^{2})^{1/2}.$
Here $G$ is the angle the vector $(\bar{C},\bar{S})$ makes with the horizontal axis and $U$ is the Euclidean length of $(\bar{C},\bar{S})$. It’s clear that $(G,U)$ is a bijection of the minimal sufficient statistic $Y$, so write $t(X)=(G,U)$. I’m using the notation $G$ and $U$ to align with that introduced in the previous section. That is, there is a group of transformations $\mathscr{G}$ that corresponds to rotations of points in the plane; technically, $\mathscr{G}$ is the special orthogonal group $\text{SO}(2)$ consisting of orthogonal matrices with unit determinant. Then $G$ can be interpreted as both an angle or as a rotation by that angle. Similarly, the interior of the unit circle can be partitioned into distinct concentric circles, which correspond to orbits, and $U$ determines which of these orbits $Y$ sits on.
The details here align with the general setup above, so I can proceed to carry out the fiducial (or default-prior Bayes) and IM analyses. It’s a standard result (e.g., Mardia and Jupp, 2000, Eq. 4.5.5) that the joint density for $t(X)=(G,U)$ factors as
$p_{\theta}(g,u)$ $=p_{\theta}(g\mid u)\,p(u)$
=$=\{2\pi I_{0}(\kappa u)\}^{-1}\exp\{\kappa u\cos(g-\theta)\}\,p(u),$
where $p(u)$ denotes the marginal density for $U$, which doesn’t depend on $\theta$ and isn’t relevant here. The expression for the conditional density of $G$, given $U=u$, is easily seen to be that of a von Mises distribution but with concentration parameter $\kappa u$. The influence of conditioning on $U$ makes intuitive sense because, as indicated above, the statistic $U$ acts like a measure of how “concentrated” the observed angles are around a common angle. It’s the conditional density above that drives both the fiducial and IM analyses. Indeed, the no-prior Bayes (and fiducial) density is
$q_{x}(\theta)\propto\exp\{\kappa u\cos(g-\theta)\},\quad\theta\in(0,2\pi],$
and, after simplification, the IM’s possibility contour is
$\pi_{g|u}(\theta)=\mathsf{P}\{\cos H\leq\cos(g-\theta)\mid U=u\},\quad\theta\in[0,2\pi),$
where the probability is with respect to the conditional distribution of $H:=G-\Theta$, given $U=u$, which is simply a von Mises distribution with mean angle 0 and concentration parameter $\kappa u$. This contour function is easy to evaluate numerically and to plot as in the main text. By setting a sufficiently fine grid on the interval $(0,2\pi)$ and then evaluating the contour on that grid, it’s easy to obtain the level set $C_{\alpha}(x)$ for any $\alpha$. From here it’s straightforward to sample from the inner probabilistic approximation, etc.
Appendix C Details from Section 4.2
C.1 Setup and regularity conditions
Certain regularity conditions are required in order to establish asymptotic concentration properties of estimators, posterior distributions, etc., and the same is true for IMs and the inner probabilistic approximations under consideration here. Roughly, these conditions ensure that the log-likelihood function is smooth enough that it can be well-approximated by a quadratic function. One common set of regularity conditions are the classical Cramér conditions (Cramér, 1946), versions of which can be found in the standard texts, including Lehmann and Casella, (1998, Theorem 3.10) and Schervish, (1995, Theorem 7.63). Here I’ll adopt the more modern set of sufficient conditions originating in Le Cam, (1956, 1960, 1970) and Hájek, (1972); see, also, Bickel et al., (1998) and van der Vaart, (1998).
The posited model specifies a class $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ of probability distributions, supported on $\mathbb{X}$, with $\mathsf{P}_{\theta}$ having a density function $p_{\theta}(x)$ relative to a $\sigma$-finite measure $\nu$ on $\mathbb{X}$. Then the data $X^{n}=(X_{1},\ldots,X_{n})$ is assumed to be iid of size $n$ with common distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the uncertain “true value” of the parameter. Following Bickel et al., (1998, Ch. 2), define the (natural) logarithm and square-root density functions:
$\ell_{\theta}(x)=\log p_{\theta}(x)\quad\text{and}\quad s_{\theta}(x)=p_{\theta}(x)^{1/2}.$
The “dot” notation, e.g., $\dot{g}_{\theta}(x)$, represents a function that behaves like the derivative of $g_{\theta}(x)$ with respect to $\theta$ for pointwise in $x$. If the usual partial derivative of $g_{\theta}(x)$ with respect to $\theta$ exists, then $\dot{g}_{\theta}(x)$ is that derivative; but suitable functions $\dot{g}_{\theta}(x)$ may exist even when the ordinary derivative fails to exist. In particular, below I’ll assume existence of a suitable “derivative” $\theta\mapsto\dot{s}_{\theta}(x)$ of the square-root density. Finally, let $\mathcal{L}_{2}(\nu)$ denote the set of measurable functions on $\mathbb{X}$ that are square $\nu$-integrable.
Assumptions.
The parameter space $\mathbb{T}$ is open and there exists a vector $\dot{s}_{\theta}(x)=\{\dot{s}_{\theta,d}(x):d=1,\ldots,D\}$, with $\dot{s}_{\theta,d}$ an element of $\mathcal{L}_{2}(\nu)$, such that the following conditions hold:
  • A1.
    the maps $\theta\mapsto\dot{s}_{\theta,d}$ from $\mathbb{T}$ to $\mathcal{L}_{2}(\nu)$ are continuous for each $d=1,\ldots,D$;
  • A2.
    at each $\theta\in\mathbb{T}$,
    $\int\bigl{|}s_{\theta+u}(x)-s_{\theta}(x)-u^{\top}\dot{s}_{\theta}(x)\bigr{|}^{2}\,\nu(dx)=o(\|u\|^{2}),\quad u\to 0\in\mathbb{R}^{D};$ (21)
  • A3.
    and the $D\times D$ matrix $\int\dot{s}_{\theta}(x)\,\dot{s}_{\theta}(x)^{\top}\,\nu(dx)$ is non-singular for each $\theta\in\mathbb{T}$.
The condition in (21) is often described as $\theta\mapsto s_{\theta}$ being differentiable in quadratic mean. Note that this condition does not require the square-root density to actually be differentiable at $\theta$, only that it be locally “linearizable” in an average sense, like a differentiable-at-$\theta$ function would be. The classical Cramér conditions assume more than two continuous derivatives, so (21), which does not even require existence of a first derivative, is significantly weaker; sufficient conditions for (21) are given in van der Vaart, (1998, Lemma 7.6). Then the score function ˙$\dot{\ell}_{\theta}(x)$ is defined in terms of $\dot{s}_{\theta}(x)$ as
[…]
and it can be shown that ˙$\int\dot{\ell}_{\theta}(x)\,\mathsf{P}_{\theta}(dx)=0$ for each $\theta$. Moreover, Condition A3 above implies non-singularity of the Fisher information matrix $I_{\theta}=\int\dot{\ell}_{\theta}(x)\ell_{\theta}(x)^{\top}\,\mathsf{P}_{\theta}(dx)$ for each $\theta\in\mathbb{T}$. Bickel et al., (1998, Prop. 2.1.1) provide sufficient conditions for A1–A3.
The above conditions are rather mild, and hold in a broad range of problems, including exponential families. One further condition is required for the possibilistic Bernstein–von Mises theorem in Section 4.2, namely, that the maximum likelihood estimator is consistent, i.e., $\hat{\theta}_{X^{n}}\to\Theta$ in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$. This, is not automatic, but holds quite broadly when the parameter space dimension is fixed, as I’m assuming here.
C.2 Proof of Theorem 5
The intuition behind the Bernstein–von Mises result itself, and the proof that follows, is that both the inner probabilistic approximation and the Gaussian that it’s merging with are fully determined by corresponding possibility contours. So, any difference in these distributions will imply a difference in the possibility contours, but the contours themselves are merging by Theorem 4 and, therefore, any difference in the distributions must be vanishing asymptotically. The following details make this argument rigorous.
I’ll start by justifying the above claim that any difference between $\mathsf{Q}_{X^{n}}^{\star}$ and the Gaussian approximation $\mathsf{G}_{X^{n}}$ is due to differences in the corresponding possibility contours. The first observation is that $\mathsf{Q}_{X^{n}}^{\star}$ has the characterization as a uniform mixture of uniform kernels by Theorem 1. The second observation is that, similarly, $\mathsf{G}_{X^{n}}$ is the inner probabilistic approximation of the Gaussian possibility measure with contour $\gamma_{X^{n}}$ and, therefore, the characterization result in Theorem 1 applies to $\mathsf{G}_{X^{n}}$ too. That is, if $\mathsf{K}_{X^{n},\pi}^{\alpha}$ and $\mathsf{K}_{X^{n},\gamma}^{\alpha}$ are the uniform kernels supported on the level sets $\{\theta:\pi_{X^{n}}(\theta)=\alpha\}$ and $\{\theta:\gamma_{X^{n}}(\theta)=\alpha\}$, then
[…]
Then, for any Borel measurable subset $H$ of $\mathbb{T}$,
$\bigl{|}\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr{|}\leq\int_{0}^{1}\bigl{|}\mathsf{K}_{X^{n},\pi}^{\alpha}(H)-\mathsf{K}_{X^{n},\gamma}^{\alpha}(H)\bigr{|}\,d\alpha,$
so $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies that there exists an $H$ such that $|\mathsf{K}_{X^{n},\pi}^{\alpha}(H)-\mathsf{K}_{X^{n},\gamma}^{\alpha}(H)|$ exceeds $\varepsilon$ on a set of $\alpha$’s having positive Lebesgue measure. This implies that there exists an open subset of $H$ and a non-empty interval $(\alpha_{1},\alpha_{2})$ such that
$\pi_{X^{n}}(\theta)\in(\alpha_{1},\alpha_{2})\quad\text{and}\quad\gamma_{X^{n}}(\theta)>\alpha_{2}\quad\text{for all $\theta$ in that subset;}$;
of course, the last inequality could be replaced by “$\gamma_{X^{n}}(\theta)<\alpha_{1}$ and the positions of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ in the above display could be reversed, but these changes don’t affect the argument that follows. What matters is that there exists a point $\theta$—depending implicitly on data $X^{n}$, on $H$, and on $\varepsilon$—such that
$|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\alpha_{2}-\alpha_{1}>0.$
That Theorem 4 implies the merging of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ uniformly means that the event in the above display has vanishing $\mathsf{P}_{\Theta}$-probability as $n\to\infty$.
What’s not fully satisfactory about the above conclusion is that the positive bound “$\alpha_{2}-\alpha_{1}$” is only implicit. An important point that the above argument reveals is that any differences in how $\mathsf{Q}_{X^{n}}^{\star}$ and $\mathsf{G}_{X^{n}}$ assign probability will manifest through the contour level sets. Since these are also the sets that determine the distributions themselves, it’s without loss of generality to consider hypotheses $H$ that take the form of lower level sets of either of the defining contours. For concreteness, consider
$H=\{\theta:\pi_{X^{n}}(\theta)\leq\tau\},\quad\tau\in(0,1).$
By definition of the inner probabilistic approximation, $\mathsf{Q}_{X^{n}}^{\star}(H)=\sup_{\theta\in H}\pi_{X^{n}}(\theta)=\tau$. Moreover, since $H$ is bounded and excludes the mode of $\gamma_{X^{n}}$, it’s easy to check that $\mathsf{G}_{X^{n}}(H)$ satisfies
$\inf_{\theta\in\partial H}\gamma_{X^{n}}(\theta)\leq\mathsf{G}_{X^{n}}(H)\leq\sup_{\theta\in\partial H}\gamma_{X^{n}}(\theta).$
Consequently,
$\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)$ $\leq\tau-\inf_{\theta\in\partial H}\gamma_{X^{n}}(\theta)$
$=\sup_{\theta\in\partial H}\{\tau-\gamma_{X^{n}}(\theta)\}$
$=\sup_{\theta\in\partial H}\{\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\}$
$\leq\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|,$
where the second-to-last line follows because $\pi_{X^{n}}(\theta)\equiv\tau$ on $\partial H$. Similarly,
$\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)$ $\geq\tau-\sup_{\theta\in\partial H}\gamma_{X^{n}}(\theta)$
$=\inf_{\theta\in\partial H}\{\tau-\gamma_{X^{n}}(\theta)\}$
$=\inf_{\theta\in\partial H}\{\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\}$
$\geq-\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|.$
Putting these two inequalities together gives
$|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|\leq\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|.$
Therefore,
$|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon\implies\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon,$
and, by Theorem 4, the latter event has vanishing $\mathsf{P}_{\Theta}$-probability as $n\to\infty$.
Putting everything together, $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies existence of a set $H$ for which $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$. This implies a non-negligible difference in the contours so, without loss of generality, the problematic $H$—whose existence is guaranteed—can be taken as a lower level set corresponding to one of the contours. When $H$ is such a lower level set, $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$ implies that $|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon$ for all $\theta$ on the boundary of $H$. Since the latter event has vanishing $\mathsf{P}_{\Theta}$-probability by Theorem 4, so must the event $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$, thus proving Theorem 5.
Appendix D Details from Section 4.3
A claim was made in Section 4.3 that confidence distributions (CDs)—or no-prior Bayes solutions in general—often transform into over-confidence distributions (oCDs) when (non-linear) marginalization is carried out. Here I show just a couple simple illustrations to justify this claim. While my details are different, the insights were inspired by the analysis in Fraser, 2011a .
The first concerns a single sample $X$ from a simple Gaussian model with known variance, say $\sigma^{2}=1$, but unknown mean $\Theta$. Of course, there is an exact confidence interval $C_{\alpha}(x)=x\pm{\tt qnorm}(1-\frac{\alpha}{2})$ and the standard CD, $\mathsf{Q}_{x}^{\star}={\sf N}(x,1)$, assigns exactly probability $1-\alpha$ to $C_{\alpha}(x)$. But now suppose the goal is marginal inference on $\Phi=m(\Theta)$ where $m$ is bounded and, hence, non-linear, e.g.,
$m(\theta)=\max\{-1,\min(1,\theta)\},\quad\theta\in\mathbb{R}.$ (22)
The claim is that the marginal CD of $m(\Theta)$ will assign more than probability $1-\alpha$ to the image $m\{C_{\alpha}(x)\}$ of the original confidence interval under $m$. Figure 9 shows the (marginal) posterior probability assigned to $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for several different realizations $x$. That the curve generally falls above the diagonal line is a demonstration of the aforementioned over-confidence.
(a) $x=0$
(b) $x=0.5$
(c) $x=1$
(d) $x=2$
Figure 9: Plots of the marginal posterior probability assigned $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for several different $x$, where $m$ is the bounded function (22).
As a second example, suppose that $X=(X_{1},X_{2})$ is a single realization from a bivariate normal with mean vector $\Theta=(\Theta_{1},\Theta_{2})$ and identity covariance matrix. Again, it’s straightforward to construct a confidence disc $C_{\alpha}(x)$ and a confidence distribution $\mathsf{Q}_{x}^{\star}={\sf N}_{2}(x,I)$. Consider marginalization via the familiar non-linear function $m(\theta)=\|\theta\|^{2}$. I say “familiar” because this is an infamous example going to back at least to Stein, (1959) and more recently in Fraser, 2011a , Balch et al., (2019), and elsewhere. Figure 10 shows the same plots of marginal posterior probability assigned to $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for different realizations $x$. The over-confidence appears for any $x$ that’s away from the origin; the discrepancy between the curve and the diagonal line is increasing in $\|x\|^{2}$ to a point, and then remains roughly constant as $\|x\|^{2}\to\infty$.
(a) $x=(0.1,0.3)$
(b) $x=(0.1,3)$
Figure 10: Plots of the marginal posterior probability assigned $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for several different $x$, where $m(\theta)=\|\theta\|^{2}$ is the squared norm.
Appendix E Computational details from Section 5
Here I review some of the relevant details presented in Cella and Martin, (2024) on the choice of $\xi=\xi(x,\alpha)$ in Section 5. In this review I’ll focus on use of a Gaussian variational family, which is motivated by the asymptotic results in Section 4.2, but similar developments can surely be made with other distributional families. The key features of the Gaussian are that its credible sets can be described in closed-form (ellipsoids) and that it’s straightforward to sample on the boundary of these credible sets.
Let $\mathsf{G}_{x}^{\xi}$ denote the Gaussian distribution ${\sf N}_{D}(\hat{\theta}_{x},(J_{x}^{\xi})^{-1})$ with mean equal to the maximum likelihood estimator and covariance matrix the inverse of the $\xi$-adjusted observed Fisher information matrix in (15). Recall that $\xi$ scales the eigenvalues of $J_{x}^{\xi}$, in particular, the eigenvalues are decreasing elementwise in $\xi$. The corresponding Gaussian credible/confidence sets are
$C_{\alpha}^{\xi}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}J_{x}^{\xi}(\theta-\hat{\theta}_{x})\leq\chi_{D,1-\alpha}^{2}\},\quad\alpha\in[0,1].$
The intuition is that $C_{\alpha}^{\xi}(x)$ can be appropriately inflated or deflated by varying $\xi$, and this is relevant because the Gaussian approximation $\mathsf{G}_{x}^{\xi}$ assigns probability at least $1-\alpha$ to the original IM’s $\alpha$-cut $C_{\alpha}(x)$ if $C_{\alpha}^{\xi}(x)\supseteq C_{\alpha}(x)$ or, equivalently, if
$\sup_{\theta\not\in C_{\alpha}^{\xi}(x)}\pi_{x}(\theta)\leq\alpha.$
Since the contour $\pi_{x}$ is itself approximately Gaussian (Martin and Williams, 2025) and the maximum likelihood estimator $\hat{\theta}_{x}$, also the mode of $\pi_{x}$, is in $C_{\alpha}^{\xi}(x)$, the above supremum should be attained on the boundary $\partial C_{\alpha}^{\xi}(x)$. Moreover, since equality in the above display implies a near-perfect match between the IM’s and the posited Gaussian $\alpha$-cuts, a reasonable goal is to find a root to the function
$g_{\alpha}(\xi):=\max_{\theta\in\partial C_{\alpha}^{\xi}(x)}\pi_{x}(\theta)-\alpha.$
Design of an iterative algorithm to find this root requires care, primarily because evaluating $\pi_{x}$ is expensive; so the goal is to evaluate $g_{\alpha}(\xi)$ with as few $\pi_{x}$ evaluations as possible. Cella and Martin, (2024) propose to represent the boundary of $C_{\alpha}^{\xi}(x)$ by $2D$-many vectors
$\vartheta_{s}^{\xi,\pm}:=\hat{\theta}_{x}\pm\bigl{\{}\chi_{D,1-\alpha}^{2}\,\xi_{s}\,/\,\lambda_{s}\}^{1/2}\,e_{s},\quad s=1,\ldots,d,$ (23)
where $(\lambda_{s},e_{s})$ is the eigenvalue–eigenvector pair corresponding to the $s^{\text{th}}$ largest eigenvalue in the spectral decomposition of $J_{x}$. Define the vector function $\hat{g}_{\alpha}$ with components
$\hat{g}_{\alpha,s}(\xi)=\max\{\pi_{x}(\vartheta_{s}^{\xi,+}),\pi_{x}(\vartheta_{s}^{\xi,-})\}-\alpha,\quad s=1,\ldots,D.$ (24)
At least intuitively, negative and positive $\hat{g}_{\alpha,s}(\xi)$ indicate that the $\alpha$-cut $C_{\alpha}^{\xi}(x)$ is too large and too small, respectively in the $e_{s}$-direction. From here, Cella and Martin, (2024) suggest applying a stochastic approximation algorithm—see Robbins and Monro, (1951) and Kushner and Yin, (2003)—to construct a sequence $(\xi^{(t)}:t\geq 1)$ of $D$-vectors that converges to a root of $\hat{g}_{\alpha}$ and, hence, an approximate root of $g_{\alpha}$. For an initial guess $\xi^{(0)}$, the specific sequence is defined as
$\xi_{s}^{(t+1)}=\xi_{s}^{(t)}+w_{t+1}\,\hat{g}_{\alpha,s}(\xi^{(t)}),\quad s=1,\ldots,D,\quad t\geq 0,$
where $(w_{t})$ is a deterministic sequence that satisfies
$\sum_{t=1}^{\infty}w_{t}=\infty\quad\text{and}\quad\sum_{t=1}^{\infty}w_{t}^{2}<\infty.$
These steps are iterated until (practical) convergence is achieved, and the limit is what I called $\xi(x,\alpha)$ in Section 5. The basic steps are outlined in Algorithm 1, but I refer to Cella and Martin, (2024) for further details and discussion.
requires: data $x$, eigen-pairs $(\lambda_{s},e_{s})$, and ability to evaluate $\pi_{x}$;
initialize: $\alpha$-level, guess $\xi^{(0)}$, step size sequence $(w_{t})$, and threshold $\varepsilon>0$;
set: stop = FALSE, $t=0$;
while !stop do
       construct the representative points $\{\vartheta_{s}^{\xi^{(t)},\pm}:s=1,\ldots,D\}$ as in (23);
       evaluate $\hat{g}_{\alpha,s}(\xi_{s}^{(t)})$ for $s=1,\ldots,D$ as in (24);
       update $\xi_{s}^{(t+1)}=\xi_{s}^{(t)}\pm w_{t+1}\,\hat{g}_{\alpha,s}(\xi_{s}^{(t)})$ for $s=1,\ldots,D$;
       if $\max_{s}|\xi_{s}^{(t+1)}-\xi_{s}^{(t)}|<\varepsilon$ then
             $\xi(z,\alpha)=\xi^{(t+1)}$;
             stop = TRUE;
      else
             $t\leftarrow t+1$;
       end if
end while
return $\xi(x,\alpha)$;
Algorithm 1 Determining $\xi(x,\alpha)$—from Cella and Martin, (2024)
Appendix F Additional examples
F.1 Bivariate normal correlation
A challenging one-parameter inference problem is that where $X=(X_{1},\ldots,X_{n})$ are iid, with $X_{i}=(X_{1i},X_{2i})$ a bivariate normal random vector with zero means, unit standard deviations, and correlation $\Theta\in\mathbb{T}=(-1,1)$ to be inferred. Ironically, the known mean and variance case is more difficult—it’s a curved exponential family so the minimal sufficient statistic is not complete and there are various ancillary statistics available to condition on (e.g., Basu, 1964). How this curvature affects asymptotic inference is detailed in Reid, (2003). A likelihood-based possibilistic IM, which is exactly valid for all sample sizes and asymptotically efficient, is easy to construct in theory (Martin, 2024a , Example 2), but naive computation is relatively expensive because there’s no pivot structure and no closed-form expression for the maximum likelihood estimator. This motivates the faster, approximate solution suggested in Martin, (2025) and Section 5 above, based on the variational approximation strategy developed in Cella and Martin, (2024).
A relevant feature of the exact IM solution is that, when $\Theta$ is relatively close the extremes, i.e., $\Theta\approx\pm 1$, the IM’s possibility contour tends to be very asymmetric. Consequently, approximation suggested in Martin, (2025) and Section 5 above, which are based on stitching together symmetric, Gaussian credible intervals, tend to be a bit conservative. Other asymmetric distributions, e.g., skewed normal (e.g., Azzalini, 1985), could potentially replace the Gaussian recommendation made in Martin, (2025), but that’s beyond the scope of the present paper. I’m mentioning this point here because the same conservative behavior will be observed below with the inner probabilistic approximation; this isn’t a shortcoming of the inner probabilistic approximation itself, but of the choice to bound the IM’s asymmetric level sets using symmetric Gaussian credible limits.
For illustration, let $X=(X_{1},\ldots,X_{n})$ denote (a centered and scaled version of) the law school admissions data analyzed in Efron, (1982), where $n=15$ and the maximum likelihood estimator of $\Theta$ is $\hat{\theta}_{x}=0.789$. Figure 11(a) shows a plot of the Bayesian posterior density, based on Jeffreys prior, along with samples from the inner probabilistic approximation. (The latter is based on first reparametrizing in terms of Fisher’s z-transformation, i.e., $\Psi=\tanh^{-1}\Theta$, obtaining the inner probabilistic approximation of $\Psi$, and then transforming back to $\Theta=\tanh\Psi$.) Both distributions are concentrated around the maximum likelihood estimator and have similar shapes, but the Bayes posterior is a little tighter than inner probabilistic approximation.
Is the Bayes posterior distribution too concentrated in the above illustration? If so, then is this aggressive behavior systematic or just a fluke in this one example? To answer these questions, I carried out a brief simulation study, involving 3000 data sets of size $n=10$ drawn from a standard bivariate normal distribution with $\Theta=0.7$. For both methods, and each data set $X$, I evaluate
$U_{X}=1-\bigl{|}2Q_{X}(\Theta)-1\bigr{|},$
where $Q_{X}(\Theta)$ is the cumulative distribution function of the derived posterior distribution, based on the Bayes posterior or the inner probabilistic approximation, evaluated at the true $\Theta=0.7$. If the random variable $U_{X}$ is stochastically no smaller than ${\sf Unif}(0,1)$, then the corresponding equal-tailed posterior credible intervals have frequentist coverage probability at or above the nominal level; if $U_{X}$ is stochastically smaller than ${\sf Unif}(0,1)$, the credible intervals undercover. Figure 11(b) plots the cumulative distribution function of $U_{X}$ for the two methods and we find that the Jeffreys-prior Bayes solution and the inner probabilistic approximation solution yield $U_{X}$ that are stochastically smaller and larger than ${\sf Unif}(0,1)$, respectively. This means that the former’s credible intervals generally undercover while the latter’s generally overcover; therefore, the Bayes posterior density displayed in Figure 11(a) might indeed be “too concentrated.”
(a) Posterior results in Efron’s example
(b) Cumulative distribution function of $U_{X}$
Figure 11: Comparison of Jeffreys-prior Bayes and the possibilistic IM’s inner probabilistic approximation. In Panel (a), the solid line is the Bayes posterior density and the bars are based on samples from the inner probabilistic approximation. In Panel (b), the black and gray lines correspond to the Bayes and inner probabilistic approximation solutions, respectively. Definition and explanation of $U_{X}$ is given in the text.
F.2 Gamma shape and scale
Inference on the parameters of a gamma model is a classical problem. When the shape parameter is unknown, this model lacks the group transformation structure that makes specification of default priors having suitable reliability properties non-trivial. The go-to no-prior Bayes solution invokes Jeffreys prior on the shape–scale parameter pair. Alternatively, a possibilistic IM solution is readily available and doesn’t require any prior specification and has reliability guarantees. This IM solution is somewhat expensive to compute, however, at least if a naive approach is taken. Martin, (2025) showed how this naive approach can be drastically improved and it’s on this more sophisticated/efficient approach that my solution is based. Indeed, here I’m proposing the inner probabilistic approximation of that possibilistic IM, and the sampling from that “posterior distribution” is based on the strategy put forward in Martin, (2025).
As a first illustration, reconsider the data analyzed in Fraser et al., (1997) on the survival time (in weeks) for $n=20$ rats exposed to a certain amount of radiation. Assuming a gamma model with unknown shape and scale parameters, the posterior distributions for the two no-prior Bayes solutions—Jeffreys prior Bayes and the IM’s inner probabilistic approximation—are shown in Figure 12, both in the original and log-transformed parametrization. The two posteriors look generally similar and, after log-transformation, they both have Gaussian-like elliptical shapes. The only notable difference is that the Bayes solution is a bit more tightly concentrated than the IM-based solution.
(a) shape–scale parametrization
(b) log(shape)–log(scale) parametrization
Figure 12: Black points are samples from the Jeffreys-prior Bayes posterior distribution in the gamma model; gray points are samples from the IM’s inner probabilistic approximation.
Like in the bivariate normal correlation example above, it’s natural to ask if the Bayes solution is actually too tightly concentrated. To assess this, I’ll do a brief simulation study. Here I draw 2500 samples of size $n=20$, with true shape and scale parameters set to be $\Theta_{1}=7$ and $\Theta_{2}=3$, respectively. Similar to the bivariate normal example in the previous section, for both no-prior Bayes solutions, define the quadratic form of the form
$\kappa_{x}(\theta)=(\log\theta-\hat{\mu}_{x})^{\top}\hat{\sigma}_{x}^{-1}(\log\theta-\hat{\mu}_{x}),$
where $\hat{\mu}_{x}$ and $\hat{\sigma}_{x}$ are estimates of the posterior mean and posterior covariance matrix of the log-transformed shape and scale parameters. Then I let
$Q_{x}(\theta)=\mathsf{Q}_{x}\{\kappa_{x}(\widetilde{\Theta})\leq\kappa_{x}(\theta)\}$
denote the distribution function of the quadratic form relative to the posterior distribution and, finally, I define
$U_{X}=1-\bigl{|}2Q_{X}(\Theta)-1\bigr{|},$
where $\Theta=(7,3)$ is the true shape and scale parameters. Figure 13 plots the distribution functions of $U_{X}$ for each of the two methods. Being above the diagonal line indicates that the elliptical credible sets derived from the method’s posterior samples have frequentist coverage probability below the nominal level; similarly, being on or below the line indicates that the credible sets have frequentist coverage probability at or above the nominal level. In this case, the Jeffreys-prior Bayes solution shows severe undercoverage while the IM’s inner probabilistic approximation shows signs of very slight undercoverage for 90 and 95% credible regions. This suggests that the Jeffreys-prior Bayes solutions displayed in Figure 12 might indeed be too tightly concentrated.
Figure 13: Comparison of Jeffreys-prior Bayes and the possibilistic IM’s inner probabilistic approximation, in particular, distribution functions of $U_{X}$ corresponding to the Bayes (black) and IM inner probabilistic approximation (gray).