### v3

# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

Abstract
When prior information is lacking, the go-to strategy for probabilistic inference is to combine a “default prior” and the likelihood via Bayes’s theorem. Objective Bayes, (generalized) fiducial inference, etc. fall under this umbrella. This construction is natural, but the corresponding posterior distributions generally only offer limited, approximately valid uncertainty quantification. The present paper takes a reimagined approach that yields posterior distributions with stronger reliability properties. The proposed construction starts with an inferential model (IM), one that takes the mathematical form of a data-driven possibility measure and features exactly valid uncertainty quantification, and then returns a so-called inner probabilistic approximation thereof. This inner probabilistic approximation inherits many of the original IM’s desirable properties, including credible sets with exact coverage and asymptotic efficiency. The approximation also agrees with the familiar Bayes/fiducial solution in applications where the model has a group invariance structure. A Monte Carlo method for evaluating the probabilistic approximation is presented, along with numerical illustrations.
Keywords and phrases: Confidence distribution; credal set; fiducial inference; p-value; possibility theory; relative likelihood; validity.
1 Introduction
The program in which science advances by formulating what’s currently known, however vague or incomplete that might be, as prior beliefs and then applying a normative procedure to update those prior beliefs in light of new data is natural and deeply rooted in all of us who’ve studied probability and statistics. Despite its appeal, this program is challenging to implement for several reasons, not the least of which is that genuine prior information is rarely available, hence no particular formulation of prior beliefs is warranted. Brad Efron, in his presentation at the 2016 Joint Statistical Meetings in Chicago, said that “scientists like to work on new problems,” meaning that there’s often no history or experience to draw from in order to build a meaningful prior distribution. For similar reasons, Efron, (2013) stated more formally that “…the most important unresolved problem in statistical inference is the use of Bayes theorem in the absence of prior information.” The present paper offers new insights, theory, and methodology with the ambitious goal of resolving this most important unresolved problem.
Fisher, (1933); Fisher, 1935a ; Fisher, 1935b and his fiducial argument was the first proposed solution to the above problem, “a bold attempt to make the Bayesian omelet without breaking the Bayesian egg” (Savage, 1961); see Zabell, (1992) and Dawid, (2024) for excellent surveys. The consensus is that Fisher’s solution failed, but even his “biggest blunder” (Efron, 1998) had a major impact, inspiring fundamental advances such as confidence limits (Neyman, 1941), imprecise probability (Dempster, 1966), and new proposed solutions, including generalized fiducial inference (Hannig et al., 2016; Murph et al., 2024), confidence distributions (Xie and Singh, 2013; Thornton and Xie, 2020; Cox, 2006; Schweder and Hjort, 2016), and objective/non-informative Bayes (Jeffreys, 1946; Berger, 2006; Berger et al., 2024). In what follows, to keep terminology simple, I’ll refer to all of these approaches as no-prior Bayes solutions.
A difficulty with Bayesian-like probabilistic inference in this no-prior setting concerns the interpretation of the posterior probabilities themselves. When genuine prior information is available, then the Bayesian posterior probabilities are the unique coherent updates of prior beliefs given the observed data. But when there’s no prior beliefs to update, and a default prior is used in its place, of course the aforementioned “updating” interpretation disappears and, consequently, it’s unclear if the corresponding posterior distribution is meaningful at all. Indeed, “[Bayes’s formula] does not create real probabilities from hypothetical probabilities” (Fraser, 2014) and, more forcefully,
…any serious mathematician would surely ask how you could use [Bayes’s theorem] with one premise missing by making up an ingredient and thinking that the conclusions of the [theorem] were still available (Fraser, 2011b ).
Fortunately, coherent updating of prior beliefs isn’t necessary for a posterior to have belief-forming inferential weight. But making this case requires, first, defining what it takes for a framework to have belief-forming inferential weight and, second, demonstrating that a proposed framework meets this requirement. To me, the only way forward is to demonstrate that the posterior probabilities are reliable, i.e., it’s a demonstrably rare event that the posterior assigns high (resp. low) probabilities to assertions about the unknown that are false (resp. true). The advantage of such a requirement is that the belief-forming inferential weight is inherited from Fisher’s inductive logic: if, e.g., assigning low probability to a true assertion is a rare event and, in the application at hand, the assertion is assigned low probability, then it’s safe to infer that the assertion is false because rare events effectively don’t happen. No-prior Bayes solutions often have some form of reliability, but the false confidence theorem (Balch et al., 2019; Martin, 2019) says that none have the strong reliability property just described. Since no-prior Bayes solutions can’t offer reliable belief assignments, there’s motivation to look beyond these existing approaches.
Inferential models (IMs) (e.g., Martin and Liu, 2013, 2015; Martin, 2025b ) are an alternative to the aforementioned probabilistic approaches, whose output is a data-dependent possibility measure that offers provably reliable uncertainty quantification about unknown parameters. Further details about possibility measures and a specific IM construction are given in Section 2. Importantly, my proposed shift from probabilistic to possibilistic uncertainty quantification facilitates strong reliability and, among other things, this strong reliability implies the usual frequentist error rate control without giving up on fully conditional Bayesian-like reasoning.
Despite the benefits of possibilistic IMs compared to probabilistic no-prior Bayes solutions, I’m under no delusion that statisticians will abandon probabilism in favor of possibilism in the near future. But that doesn’t mean IMs have to wait for the distant future to make a contribution. The “reIMagined” approach I’m proposing here starts with the provably reliable, likleihood-based possibilistic IM and then reads off and returns an inner probabilistic approximation as a novel no-prior Bayes solution. As I’ll demonstrate below, the proposed inner probabilistic approximation inherits some—but necessarily not all—of the original possibilistic IM’s strong reliability properties: in particular, the corresponding highest posterior density credible sets for the full parameter are confidence sets. The motivation for this new approach is that there’s only so much that can be achieved by constructing default priors and checking if the corresponding posterior distributions are reliable. My suggestion here is to prioritize the reliability properties and directly construct a data-dependent probability that has those desired properties.
The remainder of the paper is organized as follows. Some background on possibility theory and IMs is presented in Section 2. A characterization of the IM’s credal set is presented in Section 3, which is then used to define the corresponding inner probabilistic approximation. Various properties of the proposed inner probabilistic approximation are developed in Section 4, including an agreement with the familiar no-prior Bayes solutions in invariant models and a version of the celebrated Bernstein–von Mises theorem that establishes its asymptotic efficiency. A strategy for (approximately) computing the IM’s inner probabilistic approximation is presented in Section 5 and a new, valid, and efficient solution to the technically challenging and practically relevant Behrens–Fisher problem is presented in Section 6. Some concluding remarks are given in Section 7 and the appendices contain some additional technical details and illustrations.
2 Background
2.1 Possibility theory
Possibility measures (e.g., Dubois and Prade, 1988) are among the simplest imprecise probability models, closely linked to fuzzy set theory (e.g., Zadeh, 1978) and Dempster–Shafer theory (e.g., Shafer, 1976, 1987). Applications in statistics are described in Dubois, (2006) and Dubois and Denœux, (2010); see, also, Section 2.2.
The mathematical differences between probability and possibility theory can be succinctly summarized as follows: optimization is to possibility theory what integration is to probability theory. That is, a possibility measure $\mathsf{\Pi}$ defined on a space $\mathbb{Z}$ is determined by a function $\pi:\mathbb{Z}\to[0,1]$ with the property that $\sup_{z\in\mathbb{Z}}\pi(z)=1$. This function is called the possibility contour and the supremum-equals-1 property is a normalization condition analogous to the familiar integral-equals-1 property of probability densities. Then the possibility measure is determined by optimizing its contour, i.e., $\mathsf{\Pi}(A)=\sup_{z\in A}\pi(z)$, for any $A\subseteq\mathbb{Z}$, just like a probability measure is determined by integrating its density.
The different calculus has a number of implications. Of particular relevance for the developments in this paper is that the aforementioned supremum-equals-1 normalization condition ensures that $\mathsf{\Pi}$ is a coherent upper probability (e.g., De Cooman, 1997; De Cooman and Aeyels, 1999) in the spirit of Walley, (1991) and others. Among other things, this means that $\mathsf{\Pi}$ determines a non-empty (closed and convex) set of probabilities it dominates:
$\mathscr{C}(\mathsf{\Pi})=\{\mathsf{Q}\in\mathcal{P}(\mathbb{Z}):\mathsf{Q}\preceq\mathsf{\Pi}\},$ (1)
where $\mathcal{P}(\mathbb{Z})$ is the set of probabilities supported on the Borel $\sigma$-algebra of measurable subsets of $\mathbb{Z}$ and $\mathsf{Q}\preceq\mathsf{\Pi}$” means $\mathsf{Q}(H)\leq\mathsf{\Pi}(H)$ for all measurable HH.” The set $\mathscr{C}(\mathsf{\Pi})$ is called the credal set corresponding to $\mathsf{\Pi}$ and all (coherent) upper probabilities have one. Aside from being relatively simple, a key advantage to possibilistic uncertainty quantification is that the associated credal set has a statistically oriented characterization, which is important for the statistical developments below.
Critical to the developments in the present paper, and closely tied to the above credal sets, is a notion of approximating a probability by a possibility and vice versa. I’ll focus here on the former, and leave the latter for Section 3. Suppose that $\mathsf{Q}$ is a probability on $\mathbb{T}$ with density/mass function $q$. The probability-to-possibility transform (e.g., Dubois et al., 2004; Hose, 2022) applied to $\mathsf{Q}$ yields a possibility measure $\mathsf{\Pi}$ determined by the contour
$\pi(\theta)=\mathsf{Q}\{q(\Theta)\leq q(\theta)\},\quad\theta\in\mathbb{T}.$
This $\mathsf{\Pi}$ has two key properties: first, $\mathscr{C}(\mathsf{\Pi})$ contains $\mathsf{Q}$, so that $\mathsf{\Pi}\succeq\mathsf{Q}$; second, $\mathsf{\Pi}$ is the “least imprecise” possibilistic approximation of $\mathsf{Q}$ in the sense that it has the smallest credal set among other suitable possibilistic approximations. More formally, $\mathscr{C}(\mathsf{\Pi})=\bigcap\mathscr{C}(\mathsf{\Omega})$, where the intersection is over all those possibility measures $\mathsf{\Omega}$ that satisfy $\mathsf{\Omega}\succeq\mathsf{Q}$ and whose associated contour $\omega$ is comonotone [^2] Comonotonicity here means that the contour $\omega(\theta)$ is an increasing function of $q(\theta)$. This constraint is motivated by a basic and fundamental principle in this possibility-theory context: “what is probable must be plausible” (Dubois and Prade, 1988, p. 121). with $q$. So, $\mathsf{\Pi}$ is often called an outer possibilistic approximation of $\mathsf{Q}$ (e.g., Dubois and Prade, 1990).
2.2 Possibilistic IMs
An inferential model (IM) is a mapping from data, model, etc. to an imprecise-probabilistic quantification of uncertainty about relevant unknowns. Critical to this approach and perspective is that the IM’s uncertainty quantification be reliable, as a function of data, in a sense to be described below. It’s this insistence on reliability that implies the IM output must take on the form of an imprecise probability. The first IM developments were built up using random sets and described using belief function terminology. More recent developments directly appeal to possibility-theoretic tools and reasoning in the IM construction and interpretation, and I refer to these as possibilistic IMs; see Martin, 2025b for more details. This paper focuses on this latter brand of IMs.
Consider a model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ consisting of probability distributions supported on a sample space $\mathbb{X}$, with parameter space $\mathbb{T}$. Suppose that the observable data $X$, taking values in $\mathbb{X}$, is a sample from the distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the unknown/uncertain “true value.” Prior information about $\Theta$ is assumed to be vacuous, but see Martin, 2022b for generalizations. The model and observed data $X=x$ determine a relative likelihood function
$R(x,\theta)=\frac{p_{\theta}(x)}{\sup_{\vartheta}p_{\vartheta}(x)},\quad\theta\in\mathbb{T},$
where $p_{\theta}$ is the density of $\mathsf{P}_{\theta}$. I will assume throughout that the denominator is finite for almost all $x$.
The relative likelihood itself defines a possibility contour, i.e., a non-negative function such that $\sup_{\theta}R(x,\theta)=1$ for almost all $x$, which can be used for data-driven uncertainty quantification about $\Theta$. This has been extensively studied in (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014), and it has a number of desirable properties. What it lacks, however, is a formal calibration property justifying that the “possibilities” assigned to hypotheses about $\Theta$ have belief-forming inferential weight.
Fortunately, this reliability-focused calibration is easy to achieve, by “validifying” (Martin, 2022a ; Martin, 2022b ) the relative likelihood. That is, the likelihood-based possibilistic IM construction proceeds by applying a version of the probability-to-possibility transform to the relative likelihood, returning contour function:
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\bigl\{R(X,\theta)\leq R(x,\theta)\bigr\},\quad\theta\in\mathbb{T}.$ (2)
The corresponding possibility measure $\mathsf{\Pi}_{x}$ is defined via optimization just like I explained above:
$\mathsf{\Pi}_{x}(H)=\sup_{\theta\in H}\pi_{x}(\theta),\quad H\subseteq\mathbb{T}.$ (3)
An illustration of what a typical contour $\pi_{x}$ looks like is shown in Figure 1. Some further details about this simple example are presented at the end of Section 3 below.
Figure 1: Illustration of a typical possibility contour $\theta\mapsto\pi_{x}(\theta)$ and how the possibility measure $\mathsf{\Pi}_{x}$ is determined from it. Here, the hypothesis $H$ is the interval $[3,5]$ and the maximum value $\mathsf{\Pi}_{x}(H)$ on contour is highlighted. Horizontal dashed line at $\alpha=0.1$ determines the 90% confidence interval $C_{0.1}(x)$ in (5).
As in Section 1, an essential feature of this possibilistic IM is its validity property:
$\sup_{\theta\in\mathbb{T}}\mathsf{P}_{\theta}\bigl\{\pi_{X}(\theta)\leq\alpha\bigr\}\leq\alpha,\quad\text{$\alpha\in[0,1]$}.$ (4)
In words, (4) means that it’s a rare event for the IM to assign small possibility $\pi_{X}(\Theta)$ to the true $\Theta$, which, of course, was important to the inductive logic explained in Section 1. This has a number of important practical consequences as well. First, (4) immediately implies that the upper $\alpha$ level set of the possibility contour,
$C_{\alpha}(x)=\{\theta\in\mathbb{T}:\pi_{x}(\theta)>\alpha\},\quad\alpha\in[0,1],$ (5)
is a $100(1-\alpha)$% confidence set, i.e.,
$\sup_{\theta\in\mathbb{T}}\mathsf{P}_{\theta}\bigl\{C_{\alpha}(X)\not\ni\theta\bigr\}\leq\alpha.$
Figure 1 gives a depiction of $C_{\alpha}(x)$, with $\alpha=0.1$. Second, from (3) and (4),
$\sup_{\theta\in H}\mathsf{P}_{\theta}\bigl\{\mathsf{\Pi}_{X}(H)\leq\alpha\bigr\}\leq\alpha,\quad\text{$\alpha\in[0,1]$, $H\subseteq\mathbb{T}$}.$ (6)
In words, a valid IM assigns possibility $\leq\alpha$ to true hypotheses at rate $\leq\alpha$ as a function of data $Z$. This gives the IM its “inferential weight”—(6) implies that $\mathsf{\Pi}_{x}(H)$ is not expected to be small when $H$ is true, so one is inclined to doubt the truthfulness of a hypothesis $H$ if $\mathsf{\Pi}_{x}(H)$ is small, and the strength of this inclination is determined by the magnitude of $\mathsf{\Pi}_{x}(H)$. This of course can be turned into a formal test procedure with frequentist error rate guarantees: by (6), the test “reject $H$ if $\mathsf{\Pi}_{x}(H)\leq\alpha$” controls the Type I error probability at level $\alpha$. Third, the above property ensures that the possibilistic IM is safe from false confidence (Balch et al., 2019; Martin, 2019), unlike other no-prior Bayes solutions. Further insights about IMs are presented in Appendix A in the supplementary materials; see, also, Martin, 2025b .
The IM output is a coherent imprecise probability and, consequently, there is an associated (non-empty) credal set, $\mathscr{C}(\mathsf{\Pi}_{x})$ as in (1). The members of $\mathscr{C}(\mathsf{\Pi}_{x})$, which I’ll denote by $\mathsf{Q}_{x}$, need not correspond to Bayesian posterior distributions under any prior. Fortunately, an interpretation can be given to the members of $\mathscr{C}(\mathsf{\Pi}_{x})$, and this is key to my developments in Section 3.
I’ll end this background section with three technical remarks about the IM scope and contribution. First, the reader certainly recognizes the contour function $\pi_{x}(\theta)$ in (2) as a p-value function corresponding to the likelihood ratio statistic. Aside from p-value function, this has appeared in the literature under many other names, e.g., preference functions (Spjøtvoll, 1983), significance functions (Fraser, 1991), and confidence curves (Blaker and Spjøtvoll, 2000; Birnbaum, 1961; Schweder and Hjort, 2002, 2016; Xie and Singh, 2013). A key point distinguishing IMs from this literature is that it embraces the inherent imprecise and leverages relevant imprecise-probabilistic results. That is, an IM determines a coherent and fully conditional imprecise-probabilistic quantification of uncertainty for inference and more (Martin, 2025b ). Even the more Bayesian-leaning developments along these lines (e.g., Pereira and Stern, 2022; Grünwald, 2023; Cortinovis and Caron, 2024) don’t take full advantage of what imprecise probability has to offer.
Second, with the big-picture assessment in the previous paragraph, one might compare IMs to other imprecise-probabilistic solutions like that in Walley, (1991). A critical observation is that, when prior information about $\Theta$ is vacuous, as assumed here, Walley’s generalized Bayes posterior is also vacuous, i.e., no learning is achieved. Since Bayesian reasoning isn’t fully satisfactory in such cases, new ideas, like those reviewed above, are needed to achieve inference that’s both valid and efficient.
Finally, with the connection to p-values above, it should be clear that other rankings besides the relative likelihood can be used in the construction. My reason for focusing on the relative likelihood here is two-fold: first, this is a principled choice in the case where the full parameter $\Theta$ is of interest (Martin, 2022b , Sec. 4); second, the focus of the present paper is on connections to Bayesian inference, and some direct dependence on the model’s likelihood function is surely needed to make that connection. But there are cases when it’s prudent to replace the relative likelihood with something else, e.g., when nuisance parameters are present as in Section 6.
3 Inner probabilistic approximations
3.1 Intuition
Section 2.1 discussed approximating a given probability distribution by a possibility measure, the workhorse being the probability-to-possibility transform. The goal of the present section is to reverse this process: to approximate the possibilistic IM by a suitable probability distribution. Before giving a general description, I’ll offer some intuition using a simple example where all the calculations can be done explicitly.
Let $X\sim\mathsf{P}_{\Theta}={\sf N}(\Theta,1)$. Given $X=x$, the standard no-prior Bayes solution, corresponding to the flat, Jeffreys prior, returns $\mathsf{Q}_{x}={\sf N}(x,1)$ as the posterior distribution for $\Theta$. My proposal of approximating the possibilistic IM by a probability definitely isn’t effective if it doesn’t yield the standard no-prior Bayes solution in this simple case. For the IM construction, the relative likelihood is $R(x,\theta)=\exp\{-(x-\theta)^{2}/2\}$ and the contour function is
$\pi_{x}(\theta)$ $=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}$
$=\mathsf{P}_{\theta}\{e^{-(X-\theta)^{2}/2}\leq e^{-(x-\theta)^{2}/2}\}$
$=\mathsf{P}\{{\sf ChiSq}(1)\geq(x-\theta)^{2}\}$
$=1-F_{1}\bigl((x-\theta)^{2}\bigr),$
where $F_{\nu}$ is the ${\sf ChiSq}(\nu)$ distribution function. How are $\mathsf{Q}_{x}$ and the corresponding $\mathsf{\Pi}_{x}$ related? One way to see it, following Section 2.1, is to apply the probability-to-possibility transform to $\mathsf{Q}_{x}$ and see what comes out. In this case, with $q_{x}$ the ${\sf N}(x,1)$ density function
$\text{prob-to-poss}_{\mathsf{Q}_{x}}(\theta)$ $:=\mathsf{Q}_{x}\{q_{x}(\Theta)\leq q_{x}(\theta)\}$
$=\mathsf{Q}_{x}\{e^{-(\Theta-x)^{2}/2}\leq e^{-(\theta-x)^{2}/2}\}$
$=\mathsf{P}\{{\sf ChiSq}(1)\geq(\theta-x)^{2}\}$
$=\pi_{x}(\theta).$
That is, the outer possibilistic approximation of $\mathsf{Q}_{x}$, obtained via the probability-to-possibility transform, is precisely the possibilistic IM $\mathsf{\Pi}_{x}$. This is an outer approximation because, roughly, it’s the “least imprecise” possibility measure among those that $\preceq$-dominate $\mathsf{Q}_{x}$. It’s only natural, then, to turn this relationship around and say that $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$ in the sense of being the “maximally diffuse” probability distribution that’s $\preceq$-dominated by $\mathsf{\Pi}_{x}$. These details will be made more precise in Section 3.2 below.
As a preview, there are at least two equivalent ways to extract a probability from a possibility: one based on bounding tail probabilities and another based on level set matching. Here I’ll focus on the latter approach, formalized in Section 3.2. Since $\pi_{x}(\theta)$ is most naturally interpreted as the “plausibility” of $\Theta=\theta$ given $x$, and since posterior densities are (informally) interpreted in the same way, it makes sense that the density $q_{x}$ of $\mathsf{Q}_{x}$ would have the same level sets as $\pi_{x}$. Akin to slice sampling (e.g., Neal, 2003), a probability $\mathsf{Q}_{x}$ is determined by:
  • first sampling the level set;
  • then sampling a point on the chosen level set.
In the present case, the level sets are the confidence intervals $C_{\alpha}(x)=\{\theta:\pi_{x}(\theta)>\alpha\}$, indexed by $\alpha\in[0,1]$. Then sampling the level sets is equivalent to sampling the level itself, say $A$, from a distribution supported on $[0,1]$. If $\mathsf{Q}_{x}$ represents the distribution of $\Theta$ under the two-stage sampling scheme above, then $\Theta\in C_{\alpha}(x)$ if and only if $A>\alpha$. Then dominance $\mathsf{Q}_{x}\preceq\mathsf{\Pi}_{x}$ implies
$\mathsf{Q}_{x}(A>\alpha)$ $\equiv\mathsf{Q}_{x}\{C_{\alpha}(x)\}$
$=1-\mathsf{Q}_{x}\{C_{\alpha}(x)^{c}\}$
$\geq 1-\mathsf{\Pi}_{x}\{C_{\alpha}(x)^{c}\}$
$=1-\alpha.$
That is, dominance itself $A$ is stochastically no smaller than ${\sf Unif}(0,1)$. At least intuitively, the “maximally diffuse” probability distribution $\mathsf{Q}_{x}^{\star}$ dominated by $\mathsf{\Pi}_{x}$ corresponds to taking $A\sim{\sf Unif}(0,1)$ and, given the level $A=\alpha$, sampling $\Theta$ on $C_{\alpha}(x)$ with maximal variance, i.e., uniformly on the endpoints of $C_{\alpha}(x)$. Those endpoints are $x\pm z_{\alpha/2}$, where $z_{\beta}$ is the upper-$\beta$ quantile of ${\sf N}(0,1)$, and then the distribution function of $\mathsf{Q}_{x}^{\star}$ is
[…]
As a function of $A\sim{\sf Unif}(0,1)$, the random variable $z_{A/2}$ has the same distribution as $|Z|$ with $Z\sim{\sf N}(0,1)$. So,
$\mathsf{Q}_{x}^{\star}\bigl((-\infty,\theta]\bigr)=\tfrac{1}{2}\,\mathsf{P}(|Z|\geq x-\theta)+\tfrac{1}{2}\,\mathsf{P}(|Z|\leq\theta-x).$
Using some basic properties of the standard normal distribution, the right-hand side simplifies, leading to
$\mathsf{Q}_{x}^{\star}\bigl((-\infty,\theta]\bigr)=\mathsf{P}(Z\leq\theta-x),$
and, therefore, this “maximally diffuse” probability distribution $\mathsf{Q}_{x}^{\star}$ dominated by $\mathsf{\Pi}_{x}$ is just the standard no-prior Bayes solution, i.e., $\mathsf{Q}_{x}^{\star}={\sf N}(x,1)$.
3.2 Characterization
The goal of the present section is to formalize and generalize the intuition developed above. Start with a well-known characterization of the credal set for general possibility measures (e.g., Couso et al., 2001; Destercke and Dubois, 2014) which, applied to the present case, says
$\begin{split}\mathsf{Q}_{x}&\in\mathscr{C}(\mathsf{\Pi}_{x})\iff\mathsf{Q}_{x}\{C_{\alpha}(x)\}\geq 1-\alpha\;\;\text{for all $\alpha\in[0,1]$},\end{split}$ (7)
where $C_{\alpha}(x)$ is as defined in (5) with $\pi_{x}$ the contour corresponding to $\mathsf{\Pi}_{x}$. The IM’s validity implies that $C_{\alpha}(x)$ is a $100(1-\alpha)$% confidence set, so, since the elements $\mathsf{Q}_{x}$ of $\mathscr{C}(\mathsf{\Pi}_{x})$ assign probability at least $1-\alpha$ to the $100(1-\alpha)$% confidence set $C_{\alpha}(x)$, there’s good reason to call the elements of $\mathscr{C}(\mathsf{\Pi}_{x})$ confidence distributions. This definition of confidence distributions generalizes those commonly found in the statistics literature (Xie and Singh, 2013; Schweder and Hjort, 2013; Cunen et al., 2020) that—for reasons discussed in Section 4.3—focus mainly on scalar parameter cases.
Refinements to the characterization in (7) can be given. For example, Wasserman, 1990b (, Theorem 2.1) characterizes the credal set associated with a belief function as a collection of suitable mixture distributions. The result below gives a similar mixture characterization for possibility measures, including for the IM output $\mathsf{\Pi}_{x}$.
Theorem 1 (Martin, 2025a ).
Let $\pi_{x}(\theta)$ be the contour associated with the possibility measure $\mathsf{\Pi}_{x}$, and set $C_{\alpha}(x)=\{\theta:\pi_{x}(\theta)>\alpha\}$ as in (5). Then $\mathsf{Q}_{x}\in\mathscr{C}(\mathsf{\Pi}_{x})$ if and only if it can be represented as
$\mathsf{Q}_{x}(\cdot)=\int_{0}^{1}\mathsf{K}_{x}^{\alpha}(\cdot)\,\mathsf{M}_{x}(d\alpha),$ (8)
for some kernel $\mathsf{K}_{x}^{\alpha}$, fully supported on $C_{\alpha}(x)$, so that $\mathsf{K}_{x}^{\alpha}\{C_{\alpha}(x)\}=1$ for each $\alpha\in[0,1]$, and some probability measure $\mathsf{M}_{x}$ on $[0,1]$ whose distribution function is upper-bounded by that of ${\sf Unif}(0,1)$.
The mixture representation (8) of the credal set elements implies that a draw $\Theta$ from $\mathsf{Q}_{x}$ can be obtained via a two-step sampling procedure: draw $A\sim\mathsf{M}_{x}$ and, given $A=\alpha$, draw $\Theta\sim\mathsf{K}_{x}^{\alpha}$. The reader will recognize this two-step procedure from the discussion in Section 3.1. There are, of course, many different choices of $\mathsf{K}_{x}^{\alpha}$ and $\mathsf{M}_{x}$ that can be made, and it’s this ambiguity that makes the credal set a genuine, non-singleton set.
A relevant question is if there exists members $\mathsf{Q}_{x}^{\star}$ of $\mathscr{C}(\mathsf{\Pi}_{x})$ such that equality on the right-hand side of (7) is achieved for each $\alpha$, i.e.,
$\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}=1-\alpha,\quad\text{for all $\alpha\in[0,1]$}.$ (9)
This formalizes the notion of “maximally diffuse” in Section 3.1. The point is that each element $\mathsf{Q}_{x}$ of the credal set is a confidence distribution in the sense that it assigns $\mathsf{Q}_{x}$-probability at least $1-\alpha$ to $C_{\alpha}(x)$. Then “maximally diffuse” simply means a confidence distribution that’s not conservative, one whose probability exactly matches the confidence level. As explained in Section 3.1 above, I’ll refer to those $\mathsf{Q}_{x}^{\star}$ for which (9) holds as inner probabilistic approximations of the possibilistic IM $\mathsf{\Pi}_{x}$.
Setting aside the question about existence of an inner probabilistic approximation for the moment, note that the mixture formulation in (8) implies that the highest posterior density credible sets exactly agree with $\{C_{\alpha}(x):\alpha\in[0,1]\}$ in (5). Since validity of the IM implies that the latter are exact confidence sets, then the former are too. That is, inner probabilistic approximations exactly—not just asymptotically—achieve the “probability matching” property (e.g., Staicu and Reid, 2008; Datta and Ghosh, 1995) desired of no-prior Bayes solutions.
Theorem 2.
If $\mathsf{Q}_{x}^{\star}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$, then its associated highest posterior density credible sets are confidence sets.
There are cases in which attainment of the equality in (9) is out of reach. For instance, suppose that $\theta\mapsto\pi_{x}(\theta)$ is bounded away from 0 for some $x$. This is rare in applications, but not impossible. It may happen that the data is not particularly informative about a feature $\Phi=m(\Theta)$ of $\Theta$, and then the corresponding marginal IM (see Section 4.3) has a non-vanishing contour—that is, the data is insufficiently informative to exclude any values of $\Phi$ with high degree of confidence.
Some clues about what it takes to satisfy (9) were already presented in Section 3.1. Indeed, to get a “maximally diffuse” member of the credal set, I suggested taking the marginal distribution $\mathsf{M}_{x}$ to be ${\sf Unif}(0,1)$ and taking the kernel $\mathsf{K}_{x}^{\alpha}$ to be fully supported on the boundary, $\partial C_{\alpha}(x)$, of the level set $C_{\alpha}(x)$. To justify this, note that, for each $\alpha\in[0,1]$, a $\mathsf{Q}_{x}$ of the form (8) satisfies
$\mathsf{Q}_{x}\{C_{\alpha}(x)\}$ $=\int_{0}^{1}\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}\,\mathsf{M}_{x}(d\beta)$
$=\int_{0}^{\alpha}\underbrace{\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}}_{\geq 0}\,\mathsf{M}_{x}(d\beta)+\int_{\alpha}^{1}\underbrace{\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}}_{=1}\,\mathsf{M}_{x}(d\beta)$
$\geq\mathsf{M}_{x}([\alpha,1])$
$\geq 1-\alpha.$
Then it’s clear that equality in (9) is achieved by taking $\mathsf{M}_{x}={\sf Unif}(0,1)$ and $\mathsf{K}_{x}^{\alpha}$ fully supported on $\partial C_{\alpha}(x)$, so that $\mathsf{K}_{x}^{\beta}\{C_{\alpha}(x)\}=0$ for all $\beta<\alpha$.
How the kernel allocates its mass to $\partial C_{\alpha}(x)$ isn’t determined by (9), hence the inner probabilistic approximation of $\mathsf{\Pi}_{x}$ isn’t unique. I’ll say more about the role played by the kernel later and in Appendix B. Here I present my recommended choice of inner probabilistic approximation. Note that the sets $C_{\alpha}(x)$ are typically bounded; in fact, when the sample size is moderate to large, they’re approximately ellipsoids (Section 4.2). Then choosing the kernel $\mathsf{K}_{x}^{\alpha}$ to be a uniform distribution on $\partial C_{\alpha}(x)$, for each $\alpha\in(0,1)$, is quite natural, and consistent with both the maximum entropy (e.g., Cover and Thomas, 2006; Bernardo and Smith, 1994) and indifference principles (e.g., Keynes, 1921; Jaynes, 2003). It’ll also be seen in Section 4.2 that uniformity is needed for asymptotic efficiency of the inner probabilistic approximation.
To summarize, my proposed inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\mathsf{\Pi}_{x}$ is a uniform mixture of uniform distributions. More specifically, a sample $(\Theta\mid x)\sim\mathsf{Q}_{x}^{\star}$ can be drawn via the following two-step procedure:
  1. 1.
    sample $(A\mid x)\sim{\sf Unif}(0,1)$, and
  2. 2.
    sample $(\Theta\mid x,A=\alpha)\sim\mathsf{K}_{x}^{\alpha}:={\sf Unif}\{\partial C_{\alpha}(x)\}$.
Step 2 is far simpler to state than it is to carry out, and I’ll discuss how this computational challenge can be addressed—at least approximately—in Section 5.
3.3 Illustration
Consider a gamma model $\mathsf{P}_{\theta}={\sf Gamma}(n,\theta)$ for a single observation $X$, with $n$ a fixed constant; the more practical case of $n$ iid exponential samples where $X$ is the sum, the complete sufficient statistic, reduces to this one. Then the likelihood-based IM has contour
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\bigl\{(X/\theta)^{n}e^{-X/\theta}\leq(x/\theta)^{n}e^{-x/\theta}\bigr\},\quad\theta>0.$
The right-hand side has no simple closed-form expression but, since $X/\theta$ is a pivot under $\mathsf{P}_{\theta}$, the computations are straightforward using Monte Carlo. The contour shown in Figure 1 is actually the one above, with $n=7$ and $x=14$. Clearly, the level sets $C_{\alpha}(x)$ are intervals, say, $a(x,\alpha),b(x,\alpha)]$, where $b\geq a$, but with $a(x,1)=b(x,1)=x/n$, the maximum likelihood estimator. Then the kernel $\mathsf{K}_{x}^{\alpha}$ associated with an inner probabilistic approximation is a probability distribution supported on the endpoints $a(x,\alpha)$ and $b(x,\alpha)$, i.e.,
$\mathsf{K}_{x}^{\alpha}=w_{x,\alpha}\,\delta_{a(x,\alpha)}+(1-w_{x,\alpha})\,\delta_{b(x,\alpha)},$
where $w_{x,\alpha}\in[0,1]$ and $\delta_{z}$ denotes a probability measure that puts all its mass at $z$. The question I want to address here is: what do the inner probabilistic approximations $\mathsf{Q}_{x}^{\star}$ look like? With constant weights, i.e., $w_{x,\alpha}\equiv w$, the argument used in the Gaussian example above can be generalized as follows. For a generic $\theta$, it follows that
$\{a(x,\alpha)\geq\theta\}\;\text{or}\;\{b(x,\alpha)\leq\theta\}\iff\pi_{x}(\theta)\leq\alpha.$
Using the mixture characterization in (8), the distribution function of $\Theta\sim\mathsf{Q}_{x}^{\star}$, at $\theta$ is
$\mathsf{Q}_{x}^{\star}\bigl((-\infty,\theta]\bigr)$ $=\int_{0}^{1}w\,1\{a(x,\alpha)\leq\theta\}\,d\alpha+\int_{0}^{1}(1-w)\,1\{b(x,\alpha)\leq\theta\}\,d\alpha$
$=\begin{cases}w\,\pi_{x}(\theta)&\text{if $\theta\leq x/n$}\\ 1-(1-w)\,\pi_{x}(\theta)&\text{if $\theta>x/n$}.\end{cases}$
Two such distribution functions are shown in Figure 2(a), one based on my recommended choice $w=0.5$ (solid) and another based on $w\approx 0.45$ (dashed). A third distribution function (dotted) is also shown, namely, that of the Bayesian posterior based on the right Haar prior with density $\theta\mapsto\theta^{-1}$, which is an inverse gamma distribution. Unlike the flat (right Haar) prior Bayes posterior in the Gaussian example above, this Bayes posterior can’t be expressed via a constant-weight kernel (Appendix B); the choice $w\approx 0.45$ was my best effort to do so. But that this Bayes solution is also an inner probabilistic approximation is a general fact established in Section 4.1 below. The three distribution functions are very similar, but not the same. To confirm that all three are, indeed, inner probabilistic approximations, I plot, in Figure 2(b), the non-coverage probability functions $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$. According to (9), the non-coverage probability for each inner probabilistic approximation should be equal to $\alpha$, which is confirmed in the plot.
(a) Distribution function of $\mathsf{Q}_{x}^{\star}$
(b) Non-coverage probability of $\mathsf{Q}_{x}^{\star}$
Figure 2: Plots of three inner probabilistic approximations for the gamma illustration with $n=7$ and $x=14$: uniform kernel (solid), non-uniform kernel with $w\approx 0.45$ (dashed), and the Bayesian posterior distribution (dotted) based on the default prior $\theta\mapsto\theta^{-1}$. Panel (a) shows the distribution function $t\mapsto\mathsf{Q}_{x}^{\star}(\Theta\leq t)$ and Panel (b) shows the non-coverage probability function $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$.
4 Properties
4.1 Agrees with Bayes in group invariant models
Let $\mathscr{G}$ denote a group of transformations from $\mathbb{X}$ to $\mathbb{X}$, with function composition $\circ$ as the binary operation. As is customary, I’ll write $gx$ for the image of $x\in\mathbb{X}$ under transformation $g\in\mathscr{G}$. Since $\mathscr{G}$ is a group, it’s associative, i.e., $g_{1}\circ(g_{2}\circ g_{3})=(g_{1}\circ g_{2})\circ g_{3}$ for all $g_{1},g_{2},g_{3}\in\mathscr{G}$, it contains the identity transformation, and for every $g\in\mathscr{G}$, there exists an inverse $g^{-1}\in\mathscr{G}$ such that $g\circ g^{-1}=g^{-1}\circ g=\text{identity}$. Examples of transformations include translation, scaling, rotation, permutation, etc.
The group $\mathscr{G}$ connects to the statistical model as follows. Suppose that, for each $g\in\mathscr{G}$ and each $\theta\in\mathbb{T}$, there exists a corresponding $\bar{g}\theta\in\mathbb{T}$ such that
$\mathsf{P}_{\theta}(gX\in\cdot)=\mathsf{P}_{\bar{g}\theta}(X\in\cdot),\quad(\theta,g)\in\mathbb{T}\times\mathscr{G}.$ (10)
For example, if the distribution of $X$ depends on a location parameter $\theta$, then the distribution of $X+\tau$ depends on parameter $\theta+\tau$. When the statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ satisfies (10), it’s called an invariant statistical model—that is, the model for $X$ is also the model for $gX$, with $g\in\mathscr{G}$. Let $\overline{\mathscr{G}}$ denote the collection of transformations $\bar{g}$ on $\mathbb{T}$, corresponding to the mappings $g\in\mathscr{G}$ and parameters $\theta\in\mathbb{T}$; it’s easy to check that $\overline{\mathscr{G}}$ is a group too. More details on equivariant models can be found in Fraser, (1968), Eaton, (1989), and Schervish, (1995, Ch. 6), among others. The necessary technical background is given in Appendix C of the supplementary material.
For invariant models, a standard and broadly agreed upon no-prior Bayes solution $\mathsf{Q}_{x}^{\rho}$ is that based on combining the likelihood with a so-called right Haar prior $\rho$ using Bayes’s formula (e.g., Kass and Wasserman, 1996). This also agrees with the solution obtained via Fisher’s fiducial argument, among others; see Lindley, (1958).
To clarify the scope of the following result, I’m assuming that the group $\mathscr{G}$ with respect to which the model is invariant is unique. In that case, under very mild conditions, the right Haar measure $\rho$ and the corresponding Bayes solution $\mathsf{Q}_{x}^{\rho}$ is unique too. But there are models, including the multivariate Gaussian models studied in Sun and Berger, (2007), that are invariant with respect to more than one group. So, even though each group has a unique right Haar prior, the non-uniqueness of the group implies non-uniqueness of the right Haar prior and corresponding Bayes solution. This latter case is more nuanced and I’ll discuss these details in Appendix C.
Theorem 3.
If the model is invariant with respect to a unique group $\mathscr{G}$, then, under the conditions laid out in Appendix C, the posterior $\mathsf{Q}_{x}^{\rho}$ relative to the right Haar prior $\rho$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$.
That is, for invariant models, the standard no-prior Bayes solutions are inner probabilistic approximations of the possibilistic IM. This justifies my previous claim that the Bayes posterior in the gamma illustration, in Section 3.3, is an inner probabilistic approximation. This also serves as a general proof of concept for the no-prior Bayes solution that I’m proposing here: that is, at least in cases where the group is unique, the corresponding Bayes solutions are reasonable and widely used, so what I’m proposing must be equally reasonable and usable.
Remember that inner probabilistic approximations are not unique. So, the right Haar prior Bayes solutions might not exactly match my recommended strategy that defines $\mathsf{K}_{x}^{\beta}$ to be uniform on $\partial C_{\beta}(x)$; the illustration in Figure 2 is such a case. In fact, when there’s a Bayesian solution that’s an inner probabilistic approximation of the IM, then I might recommend the former since its computation is more familiar than the latter’s. But, like in Figure 2, the difference between each inner probabilistic approximations will be negligible and will vanish as the information in data increases; see Section 4.2.
As an illustration, suppose that observations correspond to directions—points on a unit circle in the plane—or simply angles with respect to a reference direction. Real-world applications involving these kind of data include wind and animal movement studies; see Mardia and Jupp, (2000) for a comprehensive treatment. More generally, direction measurements can be represented by points on the surface of a hypersphere. This is common in astronomy, where the position of a planet or star can be described by a point on the celestial sphere. The present example involves $n=25$ such planar directional observations, plotted in Figure 3. These data are simulated from the so-called von Mises distribution, parametrized by a mean direction/angle $\theta$ and a concentration parameter $\kappa$. Further details about this model and the corresponding solutions are given in Appendix C.2 in the supplement. Here I’ll take the concentration parameter $\kappa=4$ to be fixed and known. I’ll also interpret the data as angles (radians) so that the parameter $\theta$ acts like a location paramter. With this simplification, the $\theta$-only von Mises model is invariant with respect to addition modulo $2\pi$. Therefore, Theorem 3 implies that the Bayes solution is an inner probabilistic approximation of the possibilistic IM. Figure 4 shows, first, a plot of the possibilistic IM’s contour for these data, with its peak at the maximum likelihood estimator, $\hat{\theta}_{x}=1.35$ and, second, a histogram of samples drawn from the inner probabilistic approximation (with $\mathsf{K}_{x}^{\beta}$ a uniform distribution) and the no-prior Bayesian posterior density overlaid. The visual similarities provide confirmation of the claim in Theorem 3.
Figure 3: Plot of the directional data analyzed in Section 4.1. Points on the circle determine the direction. The small closed circle denotes the sample mean in Cartesian coordinates and the corresponding angle determined by that point, $\hat{\theta}_{x}=1.35$, is the maximum likelihood estimator of the mean angle. The length of that vector, 0.87, is a to-be-conditioned-on ancillary statistic; see Appendix C.2
(a) Possibilistic IM contour
(b) Inner probabilistic approximation
Figure 4: Analysis of the directional data in Section 4.1. Panel (a) shows the possibilistic IM contour and Panel (b) shows a histogram of samples from the corresponding inner probabilistic approximation with the flat-prior Bayes posterior density overlaid.
4.2 Asymptotic normality and efficiency
Next is a version of the celebrated Bernstein–von Mises theorem for the possibilistic IM, establishing its asymptotic normality and efficiency. That is, for large $n$, the possibilistic IM’s contour resembles a Gaussian possibility contour, defined below, with covariance matrix that matches the Cramér–Rao lower bound. This implies a traditional Bernstein–von Mises theorem for the proposed inner probabilistic approximation and, therefore, it agrees with any other reasonable no-prior Bayesian solution asymptotically. Moreover, this near-Gaussian form of the IM output offers some valuable insights and simplifications relevant to computing the inner probabilistic approximation; see Section 5. The first order of business is to define Gaussian possibility.
Let $g_{m,v}$ denote the $D$-dimensional Gaussian probability density function, parametrized by a mean vector $m\in\mathbb{R}^{D}$ and a covariance matrix $v\in\mathbb{R}_{+}^{D\times D}$. Define the corresponding Gaussian possibility measure as the outer possibilistic approximation of ${\sf N}_{D}(m,v)$, the probability-to-possibility transform applied to the aforementioned Gaussian distribution. The corresponding possibility contour is
$\gamma_{m,v}(y)=\mathsf{P}\{g_{m,v}(Y)\leq g_{m,v}(y)\},\quad Y\sim{\sf N}_{D}(m,v),$
which can easily be rewritten as
$\gamma_{m,v}(y)$ $=1-F_{D}\bigl((y-m)^{\top}v^{-1}(y-m)\bigr),$
where $F_{D}$ is the ${\sf ChiSq}(D)$ distribution function. Write $\mathsf{\Gamma}_{m,v}$ for the corresponding possibility measure obtained by maximizing the contour as described above. Other authors have used a different form of Gaussian possibility (e.g., Denœux, 2023b ; Denœux, 2023a ), but I find the one above appealing and convenient because of the following duality between ${\sf N}_{D}(m,v)$ and $\mathsf{\Gamma}_{m,v}$: the former is the inner probabilistic approximation of the latter, and the latter is the outer possibilistic approximation of the former.
For some intuition about how this is related to the possibilistic IMs, consider a simple Gaussian location problem, $X\sim\mathsf{P}_{\Theta}={\sf N}_{D}(\Theta,v)$, with known covariance matrix $v$. Then the relative likelihood is $R(x,\theta)=\exp\{-\frac{1}{2}(x-\theta)^{\top}v^{-1}(x-\theta)\}$ and the IM’s contour is
$\pi_{x}(\theta)$ $=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}$
$=1-F_{D}\bigl((x-\theta)^{\top}v^{-1}(x-\theta)\bigr),\quad\theta\in\mathbb{R}^{D}.$
This is exactly the Gaussian possibility contour with mean $x$ and covariance matrix $v$. More generally, by the famous theorem of Wilks, (1938), the relative likelihood $R(X^{n},\theta)$, as a function of iid samples $X^{n}=(X_{1},\ldots,X_{n})$ from $\mathsf{P}_{\theta}$, will be asymptotically distributed as $\exp\{-\frac{1}{2}{\sf ChiSq}(D)\}$. Then the following approximation is expected when $n$ is large,
$\pi_{x^{n}}(\theta)\approx 1-F_{D}\bigl((\hat{\theta}_{x^{n}}-\theta)^{\top}J_{x^{n}}(\hat{\theta}_{x^{n}}-\theta)\bigr),$
where $\hat{\theta}_{x^{n}}$ is the maximum likelihood estimator and $J_{x^{n}}$ is the observed Fisher information matrix. The reader will notice that the right-hand side above is exactly the Gaussian possibility contour with mean vector $\hat{\theta}_{x^{n}}$ and covariance matrix $J_{x^{n}}^{-1}$, hence the connection between the possibilistic IM and the Gaussian possibility is very natural and holds quite generally, at least approximately. Making this connection rigorous requires care, both in formulating the proper regularity conditions and in suitably bounding all the error terms. Along these lines, under regularity conditions comparable to those advanced in Le Cam, (1970), Martin and Williams, (2025) proved the following Bernstein–von Mises theorem for the possibilistic IM.
Theorem 4 (Martin and Williams, 2025).
Under the regularity conditions presented in Appendix D, as $n\to\infty$, the possibilistic IM’s contour $\pi_{X^{n}}$ satisfies
$\sup_{\theta\in\mathcal{T}}\bigl|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\bigr|\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability},$
where $\mathcal{T}$ is a fixed but arbitrary compact subset of $\mathbb{T}$,
$\gamma_{X^{n}}(\theta):=\gamma_{\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1}}(\theta)$
is the centered and scaled Gaussian possibility contour, $I_{\Theta}$ the Fisher information matrix, and $\Delta_{\Theta}(X^{n})$ a sequence that converges in distribution to ${\sf N}_{D}(0,I_{\Theta}^{-1})$. In addition, if the likelihood is twice differentiable, then, with $\mathsf{P}_{\Theta}$-probability converging to 1 as $n\to\infty$,
$\sup_{u\in\mathcal{U}}\bigl|\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}u)-\gamma(u)\bigr|\to 0,$
where $\mathcal{U}$ is a fixed but arbitrary compact subset of $\mathbb{R}^{D}$, and $\gamma$ is the standard Gaussian possibility with zero mean and identity covariance matrix.
Among other things, Theorem 4 implies consistency: $\mathsf{\Pi}_{X^{n}}(H)\to 0$ in $\mathsf{P}_{\Theta}$-probability for any $H$ with $\Theta\not\in H$. Moreover, since the covariance matrix in the Gaussian limit agrees with the Cramér–Rao lower bound, Theorem 4 implies that the possibilistic IM is asymptotically efficient and, hence, there’s no loss of efficiency associated with the IM’s exact validity (and imprecision). Martin and Williams, (2025) give numerical illustrations that highlight the accuracy of the Gaussian approximation.
A direct consequence of Theorem 4 is that the IM’s inner probabilistic approximation converges to a Gaussian probability distribution just like in the classical, Bayesian Bernstein–von Mises theorem. The proof is in Appendix D of the supplementary material.
Theorem 5.
Under the conditions of Theorem 4, the proposed inner probabilistic approximation $\mathsf{Q}_{X^{n}}^{\star}$ of the likelihood-based possibilistic IM satisfies a probabilistic Bernstein–von Mises theorem, i.e.,
$d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}}):=\sup_{H}\bigl|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr|\to 0,$
where the convergence is in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$, where
$\mathsf{G}_{X^{n}}={\sf N}_{D}(\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1})$
is the approximating Gaussian distribution, and the supremum is over all Borel measurable subsets $H$ of $\mathbb{T}$.
I’ll give a quick proof sketch for a scalar $\Theta$. As in Section 3.3, using $w=\frac{1}{2}$, the distribution function of $\mathsf{Q}_{x^{n}}^{\star}$ is
$\mathsf{Q}_{x^{n}}^{\star}\bigl((-\infty,\theta]\bigr)$ $=\begin{cases}\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta\leq\hat{\theta}_{x^{n}}$}\\ 1-\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta>\hat{\theta}_{x^{n}}$}.\end{cases}$
Switching to local coordinates, i.e., $\theta=\hat{\theta}_{x^{n}}+J_{x^{n}}^{-1/2}z$ for $z$ fixed, the case $\theta\leq\hat{\theta}_{x^{n}}$ corresponds to $z\leq 0$ and Theorem 4 implies that, with $Z\sim{\sf N}(0,1)$,
$\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)$ $\to\tfrac{1}{2}\,\gamma(z)$
$=\tfrac{1}{2}\,\mathsf{P}(Z^{2}\geq z^{2})$
$=\mathsf{P}(Z\leq z).$
Similarly, if $z>0$, then $1-\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)\to\mathsf{P}(Z\leq z)$. Therefore, with probability tending to 1, the distribution of $J_{X^{n}}^{1/2}(\Theta-\hat{\theta}_{X^{n}})$, under $\Theta\sim\mathsf{Q}_{X^{n}}^{\star}$, is converging to a standard normal, just like in the classical Bernstein–von Mises theorem.
A point that follows immediately from the above sketch is that assigning probability $\frac{1}{2}$ to the boundary points in the inner probabilistic approximation is crucial; of course, any choice that assigns probability to the boundary points that converges to $\frac{1}{2}$ works fine too. More generally, since $\mathsf{G}_{X^{n}}$ is an inner probabilistic approximation to the possibility measure with contour $\gamma_{X^{n}}$, and corresponds to taking a uniform kernel on the level sets of $\gamma_{X^{n}}$, it’s necessary that the choice of kernel defining $\mathsf{Q}_{X^{n}}^{\star}$ at least be asymptotically uniform. This observation adds further support for my recommendation above to adopt the uniform kernel in the inner probabilistic approximation.
4.3 Marginalization risks, or lack thereof
Except for the impossibly rare cases where a genuine prior distribution is available, reliable statistical inference is inherently imprecise—those familiar test and confidence procedures that control error rates all have an imprecise probabilistic characterization (Martin, 2021a ). The point is that there’s no single probability distribution that can reliably quantify uncertainty about the unknowns in a statistical model. So, insisting that uncertainty quantification be probabilistic has risks:
[Xie and Singh, (2013)] are thus recommending that we ignore the restriction to confidence sets or equivalent, and free confidence to allow the production of parameter distributions. Certainly distributions are easier to think about, are largely in accord with Fisher’s original proposal, and are more in the freedom of the Bayes approach, but they do overlook inherent risks… (Fraser, 2013)
Fraser’s “risks” primarily concern marginalization (e.g., Dawid et al., 1973; Fraser, 2011a ; Balch et al., 2019). Roughly, whatever desirable properties a no-prior Bayes solution for $\Theta$ has, the corresponding no-prior Bayes solution for $\Phi=m(\Theta)$ obtained by probabilistic marginalization generally doesn’t share those desirable properties. When Schweder and Hjort, (2013) warn that “joint [confidence distributions] should not be sought, we think, since they might easily lead the statistician astray,” they’re concerned about users being tempted to do these familiar probabilistic operations, thus creating a risk of unreliability. Fraser, 2011a and Fraser et al., (2016) highlight non-linearity in $m$ as a major contributor to this risk; see, also, Martin, 2024b and the discussion below.
It’s important to emphasize that no probabilistic uncertainty quantification is safe from all such risks, including my proposed inner probabilistic approximations. So, the goal of the present section is simply to understand how marginalization with these approximations works and to identify which marginal inferences are safe. To be safe from all such risks, one must break from the familiar probabilistic uncertainty quantification in one way or another: Grünwald, (2018) suggests explicitly restricting one’s probabilistic inferences to those that are safe, whereas I’ve suggested relaxing from probabilities to possibilities so that all inferences are safe (Martin, 2025b ).
There are different ways to transform a quantification of uncertainty about $\Theta$ to a one about $\Phi=m(\Theta)$. Start with a confidence set $C_{\alpha}(x)$. The image $C_{\alpha}^{m}(x)=m\{C_{\alpha}(x)\}$ is a confidence set for $\Phi$ because
$\Theta\in C_{\alpha}(x)\implies m(\Theta)\in C_{\alpha}^{m}(x).$ (11)
The opposite implication generally doesn’t hold, which is the origin of the aforementioned risk. So, while the transformed confidence set for $\Phi$ maintains the nominal coverage probability enjoyed by the original confidence for $\Theta$, it can be conservative.
Next, take a no-prior Bayes posterior $\mathsf{Q}_{x}$ for $\Theta$ and map it to the corresponding marginal posterior $\mathsf{Q}_{x}^{m}$ for $\Phi=m(\Theta)$ via the probability calculus. An advantage of this approach is its convenience: draw samples of $\Phi$ from $\mathsf{Q}_{x}^{m}$ by first sampling $\Theta$ from $\mathsf{Q}_{x}$ and then transforming them via $m$. But, by (11):
$\mathsf{Q}_{x}^{m}\{C_{\alpha}^{m}(x)\}\geq\mathsf{Q}_{x}\{C_{\alpha}(x)\},\quad\alpha\in[0,1].$ (12)
That is, $\mathsf{Q}_{x}^{m}$ is inclined to assign more of its mass to the marginal confidence sets than $\mathsf{Q}_{x}$ assigned to the original confidence sets. So, if $\mathsf{Q}_{x}$ was a confidence distribution (CD), i.e., it assigns probability $1-\alpha$ to $C_{\alpha}(x)$ for each $\alpha\in[0,1]$, then (12) suggests calling $\mathsf{Q}_{x}^{m}$ an over-confidence distribution (oCD). Illustrations of this over-confidence phenomenon are presented in Appendix E of the supplement. This can be problematic because surely the user of $\mathsf{Q}_{x}^{m}$ will disregard the safe confidence set $C_{\alpha}^{m}(x)$ and, instead, construct his/her own quantification of uncertainty directly based on the over-confident $\mathsf{Q}_{x}^{m}$, risking unreliability. This begs the question: when does $\mathsf{Q}_{x}^{m}$’s credible set agree with the safe $C_{\alpha}^{m}(x)$?
Let $\mathsf{\Pi}_{x}$ be the possibilistic IM as described above, quantifying uncertainty about $\Theta$, given $X=x$. One way—but not the only way, see Section 6—to carry out possibilistic marginalization is by applying the general extension operation (e.g., Zadeh, 1975, 1978). Extension yields a new marginal IM $\mathsf{\Pi}_{x}^{m}$ for $\Phi=m(\Theta)$ with possibility contour
$\pi_{x}^{m}(\phi)=\sup_{\theta:m(\theta)=\phi}\pi_{x}(\theta),\quad\phi\in m(\mathbb{T}).$ (13)
If the confidence set $C_{\alpha}(x)$ in question is determined by the contour $\pi_{x}$ as in (5), then its image $C_{\alpha}^{m}(x)$ under $m$ can be expressed as
$C_{\alpha}^{m}(x)=\{m(\theta):\pi_{x}(\theta)>\alpha\}=\{\phi:\pi_{x}^{m}(\phi)>\alpha\}.$
That the possibilistic marginalization strategy appears organically in the common-sense treatment of confidence sets is no coincidence; see Martin, 2021a . The inner probabilistic approximations of $\mathsf{\Pi}_{x}$ are confidence distributions, and these are characterized as mixtures of kernels supported on the boundaries of the confidence sets. Then the derived marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ clearly also can be characterized by a similar mixture, but $\mathsf{Q}_{x}^{m}$ would only be the inner probabilistic approximation of $\mathsf{\Pi}_{x}^{m}$ if the boundaries $\partial C_{\alpha}(x)$ of the original confidence sets get mapped via $m$ to the boundaries $\partial C_{\alpha}^{m}(x)$ of the marginal confidence sets. Define $m$ to be boundary-preserving if $m\{\partial C_{\alpha}(x)\}=\partial C_{\alpha}^{m}(x)$ for each $\alpha\in[0,1]$. Among other forms of mappings, linear transformations are boundary preserving because they attain their extrema on the boundary of their domain.
Theorem 6.
If $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$, and $m$ is a boundary-preserving mapping, then the marginal distribution $\mathsf{Q}_{x}^{m}$ for $\Phi=m(\Theta)$ derived from $\mathsf{Q}_{x}$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}^{m}$.
That the probability calculus leads to reliable marginalization effectively only when the transformation is linear doesn’t imply that reliable marginalization is out of reach in the non-linear case. It just means non-probabilistic marginalization strategies are needed. This isn’t unfamiliar: Bayesians choose default prior distributions tailored to the particular target $\Phi$ (e.g., Bernardo, 1979; Datta and Ghosh, 1995), and fiducialists apply various manipulations akin to those used to construct approximate pivotal quantities. These insights motivate the following suggestion: first carry out the appropriate possibilistic marginalization to get $\pi_{x}^{m}$ and $\mathsf{\Pi}_{x}^{m}$, and then construct the inner probabilistic approximation. An example of this using the extension-based marginalization is shown below. Extension-based marginalization is a simple and very general approach, but is not the most efficient. Typically, sharper inferences can be drawn when marginalization is carried out by “validifying” a relative profile likelihood. Details of this profile-based approach will be presented in the context of a concrete example in Section 6 below.
Here I’ll revisit the directional data example presented in Section 4.1 above, where the quantity of interest is $\Phi=\cos(a\Theta)$, for two different values of $a$. Start with the case of $a=1$. Since the Bayes/fiducial distribution—which is also an inner probabilistic approximation of the possibilistic IM by Theorem 3—is almost fully supported on the interval $(0,\pi)$, where $\theta\mapsto\cos(\theta)$ is one-to-one, this mapping is effectively boundary-preserving. Then Theorem 6 implies that the marginal Bayes/fiducial distribution for $\Phi$ is an inner probabilistic approximation of the marginal possibilistic IM for $\Phi$. Figure 5 shows, for $a=1$, the extension-based marginal IM contour for $\Phi$, samples of the marginal Bayes/fiducial with the corresponding density function overlaid, and the non-credibility function $\alpha\mapsto 1-\mathsf{Q}_{x}^{m}\{C_{\alpha}^{m}(x)\}$, for $\mathsf{Q}_{x}^{m}$ equal to the Bayes/fiducial marginal distribution and to the inner probabilistic approximation of the extension-based marginal IM. The key observation is that the contour and the posterior distribution have similar shapes and, as Theorem 6 predicts, the non-credibility function behaves exactly as it should for an inner probabilistic approximation, according to the definition (9). Again, this all happens because $\theta\mapsto\cos(a\theta)$ with $a=1$ is approximately boundary-preserving on the interval where the relevant posterior distributions are supported.
(a) Marginal IM contour
(b) Bayes marginal
(c) Non-credibility
Figure 5: Marginal inference results for $\Phi=\cos(a\Theta)$, with $a=1$, in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
Next, take $a=1.5$. Then the “effectively boundary-preserving” property no longer holds and, surprisingly, this small change in the quantity of interest has non-negligible impact on the results. Figure 6 shows the same plots as in Figure 5, this time for $a=1.5$. The key observation is that, while the general shape of the marginal IM contour and the Bayes posterior distribution were similar in Figure 5, they are quite different in Figure 6; this is a consequence of the entirely different style of marginalization, from $\Theta$ to $\Phi$, with the IM’s possibilistic marginalization preserving the validity properties enjoyed by both approaches’ inference on $\Theta$ whereas the Bayesians’ probabilistic marginalization doesn’t. The IM’s inner probabilistic approximation (not shown) looks similar to the Bayesian posterior, but it’s not the same. Indeed, Figure 6(c) shows the non-credibility for the Bayes posterior distribution and the IM’s inner probabilistic approximation, and there are clear signs of the Bayesians’ over-confidence. The IM’s inner approximation also shows (milder) signs of over-confidence, but there’s a perfectly reasonable explanation for this. Recall from Section 3 that exact inner probabilistic approximations need not exist, e.g., if the possibility contour is bounded away from 0. In Figure 6(a), note that the contour is in fact bounded away from 0 and, consequently, the sets $C_{\alpha}^{m}(x)$ are equal to the full space [1,1][-1,1] for sufficiently small $\alpha$. Since any probability distribution $\mathsf{Q}_{x}^{m}$ supported on [1,1][-1,1] must assign probability $1>1-\alpha$ to $C_{\alpha}^{m}(x)=[-1,1]$ for all sufficiently small $\alpha$, there is no $\mathsf{Q}_{x}^{\star}$ that satisfies (9) for all $\alpha$. But the IM’s pseudo inner probabilistic approximation—“pseudo” because the inner probabilistic approximation strictly doesn’t exist—quickly corrects for this over-confidence at small $\alpha$ and achieves the objective (9) for all $\alpha$ that are not too small.
(a) Marginal IM contour
(b) Bayes marginal
(c) Non-credibility
Figure 6: Marginal inference results for $\Phi=\cos(a\Theta)$, with $a=1.5$, in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
5 Computation
Until recently, only naive/inefficient strategies for computing the IM contour were available. In particular, the go-to strategy was to approximate $\pi_{x}$ via
$\pi_{x}(\theta)\approx\frac{1}{M}\sum_{m=1}^{M}1\{R(X_{m,\theta},\theta)\leq R(x,\theta)\},$ (14)
where $X_{m,\theta}$ are independent copies of the data $X$, drawn from $\mathsf{P}_{\theta}$, for $m=1,\ldots,M$. The above computation is feasible at a few different $\theta$ values, but all the practically relevant calculations—e.g., identifying the confidence set in (5)—require evaluation over a sufficiently fine grid covering the relevant portions of $\mathbb{T}$, and this is expensive. Recently, Martin, 2025a developed a new and more efficient strategy for doing the relevant IM computations that replaces most of the naive contour evaluations as in (14) with Monte Carlo sampling from a sort of “posterior distribution.” That distribution is, in fact, the inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ of $\mathsf{\Pi}_{x}$ defined in Section 3 above. Therefore, the same computational strategy for (approximately) sampling from $\mathsf{Q}_{x}^{\star}$ advanced in Martin, 2025a can be reIMagined here as an implementation of the new no-prior Bayes solution.
The strategy in Martin, 2025a is, roughly, to stitch together a suitable collection of Gaussian possibilistic IMs. While the “stitched” collection of Gaussian possibilities isn’t a Gaussian possibility, the use of a basic Gaussian building block can be justified by the asymptotic normality result in Section 4.2. Interest here is solely in the inner probabilistic approximation, so it’s not necessary for me to describe all the details involved in the “stitched IM.” For $\sigma$ a generic $D$-vector of positive numbers, consider a collection of Gaussian confidence/level sets
$C_{\alpha}^{\sigma}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}\,J_{x}^{\sigma}\,(\theta-\hat{\theta})\leq\chi_{D,1-\alpha}^{2}\},$
where $\chi_{D,1-\alpha}^{2}$ is the $1-\alpha$ quantile of the ${\sf ChiSq}(D)$ distribution, and $J_{x}^{\sigma}$ is an embellished version of the observed Fisher information matrix with spread at least partially controlled by the parameters in $\sigma$. Specifically, if $J_{x}$ has spectral decomposition $J_{x}=E\Lambda E^{\top}$, then the embellished version $J_{x}^{\sigma}$ is defined as
$J_{x}^{\sigma}=E\,\text{diag}(\sigma^{-1})\,\Lambda\,\text{diag}(\sigma^{-1})\,E^{\top},$ (15)
where $\text{diag}(\sigma^{-1})$ is the diagonal matrix with elementwise reciprocal of the $\sigma$ vector on the diagonal. Then $\sigma$ controls confidence set’s size by manipulating the eigenvalues of the underlying observed Fisher information matrix, e.g., a large (resp. small) entry in $\sigma$ stretches (resp. contracts) the level set in the direction of the corresponding eigenvector. Theorem 4 says that the original $C_{\alpha}(x)$ will be roughly elliptical, so it’s safe to assume that there exists a vector $\sigma$ such that the following inclusion holds snugly:
$C_{\alpha}(x)\subseteq C_{\alpha}^{\sigma}(x).$ (16)
Of course, this vector $\sigma$ generally depends on $(x,\alpha)$, so I’ll denote this as $\sigma(x,\alpha)$. A cartoon depiction of the proposed strategy is shown in Figure 7. The motivation for bounding $C_{\alpha}(x)$ inside $C_{\alpha}^{\sigma}(x)$ is that it’s not possible to describe the former exactly but the latter is very simple, at least for fixed $\sigma$. The data-driven algorithm recommended in Martin, 2025a , relying on the foundations laid in Cella and Martin, (2025), for selecting $\sigma(x,\alpha)$ is presented in Appendix F in the supplement. Once this $\sigma(x,\alpha)$ has been determined, the inner probabilistic approximation uses the kernel
$\mathsf{K}_{x}^{\alpha}={\sf Unif}\{\partial C_{\alpha}^{\sigma(x,\alpha)}(x)\},\quad\alpha\in(0,1).$
Thanks to the direction of the inclusion in (16), the proposed implementation of the inner probabilistic approximation is conservative. So, the relevant properties enjoyed in theory by the inner probabilistic approximation of the IM should be at least approximately enjoyed by the practical proposal above; see Section 6.
$\hat{\theta}_{x}$$C_{\alpha}(x)$$\partial C_{\alpha}^{\sigma(x,\alpha)}(x)$
Figure 7: Approximating the IM contour’s $\alpha$-level set $C_{\alpha}(x)$ (gray) by an ellipse $C_{\alpha}^{\sigma}(x)$ with a “good” choice of $\sigma$ (solid) and with a “bad” choice of $\sigma$ (dashed).
6 Example: Behrens–Fisher problem
For brevity, I only give one detailed illustration of the inner probabilistic approximation here, but two additional examples are presented in Appendix G.
Few examples have deeper practical and historical relevance than the Behrens–Fisher problem (Fisher, 1935a ; Fisher, 1939). The original problem is simple to state: independent samples of size $n_{1}$ and $n_{2}$, respectively, are available from two distinct normal populations, ${\sf N}(\Theta_{11},\Theta_{12}^{2})$ and ${\sf N}(\Theta_{21},\Theta_{22}^{2})$, with $\Theta=(\Theta_{11},\Theta_{12},\Theta_{21},\Theta_{22})$ unknown, and the goal is marginal inference on the difference $\Phi=m(\Theta)=\Theta_{21}-\Theta_{11}$ of the two means. The problem is straightforward if the two variances are known or if their ratio is known, but the fully unknown variance case has remained elusive, i.e., lots of candidate solutions are available but, despite its apparent simplicity and practical relevance, there is no consensus on which solution is “right” or “best.” For a review of the problem and various solutions, see Kim and Cohen, (1998).
By far the most widely used solution to the Behrens–Fisher problem is the modification of the basic Student-t pivot with the degrees of freedom approximations due to Welch, (1938, 1947); this is implemented in R’s t.test function. Other standard approaches include the simple-but-conservative solution proposed by Hsu, (1938) and Scheffé, (1970), and the Bayesian solution proposed by Jeffreys, (1940), based on the right Haar prior for $\Theta$, which is mathematically equivalent to Fisher’s fiducial solution. Ironically, the solution proposed by Jeffreys is different from the Bayesian solution based on the Jeffreys prior, both in its construction and its performance.
To set the scene, start by considering inference on the full parameter $\Theta$. This model has lots of structure so the relative likelihood-based, possibilistic IM construction for inference on $\Theta$ is conceptually and computationally straightforward. Thanks to the model’s underlying affine group invariance, it follows from Theorem 3 that the corresponding inner probabilistic approximation for $\Theta$ is exactly the Bayesian posterior distribution for $\Theta$ based on the right Haar prior, which is also Fisher’s fiducial distribution for $\Theta$. Since the mapping from $\Theta$ to $\Phi$ is linear, it follows from Theorem 6 that the (mathematically equivalent) Bayes and fiducial solutions proposed by Jeffreys and Fisher, respectively, correspond to the marginal distribution for $\Phi$ derived from the IM’s inner probabilistic approximation; moreover, these are also the inner probabilistic approximation to the marginal IM based on following the extension principle.
The new approach I’m proposing here first carries out a different—and generally more efficient—IM marginalization strategy, based on validifying the profile relative likelihood. This idea was first developed in Martin, 2023b and an illustration of its use in the Behrens–Fisher problem is considered in Example 5 there. Roughly speaking, the difference between this profile likelihood-based marginal IM construction and the aforementioned extension-based construction is where the marginalization is carried out: the former first eliminates nuisance parameters in the relative likelihood and then constructs the IM contour directly for $\Phi$, whereas the latter first constructs the IM contour for $\Theta$ and then marginalizes to $\Phi$. Martin and Williams, (2025) recently confirmed that, at least asymptotically, the profile likelihood-based marginal IM construction was more efficient than the extension-based construction. My proposal here is to extract the inner probabilistic approximation of this profile likelihood-based marginal possibilistic IM for $\Phi$. This requires no change in the computational approach outlined in Section 5, and the exact validity of the marginal IM implies, e.g., that credible intervals for $\Phi$ based on this “posterior” are exact confidence intervals. The only wrinkle is that the relative profile likelihood doesn’t have a closed-form expression, and its distribution depends on a nuisance parameter; this is precisely what makes the Behrens–Fisher problem challenging. This increases the computational cost of the inner probabilistic approximation compared to other examples (see Appendix G), but this pays off in terms of efficiency gains as I demonstrate below.
The go-to real-data illustration of the Behrens–Fisher problem is the example in Lehmann, (1975, p. 83) on travel times to work via two different routes. The relevant summary statistics—sample sizes, sample means, and sample standard deviations—are as follows: $n_{1}=5$, $\hat{\theta}_{11}=7.580$, and $\hat{\theta}_{12}=2.237$; $n_{2}=11$, $\hat{\theta}_{21}=6.136$, and $\hat{\theta}_{22}=0.073$. The wide discrepancy between the two standard deviations, $\hat{\theta}_{12}$ and $\hat{\theta}_{22}$, would make it difficult to justify assuming the two variances are equal. Figure 8 shows a histogram of the samples of $\Phi$ from the inner probabilistic approximation of the marginal IM, with (kernel estimates) of the density functions from the right Haar and Jeffreys prior Bayes solutions overlaid; the former also agrees with Fisher’s fiducial distribution. The point is that all three distributions are similar, with the new inner probabilistic approximation and the right Haar prior posterior being a bit more diffuse than the Jeffreys prior posterior.
Figure 8: Histogram summarizing the distribution of $\Phi$ in the Behrens–Fisher problem based on the marginal IM’s inner probabilistic approximation. The solid and dashed lines are (kernel estimates of) the Bayesian posterior distributions for $\Phi$ based on the right Haar and Jeffreys priors, respectively.
For further comparison, I run a small simulation study. I focus on a rather severely imbalanced case—with $n_{1}=3$ and $n_{2}=20$—to ensure that the differences in performance are apparent. Otherwise, the simulation settings are standard: $\Theta_{11}=2$, $\Theta_{21}=0$, $\Theta_{12}^{2}=1$, and $\Theta_{22}^{2}=2$. I generated 10000 samples under this setup, and Table 1 presents the coverage probability and expected length of the various 90% confidence intervals for $\Phi$. Notably, only the marginal possibilistic IM’s inner probabilistic approximation is able to hit the target coverage probability nearly on the nose, and, as desired, it’s more efficient than the valid-but-conservative Bayes/fiducial solution based on the right Haar prior.
Method Coverage Length
Bayes with Jeffreys prior 0.861 2.31
Bayes with right Haar prior 0.929 3.26
t.test with Welch df 0.881 2.79
Marginal IM’s inner approximation 0.904 2.71
Table 1: Coverage probabilities and mean lengths for the 90% confidence intervals returned by the various methods. Standard errors for all of the coverage probabilities and mean lengths are roughly 0.003 and 0.01, respectively.
7 Conclusion
This paper offers a new perspective on no-prior Bayes inference, one that has certain close ties to Bayesian inference but is not actually Bayesian at all. My approach starts with an inferential model (IM) framework for data-driven uncertainty quantification that prioritizes reliability, insisting on calibration of its data-dependent degrees of belief. The kind of calibration IMs require makes them incompatible with probabilistic Bayesian inference, and it’s for this reason that the IM’s output is possibilistic, i.e., takes the mathematical form of a possibility measure. However, if one desires probabilistic uncertainty quantification, then the IM can accommodate this with an inner probabilistic approximation of its possibilistic output. Aside from being exactly probability matching, the proposed solution agrees with existing no-prior Bayes solutions in group equivariant problems where there’s an overall consensus, and is also asymptotically efficient by a version of the celebrated Bernstein–von Mises theorem. And there’s literally no choice of prior distribution required: for the given model and data, the possibilistic IM is what it is, so it’s just a matter of extracting a suitable inner probabilistic approximation.
The characterization of the contents of the IM’s credal set in Theorem 1 suggests a strategy for evaluating, at least approximately, the inner probabilistic approximation via Monte Carlo. I explored this thoroughly in Martin, 2025a , but mostly in the context of possibilistic IMs. Here I use this same strategy for my proposed reIMagined no-prior Bayesian inference, and I show that this leads to a new and broadly reliable solution to the technically challenging and practically relevant Behrens–Fisher problem. Nothing about this approach was tailored to the Behrens–Fisher problem specifically, so I fully expect similarly good performance in many other important applications. Surely, the computational efficiency of my proposed solution can be improved, so I welcome advancements along these lines from computationally savvy readers.
The focus here has been on no-prior solutions, but there are cases in which incomplete or partial prior information is available, including high-dimensional problems in which structural assumptions like sparsity are common. Bayesian solutions can’t work with incomplete prior information—a prior distribution, perhaps diffuse, must be specified for every aspect of the unknown $\Theta$. Genuine partial-prior possibilistic IMs have recently been developed (e.g., Martin, 2022b ), and an interesting idea would be to apply the same inner probabilistic approximation proposed here in the no-prior case to the partial-prior case in the above references. Just as in this paper, these inner probabilistic approximations would enjoy some of the inherent reliability properties satisfied by the partial-prior IM, but the details remain to be investigated.
Acknowledgments
The author thanks the Editor, Associate Editor, and two anonymous reviewers for their valuable comments on an earlier version of this paper. This work is partially supported by the U.S. National Science Foundation, grants SES–2051225 and DMS–2412628.
Appendix A Further IM insights, intuition, etc.
In the main text, I gave only the mathematical definition and briefly explanation of the possibilistic inferential model (IM) and its basic properties. But these developments are deeper and more insightful than can be communicated in such a brief explanation, so I’ll share a bit more here about the IM intuition and logic. I’ll organize this presentation as a list of answers to some common/frequently-asked questions.
Question 1.
Why is imprecision needed? The false confidence theorem (Balch et al., 2019) says that for any precise data-dependent probability distribution $\mathsf{Q}_{X}$, e.g., a Bayesian or fiducial posterior, and for any thresholds $\rho$ and $\tau$, there exists a hypothesis $H$ such that
$H\not\ni\Theta\quad\text{and}\quad\mathsf{P}_{\Theta}\{\mathsf{Q}_{X}(H)>\tau\}>\rho.$
This is called false confidence because, with $\tau$ large and $\rho$ not small, the posterior tends (at non-negligible rate $\rho$) to be confident (high probability $\tau$) in the truthfulness of $H$ even though $H$ is actually false. There’s a sense in which this result is trivial: take $H$ equal to the complement of $\Theta$, so that $H$ is false and, at least in the absolutely continuous case, $\mathsf{Q}_{X}(H)=1$ almost surely. But there’s more nuance here because it’s not just for these extreme/trivial $H$ that false confidence manifests; it’s not even just for “big” sets that happen to be false. The shape of $H$ and the effects of non-linearity, just like in Section 4.3 in the present paper, play an important albeit not yet well understood role.
In addition, there’s no reason why the size of a hypothesis should matter at all—data are compatible with a hypothesis as soon as they’re compatible with any single point in the hypothesis. Size/measure only becomes relevant when integration enters the picture, but there’s no justification for making this step except that statisticians can’t help it, the theory of integration is burned into our minds. Ideally, integration could be replaced by some other calculus that doesn’t rely on notions of size, no differential element.
Fortunately, both of the issues identified above can be resolved at once if ordinary probability can be replaced by a suitable imprecise probability. The proof of the false confidence theorem leverages the additivity of the posterior, so, if the additivity assumption is dropped, then the proof breaks down. But, also, certain data-dependent imprecise probabilities, namely, (possibilistic) IMs, have been shown to be immune to false confidence. Moreover, possibility theory specifically operates under a different calculus than integration: the possibilistic calculus is optimization, so there is no differential element and the size of hypotheses is irrelevant when it comes to if and to what extent they are compatible with data; see Martin, 2024a .
Question 2.
Why the possibilistic form specifically? It helps to start with being clear about what the objectives are when it comes to the formulation of a new and improved framework for statistical uncertainty quantification, so I list here three desiderata:
  1. 1.
    Mathematically rigorous. Statistical reasoning shouldn’t be based on a patchwork disparate pieces but, rather, a coherent whole from which an answer to any question about the relevant unknowns can be obtained.
  2. 2.
    As simple and “familiar-looking” as possible. Much of the theory that is currently being used for statistical inference (e.g., p-values for tests) generally works. So, a new framework doesn’t need to be completely different and, in fact, it shouldn’t give altogether different solutions to classical inference problems.
  3. 3.
    Reliable. Not only should procedures (tests, confidence sets, etc) derived from ones uncertainty quantification be reliable in the sense that they control frequentist error rates, but the (lower and upper) probabilities and expectations that it assigns to relevant hypotheses or assessments should be calibrated too. In fact, it’s this reliability property that gives the uncertainty quantification its meaning.
To my knowledge, the only options for a mathematically rigorous framework of uncertainty quantification are those that can be described by an imprecise probability of some sort. Possibility theory is arguably one of the simplest imprecise probability theories and, since the possibilistic IM framework that I’ve been advancing has close connections to p-values for likelihood ratio tests, confidence distributions, and now no-prior Bayes solutions, it clearly satisfies the first two desiderata.
The main text highlighted some of the possibilistic IM framework’s reliability properties, but there is a more to say as it concerns the topic of this section: why the possibilistic form? The property I referred to as validity in the main paper—Equation (4)—is actually quite strong. The frequentist error rate control guarantees it implies for test and confidence set procedures derived from the IM output are important, but doesn’t do justice to the strength of this property. As Martin, 2022a and Cella and Martin, (2023) have shown, the validity property in Equation (4) of the main paper is equivalent to
$\sup_{\Theta\in\mathbb{T}}\mathsf{P}_{\Theta}\{\mathsf{\Pi}_{X}(H)\leq\alpha\text{ for some $H$ with $H\ni\Theta$}\}\leq\alpha,\quad\alpha\in[0,1].$ (17)
This is similar to the property stated in Equation (6) of the main text, but with one key difference: the property in (6) establishes that the IM is calibrated “pointwise” in $H$, whereas (17) shows that, in fact, the IM is calibrated “uniformly” in $H$. This is relevant because modern statisticians and data scientists aren’t quantifying uncertainty for themselves, they’re developing general tools and software to assist other users in quantifying their statistical uncertainty. Since I can’t possibly know how a user might want to do with the uncertainty quantification framework that I offer them, if I want to guarantee that it’s safe—which is important both for science/society and for me as the developer of such a tool—then I must ensure that calibration is maintained even if the user is, e.g., peeking at the data first to decide what hypotheses to consider. The pointwise validity result in (6) of the main paper doesn’t guarantee this, but the uniform validity result in (17) does.
There are mathematical obstacles that would make it difficult for any non-possibilistic quantification of uncertainty to satisfy the strong validity property in Equation (4) of the main paper, and hence the uniform calibration property in (17) above. Indeed, if $\mathsf{\Omega}_{x}$ was some other upper probability quantifying uncertainty about $\Theta$, given $X=x$, then one could define the analogue of its “contour,” i.e., $\theta\mapsto\mathsf{\Omega}_{x}(\{\theta\})$. But this may not meet the requirement of a possibility contour, that is, generally,
$\sup_{\theta\in\mathbb{T}}\mathsf{\Omega}_{x}(\{\theta\})<1,\quad\text{and often $\ll 1$}.$
Nothing in imprecise probability theory says that this “contour” must satisfy this supremum-equals-1 property and, in fact, in many (non-statistical) cases there are good reasons why such a property would be undesirable. But, as Shafer, (1976, p. 226) explains, statistical inference problems are special and well-suited for this supremum-equals-1 property. What I can add to Shafer’s justification is that it is virtually impossible for anything other than a possibilistic quantification of uncertainty, one for which the supremum-equals-1 property holds, to satisfy the validity property as stated in Equation (4) of the main paper and, therefore, the equivalent uniform calibration property in (17) above. The point is that, if $\mathsf{\Omega}_{X}(\{\Theta\})$ is bounded away from 1 for almost all $X$, then it can’t be stochastically no less than ${\sf Unif}(0,1)$. Moreover, if one was able to find a non-possibilistic IM $(\underline{\Omega}_{x},\mathsf{\Omega}_{x})$ that does satisfy the validity property in Equation (4) in the main text, then there is a possibilistic IM that is valid and no less efficient. So, there is at least no loss of generality in focusing on IMs that take on the possibilistic form.
Question 3.
Beyond assigning beliefs to hypotheses about model parameters? Classical inference on the parameters is a fully-specified statistical model is the right context to describe this paper’s contribution, namely, a new perspective on no-prior Bayesian inference, but there are many other kinds of problems and statistical tasks that scientists are interested in. Fortunately, IMs can contribute along these lines too, and they already have. First, it’s often the case that the quantity of interest is not best described as the parameter of a statistical model, but as some feature (or functional) of the underlying distribution $\mathsf{P}$, which doesn’t have a familiar parametric form. The construction of approximately valid IMs for such cases is presented in Cella and Martin, 2022a ; work on removing the “approximately” adjective is ongoing. Prediction is also an important problem and usually requires a nonparametric touch. IM advances along these lines can be found in Cella and Martin, 2022b , with a more possibility-centric perspective offered in Martin, 2023b (, Sec. 6.3). Sometimes inference isn’t the objective, the focus is on decision-making, i.e., choosing an appropriate (data-driven) action. In Martin, 2021b , I propose the use of Choquet integration to use the IM output to evaluate upper (and lower) expected losses and recommend choosing actions that minimize the upper expected loss. I also show that the IM’s inherent reliability in inference carries over to the decision problem as well: it’s a provably rare event that the IM’s upper expected loss assessment is substantially more optimistic than an oracle’s, uniformly over actions. Finally, novel and provably reliable IM solutions for model assessment and selection, information fusion (e.g., meta-analysis), etc. are currently in development.
Appendix B Details from Section 3.2
A point emphasized in Section 3.2 is that inner probabilistic approximations of a given IM $\mathsf{\Pi}_{x}$ aren’t unique. The reason is that the characterization result and the “maximally diffuse” objective don’t determine exactly how the kernel $\mathsf{K}_{x}^{\alpha}$ allocates its mass to the boundary $\partial C_{\alpha}(x)$. The goal of the present section is to expand a bit on this ambiguity to better understand the role played by the kernel.
To start, I’ll focus on the case of a scalar parameter, like in the illustrations presented in Section 3. In such a case, the kernel can be expressed generally as
$\mathsf{K}_{x}^{\alpha}=w_{x,\alpha}\,\delta_{a(x,\alpha)}+(1-w_{x,\alpha})\,\delta_{b(x,\alpha)},$
where $w_{x,\alpha}$ is a data and level-dependent weight in $[0,1]$, $a(x,\alpha)$ and $b(x,\alpha)$ are the lower and upper endpoints of the confidence interval $C_{\alpha}(x)$, and $\delta_{z}$ is a probability distribution concentrating all of its mass on $z$. In this case, the kernel is fully determined up to the weight $w_{x,\alpha}$. My general recommendation was to take $w_{x,\alpha}$ to be constant equal to $\frac{1}{2}$, which has a two-fold motivation: it’s consistent with a version of the indifference principle and weights converging to $\frac{1}{2}$ is necessary in order for the inner probabilistic approximation to satisfy a Bernstein–von Mises theorem. Nevertheless, there’s still a choice of weight that can be made, so it’s worth having some understanding of what this entails.
For the gamma model $X\sim\mathsf{P}_{\Theta}={\sf Gamma}(n,\Theta)$ in Section 3.3 of the main text, Figure 2(a) shows the distribution functions associated with three different inner probabilistic approximations, each determined by a choice of the weight $w_{x,\alpha}$. Two of the three were based on constant-weight kernels: one with $w=0.5$ and one with $w\approx 0.45$; my rationale behind the choice $w\approx 0.45$ is described below. Since these weights are numerically very similar, it’s not surprising that the corresponding distribution functions are likewise similar. The third of the three distribution functions in Figure 2(a) is that based on the Bayes solution that employs the default right Haar prior with density $\theta\mapsto\theta^{-1}$ with respect to Lebesgue measure. I claimed in the main text that this Bayesian solution can’t be expressed in terms of the kernel above with constant weight; a non-constant weight $w_{x,\alpha}$ is needed. This begs the question: what non-constant weight corresponds to the right Haar prior Bayes solution?
In general, suppose that $Q_{x}(\theta)$ is the distribution function corresponding to an inner probabilistic approximation $\mathsf{Q}_{x}$; let $q_{x}$ be the corresponding density with respect to Lebesgue measure. Let $\hat{\theta}_{x}$ be the maximum likelihood estimator. Following the derivation in Section 3.3, for $\theta\leq\hat{\theta}_{x}$, the distribution function $Q_{x}(\theta)$ can be expressed as
$Q_{x}(\theta)=\int_{0}^{1}w_{\alpha}\,1\{a(x,\alpha)\geq\theta\}\,d\alpha=\int_{0}^{\pi_{x}(\theta)}w_{\alpha}\,d\alpha,$
where I’ve dropped the subscript “xx” in $w_{x,\alpha}$ for simplicity. Applying Leibniz’s integral rule gives the identity
$q_{x}(\theta)=w_{\pi_{x}(\theta)}\,\dot{\pi}_{x}(\theta),\quad\theta\leq\hat{\theta}_{x},$
where $\dot{\pi}_{x}(\theta)$ is the partial derivative of $\pi_{x}(\theta)$ with respect to $\theta$. This implies that
$w_{\alpha}=\frac{q_{x}(\theta_{x,\alpha})}{\dot{\pi}_{x}(\theta_{x,\alpha})},\quad\alpha\in[0,1],$ (18)
where $\theta_{x,\alpha}$ is such that $\theta_{x,\alpha}\leq\hat{\theta}_{x}$ and $\pi_{x}(\theta_{x,\alpha})=\alpha$. This calculation can’t be done analytically for the gamma illustration in Section 3.3, but it can be done numerically, and the plot of $\alpha\mapsto w_{\alpha}$ is shown in Figure 9. Notice that this curve ranges from below to above the $w=0.45$ line, with $w_{0.5}\approx 0.45$; this explains my choice of weight $w\approx 0.45$ as the best constant-weight approximation of the right Haar prior Bayes posterior.
Figure 9: Plot of $\alpha\mapsto w_{\alpha}$ as defined in (18) for the gamma model with $n=7$ and $x=14$.
The formula (18) for the non-constant weight is a bit abstract and can only be evaluated numerically in the gamma example above. So it may help to provide a simpler, proof-of-concept illustration—in particular, the Gaussian location model. In that case, the posterior density is $q_{x}(\theta)=(2\pi)^{-1/2}e^{-(\theta-x)^{2}/2}$, and the contour derivative is
$\dot{\pi}_{x}(\theta)=2(\theta-x)\times f_{1}((\theta-x)^{2}),$
where $f_{1}(z)=(2\pi)^{-1/2}z^{-1/2}e^{-z/2}$ is the ${\sf ChiSq}(1)$ density function. Then the ratio is
$\frac{q_{x}(\theta)}{\dot{\pi}_{x}(\theta)}=\frac{1}{2},\quad\theta<\hat{\theta}_{x}=x.$
This gives further justification for the choice of a constant weight $w=\frac{1}{2}$ in the IM’s inner probabilistic approximation’s kernel for the Gaussian location model.
More extreme choices of weights can be considered. For example, in the scalar parameter case, an extreme choice is taking $w=0$ or $w=1$. The large-sample results in Section 4.2 strongly suggest that such a choice isn’t warranted, so I wouldn’t recommend this. The point is that $w=0$ and $w=1$ both introduce bias that can’t be beneficial. When the parameter $\Theta$ is a vector or something more complex, then there are more choices of “extreme” kernels, some of which might not be entirely unreasonable. For instance, if $\Phi=m(\Theta)$ is a scalar feature of interest, then one might consider an “extreme” kernel that puts all of its weight on the extremes of the function $m$ on each confidence set boundary. That is, one might consider a kernel
$\mathsf{K}_{x}^{\alpha}=\tfrac{1}{2}\,\delta_{\underline{m}_{\alpha}}+\tfrac{1}{2}\,\delta_{\overline{m}_{\alpha}},$
where
$\underline{m}_{\alpha}=\min_{\theta\in\partial C_{\alpha}(x)}m(\theta)\quad\text{and}\quad\overline{m}_{\alpha}=\max_{\theta\in\partial C_{\alpha}(x)}m(\theta).$
With this choice, the marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ derived from the inner probabilistic approximation with the above kernel is my recommended inner probabilistic approximation of the marginal IM $\mathsf{\Pi}_{x}^{m}$ derived from $\Pi_{x}$ via extension. The point is that “extreme choices” of kernels might not be unreasonable, and the justification of such a choice would be based on marginalization considerations, when there’s a particular feature $\Phi=m(\Theta)$ of $\Theta$ of interest. When there’s no single feature that’s of high priority, it makes sense to make a neutral choice of kernel, and, aside from the large-sample properties, it’s precisely the uniform kernel’s neutrality that makes it a good choice.
Appendix C Details from Section 4.1
C.1 Theory
This section provides some additional background and details about the invariant model setting and the result in Theorem 3.
Recall that $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ is a family of probability distributions supported on $\mathbb{X}$ that is also invariant with respect to a group of transformations $\mathscr{G}$ from $\mathbb{X}$ to itself in the sense of Equation (10) in the main text. If the distributions $\mathsf{P}_{\theta}$ have densities with respect to some underlying dominating measure, which I’ll assume here, then invariance means
$p_{\theta}(x)=p_{\bar{g}\theta}(gx)\,\chi(g),\quad x\in\mathbb{X},\;\theta\in\mathbb{T},\;g\in\mathscr{G},$ (19)
where $\chi(g)$ is the “multiplier,” a change-of-variables Jacobian term; see Schervish, (1995, Ch. 6) and Eaton, (1989) for details. A further simplification that I’ll adopt here, common in the literature (e.g., Schervish, 1995, p. 371), is to assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$. Recall that I’m also assuming here that the group $\mathscr{G}$ with respect to which the model is invariant is unique. There are cases, however, where the model is invariant with respect to more than one group, and I’ll elaborate on this case at the end of this subsection.
The set $\mathscr{G}x=\{gx:g\in\mathscr{G}\}\subseteq\mathbb{X}$ is called the orbit of $\mathscr{G}$ corresponding to $x$. The orbits partition $\mathbb{X}$ into equivalence classes, so every point $x\in\mathbb{X}$ falls on exactly one orbit. This partition can be used to construct a new coordinate system on $\mathbb{X}$ which will be useful for us in what follows. Identify $x\in\mathbb{X}$ with $(g_{x},u_{x})$, where $u_{x}\in\mathbb{U}$ denotes the label of orbit $\mathscr{G}x$ and $g_{x}\in\mathscr{G}$ denotes the position of $x$ on the orbit $\mathscr{G}x$.
Assumptions.
Suppose that the model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ is invariant with respect to a (unique) group $\mathscr{G}$, and that the corresponding density functions $\{p_{\theta}:\theta\in\mathbb{T}\}$ satisfy (19). For simplicity, assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$. In addition, assume that the following hold:
  • A1.
    The left Haar measure $\lambda$ and the corresponding right Haar measure $\rho$ on (the Borel $\sigma$-algebra of) $\mathscr{G}$ exist and are unique up to scalar multiples.
  • A2.
    There exists a bijection $t:\mathbb{X}\to\mathscr{G}\times\mathbb{U}$, with both $t$ and $t^{-1}$ measurable, that maps $x\in\mathbb{X}$ to its position–orbit coordinates $(g_{x},u_{x})\in\mathscr{G}\times\mathbb{U}$.
  • A3.
    The distribution of $t(X)=(G,U)\in\mathscr{G}\times\mathbb{U}$ induced by the distribution of $X\sim\mathsf{P}_{\theta}$ has a density with respect to $\lambda\times\mu$ for some measure $\mu$ on $\mathbb{U}$.
A couple quick remarks are in order. First, for A1, existence and uniqueness of the left and associated right Haar measures is a classical result, which holds under the mild condition that $\mathscr{G}$ be a locally compact topological group; see Halmos, (1950) and Nachbin, (1965). For A3, existence of a joint density with respect to a product measure ensures that there’s no difficulty in defining a conditional distribution for $G$, given $U=u$, which is needed below. Finally, $U=U_{X}$ is an ancillary statistic as a function of $X\sim\mathsf{P}_{\theta}$, meaning that its marginal distribution doesn’t depend on $\theta$.
The simplest example is a location model where $\mathscr{G}=\overline{\mathscr{G}}=(\mathbb{R},+)$. This is an abelian group, so the left and right Haar measures are the same and both equal to Lebesgue measure. The function $x\mapsto t(x)$ in A2 consists of two components: in its $g_{x}$” coordinate an equivariant function of $x$ that estimates the location and, in its $u_{x}$” component, an invariant function of $x$, such as residuals. For example, $g_{x}=\bar{x}$ the arithmetic mean of $x=(x_{1},\ldots,x_{n})$ and $u_{x}=\{x_{i}-\bar{x}:i=1,\ldots,n\}$. Note that the $u_{x}$ coordinate satisfies a constraint, so, after it’s represented in a suitable lower-dimensional space $\mathbb{U}$, $\mu$ can be taken as Lebesgue measure there. There are many other problems that fit this general form; see, e.g., Chapters 1–2 of Fraser, (1968), including the exercises.
Since I’m assuming $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$, generic $\theta\in\mathbb{T}$ can be identified as transformations in $\mathscr{G}$; the same goes for the uncertain $\Theta$. That means $\theta$ can be inverted to $\theta^{-1}$ and can be composed with other transformations in $\mathscr{G}$. A key result (e.g., Schervish, 1995, Cor. 6.64) is that the density of $X$ or, equivalently, of $t(X)=(G,U)$, under $\mathsf{P}_{\theta}$, is given by
$p_{\theta}(g,u)=f(\theta^{-1}\circ g,u),$ (20)
where the function $f:\mathscr{G}\times\mathbb{U}\to\mathbb{R}$ doesn’t directly depend on $\theta$. The particular form of $f$ isn’t important for what follows.
  • Lemma 6.65 in Schervish, (1995) gives a Bayesian posterior distribution $\mathsf{Q}_{x}^{\rho}$ for $\Theta$ that has the following density with respect to the right Haar prior measure $\rho$:
    $q_{x}^{\rho}(\theta)=\frac{d\mathsf{Q}_{x}^{\rho}}{d\rho}(\theta)\propto f(\theta^{-1}\circ g,u),\quad\theta\in\mathbb{T},\quad x\equiv(g,u).$ (21)
    This also agrees with the accepted fiducial distribution for $\Theta$, given $x$.
  • Proposition 2 in Martin, 2023a shows that the relative likelihood is given by
    $R(x,\theta)=d_{u}\,f(\theta^{-1}\circ g,u),\quad\theta\in\mathbb{T},\quad x\equiv(g,u),$
    where $d_{u}$ depends only on the $u$-component of $x$.
The above points are crucial to the following calculation. Indeed, since the $U$ component of $X$ is ancillary, the minimum complexity principle in Martin, 2022b says the IM construction should condition on the observation $U=u$. Write $\pi_{g|u}(\theta)$ for the possibilistic IM’s contour, where the subscript is meant to emphasize that the conditional distribution of $X\equiv(G,U)$, given the observed value of $u$, is used in the validification step (2). Then the IM contour is defined as
$\pi_{g|u}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\mid U=u\},\quad\theta\in\mathbb{T}.$ (22)
Then the following identities are consequences of the two results above that link the relative likelihood and the posterior density:
$\pi_{g|u}(\theta)$ $=\mathsf{P}_{\theta}\{f(\theta^{-1}\circ G,u)\leq f(\theta^{-1}\circ g,u)\mid U=u\}$
$=\mathsf{P}\{f(H,u)\leq f(\theta^{-1}\circ g,u)\mid U=u\}$
$=\mathsf{Q}_{x}^{\rho}\{f(H,u)\leq f(\theta^{-1}\circ g,u)\}$
$=\mathsf{Q}_{x}^{\rho}\{q_{x}^{\rho}(\Theta)\leq q_{x}^{\rho}(\theta)\},$
where the first line is by simplification of the definition of $\pi_{g|u}$; the second line is by the fact that $H:=\theta^{-1}\circ G$ is a pivot with respect to the conditional distribution of $G$, given $U=u$, under $\mathsf{P}_{\theta}$ (e.g., Schervish, 1995, Cor. 6.67); the third line is by the fact that $H:=\Theta^{-1}\circ g$ is also a pivot with respect to the Bayes posterior $\Theta\sim\mathsf{Q}_{x}^{\rho}$ and has the same distribution as $H$ in previous line; and the last line is by simplification via (21).
The last line in the above display can be recognized as the probability-to-possibility transform applied to the Bayesian/fiducial posterior distribution. This identifies the posterior distribution $\mathsf{Q}_{x}^{\rho}$ as an inner probabilistic approximation of the possibilistic IM $\mathsf{\Pi}_{x}=\mathsf{\Pi}_{g|u}$. This provides a new/different proof of the main result in Martin, 2023a , of which Theorem 3 above is a generalization.
As mentioned above, there are cases in which the group $\mathscr{G}$ with respect to which the model is invariant is not unique. A key example of this is in the case of Gaussian models, which are invariant with respect to linear (or affine) transformations $gx=Ax+b$, indexed by pairs $(A,b)$ consisting of a matrix $A$ and a vector $b$. Since the group indexed by all such matrices $A$ is “too large” (in the sense that the underlying group is not amenable), it’s natural to restrict the transformations to those indexed by certain sub-classes of matrices, corresponding to one or another lower-dimensional matrix factorization. It turns out that the right Haar prior depends on which class of matrix factorizations is used. For example, in the bivariate normal model indexed by means $\mu_{1}$ and $\mu_{2}$, standard deviations $\sigma_{1}$ and $\sigma_{2}$, and correlation $\rho$, there are two right Haar priors (Sun and Berger, 2007, Sec. 5), with respective density functions
$(\mu_{1},\mu_{2},\sigma_{1},\sigma_{2},\rho)\mapsto\frac{1}{\sigma_{2}^{2}(1-\rho^{2})}\quad\text{and}\quad(\mu_{1},\mu_{2},\sigma_{1},\sigma_{2},\rho)\mapsto\frac{1}{\sigma_{1}^{2}(1-\rho^{2})}.$
Of course, these two distinct priors lead to (slightly) different Bayes solutions with different statistical properties. Some might find this troubling: since the model is what determines the statistical inference problem, not the choice of group representation, it’s problematic if the Bayesian solution depends on a somewhat ad hoc choice between one group/right Haar prior or another.
In the construction of the IM solution described above, the dependence on the group structure is implicit: it’s in the identification of the $(G,U)$ components, where $G$ is an equivariant function of $X$ and $U$ is a (maximal) invariant function of $X$ which, of course, depend on $\mathscr{G}$. So, in general, different groups with respect to which the model is invariant would suggest a different “(G,U)(G,U)” representation. Since this representation determines how the conditioning is carried out in (22), this IM solution also depends on the particular choice of group $\mathscr{G}$. But I want to emphasize that conditioning on an ancillary statistic isn’t required for the IM construction; this is only recommended when an appropriate ancillary statistic is available. If the group isn’t unique, then there are multiple ancillary statistics available, yielding distinct conditional IMs. There are some natural ways one might remove the group-specific conditional IM’s undesirable dependence on the choice of group, but this is not really relevant to the present paper, so I’ll save this investigation for future work; see, also, the next paragraph. Alternatively, when there are multiple ancillary statistics, arguably none of them are fully “appropriate.” In that case, one might prefer to ignore conditioning and proceed with the basic likelihood-based IM construction as in Equation (2) of the main text. A corresponding inner probabilistic approximation retains the desirable probability matching and Bernstein–von Mises properties, but whether it corresponds to any Bayesian posterior distribution is an interesting open question.
To conclude this discussion about non-unique groups, let me consider the aforementioned Gaussian case specifically. There, the maximum likelihood estimator is a complete and sufficient statistic, so, by Basu’s theorem, whether one conditions on an ancillary statistic or not has no effect on the IM solution. Consequently, the same argument used above to prove Theorem 3 applied to the present case gives
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}=\mathsf{Q}_{x}^{\rho}\{q_{x}^{\rho}(\Theta)\leq q_{x}^{\rho}(\theta)\},\quad\theta\in\mathbb{T},$
for any right Haar prior $\rho$. Therefore, in the multivariate Gaussian case, it follows that all right Haar prior Bayes solutions are inner probabilistic approximations of the IM with contour as defined in Equation (2) of the main text. Compare this with the results in Eaton and Sudderth, (2012) which imply, for example, that highest posterior density regions corresponding to different right Haar priors in these multivariate normal models are exactly probability matching. Outside the Gaussian family, completeness typically doesn’t hold, so I currently don’t know if the above claim concerning “all right Haar prior Bayes solutions” applies more broadly.
C.2 Directional data example
Let $X=(X_{1},\ldots,X_{n})$ be the observable angles for $n$ independent spins of a roulette wheel. The model assumed here is the von Mises distribution, with density
$p_{\theta}(x)=\{2\pi I_{0}(\kappa)\}^{-1}\exp\{\kappa\cos(x-\theta)\},\quad x,\theta\in[0,2\pi).$
Here $\kappa>0$ is a known concentration parameter, and $I_{0}$ denotes the Bessel function of order 0. The unknown mean angle $\Theta$ is the parameter to be inferred from the data.
Digging into this model further, using the cosine formula
$\cos(x-\theta)=\cos x\cos\theta+\sin x\sin\theta,$
it’s easy to see that the minimal sufficient statistic for this model is the pair
$Y=(\bar{C},\bar{S})=\Bigl(\frac{1}{n}\sum_{i=1}^{n}\cos X_{i},\,\frac{1}{n}\sum_{i=1}^{n}\sin X_{i}\Bigr).$
Relative to the posited von Mises model, $Y$ can be treated as “the data.” Since $(C_{i},S_{i})=(\cos X_{i},\sin X_{i})$ denotes the Cartesian coordinates of the point on the unit circle corresponding to the angle $X_{i}$, it’s clear that $Y=(\bar{C},\bar{S})$ is generally inside the unit circle. The more “concentrated” the original data points are around a particular angle, the closer $Y$ will be to the boundary of the circle. Convert this average position to polar coordinates:
$G=\arctan(\bar{S}/\bar{C})\quad\text{and}\quad U=(\bar{C}^{2}+\bar{S}^{2})^{1/2}.$
Here $G$ is the angle the vector $(\bar{C},\bar{S})$ makes with the horizontal axis and $U$ is the Euclidean length of $(\bar{C},\bar{S})$. It’s clear that $(G,U)$ is a bijection of the minimal sufficient statistic $Y$, so write $t(X)=(G,U)$. I’m using the notation $G$ and $U$ to align with that introduced in the previous subsection. That is, there’s a group of transformations $\mathscr{G}$ that corresponds to rotations of points in the plane; technically, $\mathscr{G}$ is the special orthogonal group $\text{SO}(2)$ consisting of orthogonal matrices with unit determinant. Then $G$ can be interpreted as both an angle or as a rotation by that angle. Similarly, the interior of the unit circle can be partitioned into distinct concentric circles, which correspond to orbits, and $U$ determines which of these orbits $Y$ sits on.
The details here align with the general setup above, so I can proceed to carry out the fiducial or default-prior Bayes and IM analyses. It’s a standard result (e.g., Mardia and Jupp, 2000, Eq. 4.5.5) that the joint density for $t(X)=(G,U)$ factors as
$p_{\theta}(g,u)$ $=p_{\theta}(g\mid u)\,p(u)$
$=\{2\pi I_{0}(\kappa u)\}^{-1}\exp\{\kappa u\cos(g-\theta)\}\,p(u),$
where $p(u)$ denotes the marginal density for $U$, which doesn’t depend on $\theta$ and isn’t relevant here. The expression for the conditional density of $G$, given $U=u$, is easily seen to be that of a von Mises distribution but with concentration parameter $\kappa u$. The influence of conditioning on $U$ makes intuitive sense because, as indicated above, the statistic $U$ acts like a measure of how “concentrated” the observed angles are around a common angle. It’s the conditional density above that drives both the fiducial and IM analyses. Indeed, the no-prior Bayes (and fiducial) density is
$q_{x}(\theta)\propto\exp\{\kappa u\cos(g-\theta)\},\quad\theta\in(0,2\pi],$
and, after simplification, the IM’s possibility contour is
$\pi_{g|u}(\theta)=\mathsf{P}\{\cos H\leq\cos(g-\theta)\mid U=u\},\quad\theta\in[0,2\pi),$
where the probability is with respect to the conditional distribution of $H:=G-\Theta$, given $U=u$, which is simply a von Mises distribution with mean angle 0 and concentration parameter $\kappa u$. This contour function is easy to evaluate numerically and to plot as in the main text. By setting a sufficiently fine grid on the interval $(0,2\pi)$ and then evaluating the contour on that grid, it’s easy to obtain the level set $C_{\alpha}(x)$ for any $\alpha$. From here it’s straightforward to sample from the inner probabilistic approximation, etc.
Appendix D Details from Section 4.2
D.1 Setup and regularity conditions
Certain regularity conditions are required in order to establish asymptotic concentration properties of estimators, posterior distributions, etc., and the same is true for IMs and the inner probabilistic approximations under consideration here. Roughly, these conditions ensure that the log-likelihood function is smooth enough that it can be well-approximated by a quadratic function. One common set of regularity conditions are the classical Cramér conditions (Cramér, 1946), versions of which can be found in the standard texts, including Lehmann and Casella, (1998, Theorem 3.10) and Schervish, (1995, Theorem 7.63). Here I’ll adopt the more modern set of sufficient conditions originating in Le Cam, (1956, 1960, 1970) and Hájek, (1972); see, also, Bickel et al., (1998) and van der Vaart, (1998).
The posited model specifies a class $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ of probability distributions, supported on $\mathbb{X}$, with $\mathsf{P}_{\theta}$ having a density function $p_{\theta}(x)$ relative to a $\sigma$-finite measure $\nu$ on $\mathbb{X}$. Then the data $X^{n}=(X_{1},\ldots,X_{n})$ is assumed to be iid of size $n$ with common distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the uncertain “true value” of the parameter. Following Bickel et al., (1998, Ch. 2), define the (natural) logarithm and square-root density functions:
$\ell_{\theta}(x)=\log p_{\theta}(x)\quad\text{and}\quad s_{\theta}(x)=p_{\theta}(x)^{1/2}.$
The “dot” notation, e.g., $\dot{g}_{\theta}(x)$, represents a function that behaves like the derivative of $g_{\theta}(x)$ with respect to $\theta$ for pointwise in $x$. If the usual partial derivative of $g_{\theta}(x)$ with respect to $\theta$ exists, then $\dot{g}_{\theta}(x)$ is that derivative; but suitable functions $\dot{g}_{\theta}(x)$ may exist even when the ordinary derivative fails to exist. In particular, below I’ll assume existence of a suitable “derivative” $\theta\mapsto\dot{s}_{\theta}(x)$ of the square-root density. Finally, let $\mathcal{L}_{2}(\nu)$ denote the set of measurable functions on $\mathbb{X}$ that are square $\nu$-integrable.
Assumptions.
The parameter space $\mathbb{T}$ is open and there exists a vector $\dot{s}_{\theta}(x)=\{\dot{s}_{\theta,d}(x):d=1,\ldots,D\}$, with $\dot{s}_{\theta,d}$ an element of $\mathcal{L}_{2}(\nu)$, such that the following conditions hold:
  • A1.
    the maps $\theta\mapsto\dot{s}_{\theta,d}$ from $\mathbb{T}$ to $\mathcal{L}_{2}(\nu)$ are continuous for each $d=1,\ldots,D$;
  • A2.
    at each $\theta\in\mathbb{T}$,
    $\int\bigl|s_{\theta+u}(x)-s_{\theta}(x)-u^{\top}\dot{s}_{\theta}(x)\bigr|^{2}\,\nu(dx)=o(\|u\|^{2}),\quad u\to 0\in\mathbb{R}^{D};$ (23)
  • A3.
    and the $D\times D$ matrix $\int\dot{s}_{\theta}(x)\,\dot{s}_{\theta}(x)^{\top}\,\nu(dx)$ is non-singular for each $\theta\in\mathbb{T}$.
The condition in (23) is often described as $\theta\mapsto s_{\theta}$ being differentiable in quadratic mean. Note that this condition does not require the square-root density to actually be differentiable at $\theta$, only that it be locally “linearizable” in an average sense, like a differentiable-at-$\theta$ function would be. The classical Cramér conditions assume more than two continuous derivatives, so (23), which does not even require existence of a first derivative, is significantly weaker; sufficient conditions for (23) are given in van der Vaart, (1998, Lemma 7.6). Then the score function $\dot{\ell}_{\theta}(x)$ is defined in terms of $\dot{s}_{\theta}(x)$ as
$\dot{\ell}_{\theta}(x)=\frac{2\dot{s}_{\theta}(x)}{s_{\theta}(x)}\,1\{s_{\theta}(x)>0\},$
and it can be shown that $\int\dot{\ell}_{\theta}(x)\,\mathsf{P}_{\theta}(dx)=0$ for each $\theta$. Moreover, Condition A3 above implies non-singularity of the Fisher information matrix $I_{\theta}=\int\dot{\ell}_{\theta}(x)\ell_{\theta}(x)^{\top}\,\mathsf{P}_{\theta}(dx)$ for each $\theta\in\mathbb{T}$. Bickel et al., (1998, Prop. 2.1.1) provide sufficient conditions for A1–A3.
The above conditions are rather mild, and hold in a broad range of problems, including exponential families. One further condition is required for the possibilistic Bernstein–von Mises theorem in Section 4.2, namely, that the maximum likelihood estimator is consistent, i.e., $\hat{\theta}_{X^{n}}\to\Theta$ in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$. This, is not automatic, but holds quite broadly when the parameter space dimension is fixed, as I’m assuming here.
D.2 Proof of Theorem 5
The intuition behind the Bernstein–von Mises result itself, and the proof that follows, is that both the inner probabilistic approximation and the Gaussian that it’s merging with are fully determined by corresponding possibility contours. So, any difference in these distributions will imply a difference in the possibility contours, but the contours themselves are merging by Theorem 4 and, therefore, any difference in the distributions must be vanishing asymptotically. The following details make this argument rigorous.
I’ll start by justifying the above claim that any difference between $\mathsf{Q}_{X^{n}}^{\star}$ and the Gaussian approximation $\mathsf{G}_{X^{n}}$ is due to differences in the corresponding possibility contours. The first observation is that $\mathsf{Q}_{X^{n}}^{\star}$ has the characterization as a uniform mixture of uniform kernels by Theorem 1. The second observation is that, similarly, $\mathsf{G}_{X^{n}}$ is the inner probabilistic approximation of the Gaussian possibility measure with contour $\gamma_{X^{n}}$ and, therefore, the characterization result in Theorem 1 applies to $\mathsf{G}_{X^{n}}$ too. That is, if $\mathsf{K}_{X^{n},\pi}^{\alpha}$ and $\mathsf{K}_{X^{n},\gamma}^{\alpha}$ are the uniform kernels supported on the level sets $\{\theta:\pi_{X^{n}}(\theta)=\alpha\}$ and $\{\theta:\gamma_{X^{n}}(\theta)=\alpha\}$, then
$\mathsf{Q}_{X^{n}}^{\star}(\cdot)=\int_{0}^{1}\mathsf{K}_{X^{n},\pi}^{\alpha}(\cdot)\,d\alpha\quad\text{and}\quad\mathsf{G}_{X^{n}}(\cdot)=\int_{0}^{1}\mathsf{K}_{X^{n},\gamma}^{\alpha}(\cdot)\,d\alpha.$
Then, for any Borel measurable subset $H$ of $\mathbb{T}$,
$\bigl|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr|\leq\int_{0}^{1}\bigl|\mathsf{K}_{X^{n},\pi}^{\alpha}(H)-\mathsf{K}_{X^{n},\gamma}^{\alpha}(H)\bigr|\,d\alpha,$
so $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies that there exists an $H$ such that $|\mathsf{K}_{X^{n},\pi}^{\alpha}(H)-\mathsf{K}_{X^{n},\gamma}^{\alpha}(H)|$ exceeds $\varepsilon$ on a set of $\alpha$’s having positive Lebesgue measure. This implies that there exists an open subset of $H$ and a non-empty interval $(\alpha_{1},\alpha_{2})$ such that
$\pi_{X^{n}}(\theta)\in(\alpha_{1},\alpha_{2})\quad\text{and}\quad\gamma_{X^{n}}(\theta)>\alpha_{2}\quad\text{for all $\theta$ in that subset;}$
of course, the last inequality could be replaced by $\gamma_{X^{n}}(\theta)<\alpha_{1}$ and the positions of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ in the above display could be reversed, but these changes don’t affect the argument that follows. What matters is that there exists a point $\theta$—depending implicitly on data $X^{n}$, on $H$, and on $\varepsilon$—such that
$|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\alpha_{2}-\alpha_{1}>0.$
That Theorem 4 implies the merging of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ uniformly means that the event in the above display has vanishing $\mathsf{P}_{\Theta}$-probability as $n\to\infty$.
What’s not fully satisfactory about the above conclusion is that the positive bound $\alpha_{2}-\alpha_{1}$” is only implicit. An important point that the above argument reveals is that any differences in how $\mathsf{Q}_{X^{n}}^{\star}$ and $\mathsf{G}_{X^{n}}$ assign probability will manifest through the contour level sets. Since these are also the sets that determine the distributions themselves, it’s without loss of generality to consider hypotheses $H$ that take the form of lower level sets of either of the defining contours. For concreteness, consider
$H=\{\theta:\pi_{X^{n}}(\theta)\leq\tau\},\quad\tau\in(0,1).$
By definition of the inner probabilistic approximation, $\mathsf{Q}_{X^{n}}^{\star}(H)=\sup_{\theta\in H}\pi_{X^{n}}(\theta)=\tau$. Moreover, since $H$ is bounded and excludes the mode of $\gamma_{X^{n}}$, it’s easy to check that $\mathsf{G}_{X^{n}}(H)$ satisfies
$\inf_{\theta\in\partial H}\gamma_{X^{n}}(\theta)\leq\mathsf{G}_{X^{n}}(H)\leq\sup_{\theta\in\partial H}\gamma_{X^{n}}(\theta).$
Consequently,
$\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)$ $\leq\tau-\inf_{\theta\in\partial H}\gamma_{X^{n}}(\theta)$
$=\sup_{\theta\in\partial H}\{\tau-\gamma_{X^{n}}(\theta)\}$
$=\sup_{\theta\in\partial H}\{\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\}$
$\leq\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|,$
where the second-to-last line follows because $\pi_{X^{n}}(\theta)\equiv\tau$ on $\partial H$. Similarly,
$\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)$ $\geq\tau-\sup_{\theta\in\partial H}\gamma_{X^{n}}(\theta)$
$=\inf_{\theta\in\partial H}\{\tau-\gamma_{X^{n}}(\theta)\}$
$=\inf_{\theta\in\partial H}\{\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\}$
$\geq-\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|.$
Putting these two inequalities together gives
$|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|\leq\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|.$
Therefore,
$|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon\implies\sup_{\theta\in\partial H}|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon,$
and, by Theorem 4, the latter event has vanishing $\mathsf{P}_{\Theta}$-probability as $n\to\infty$.
Putting everything together, $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies existence of a set $H$ for which $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$. This implies a non-negligible difference in the contours so, without loss of generality, the problematic HH—whose existence is guaranteed—can be taken as a lower level set corresponding to one of the contours. When $H$ is such a lower level set, $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$ implies that $|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon$ for all $\theta$ on the boundary of $H$. Since the latter event has vanishing $\mathsf{P}_{\Theta}$-probability by Theorem 4, so must the event $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$, thus proving Theorem 5.
Appendix E Details from Section 4.3
A claim was made in Section 4.3 that confidence distributions (CDs)—or no-prior Bayes solutions in general—often transform into over-confidence distributions (oCDs) when (non-linear) marginalization is carried out. Here I show just a couple simple illustrations to justify this claim. While my details are different, the insights were inspired by the analysis in Fraser, 2011a .
The first concerns a single sample $X$ from a simple Gaussian model with known variance, say $\sigma^{2}=1$, but unknown mean $\Theta$. Of course, there is an exact confidence interval $C_{\alpha}(x)=x\pm{\tt qnorm}(1-\frac{\alpha}{2})$ and the standard CD, $\mathsf{Q}_{x}^{\star}={\sf N}(x,1)$, assigns exactly probability $1-\alpha$ to $C_{\alpha}(x)$. But now suppose the goal is marginal inference on $\Phi=m(\Theta)$ where $m$ is bounded and, hence, non-linear, e.g.,
$m(\theta)=\max\{-1,\min(1,\theta)\},\quad\theta\in\mathbb{R}.$ (24)
The claim is that the marginal CD of $m(\Theta)$ will assign more than probability $1-\alpha$ to the image $m\{C_{\alpha}(x)\}$ of the original confidence interval under $m$. Figure 10 shows the (marginal) posterior probability assigned to $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for several different realizations $x$. That the curve generally falls above the diagonal line is a demonstration of the aforementioned over-confidence.
(a) $x=0$
(b) $x=0.5$
(c) $x=1$
(d) $x=2$
Figure 10: Plots of the marginal posterior probability assigned $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for several different $x$, where $m$ is the bounded function (24).
As a second example, suppose that $X=(X_{1},X_{2})$ is a single realization from a bivariate normal with mean vector $\Theta=(\Theta_{1},\Theta_{2})$ and identity covariance matrix. Again, it’s straightforward to construct a confidence disc $C_{\alpha}(x)$ and a confidence distribution $\mathsf{Q}_{x}^{\star}={\sf N}_{2}(x,I)$. Consider marginalization via the familiar non-linear function $m(\theta)=\|\theta\|^{2}$. I say “familiar” because this is an infamous example going to back at least to Stein, (1959) and more recently in Fraser, 2011a , Balch et al., (2019), and elsewhere. Figure 11 shows the same plots of marginal posterior probability assigned to $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for different realizations $x$. The over-confidence appears for any $x$ that’s away from the origin; the discrepancy between the curve and the diagonal line is increasing in $\|x\|^{2}$ to a point, and then remains roughly constant as $\|x\|^{2}\to\infty$.
(a) $x=(0.1,0.3)$
(b) $x=(0.1,3)$
Figure 11: Plots of the marginal posterior probability assigned $m\{C_{\alpha}(x)\}$ as a function of the level $\alpha$ for several different $x$, where $m(\theta)=\|\theta\|^{2}$ is the squared norm.
Appendix F Computational details from Section 5
Here I review some of the relevant details presented in Cella and Martin, (2025) on the choice of $\sigma=\sigma(x,\alpha)$ in Section 5. In this review I’ll focus on use of a Gaussian variational family, which is motivated by the asymptotic results in Section 4.2, but similar developments can surely be made with other distributional families. The key features of the Gaussian are that its credible sets can be described in closed-form (ellipsoids) and that it’s straightforward to sample on the boundary of these credible sets.
Let $\mathsf{G}_{x}^{\sigma}$ denote the Gaussian distribution ${\sf N}_{D}(\hat{\theta}_{x},(J_{x}^{\sigma})^{-1})$ with mean equal to the maximum likelihood estimator and covariance matrix the inverse of the $\sigma$-adjusted observed Fisher information matrix in (15). Recall that $\sigma$ scales the eigenvalues of $J_{x}^{\sigma}$, in particular, the eigenvalues are decreasing elementwise in $\sigma$. The corresponding Gaussian credible/confidence sets are
$C_{\alpha}^{\sigma}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}J_{x}^{\sigma}(\theta-\hat{\theta}_{x})\leq\chi_{D,1-\alpha}^{2}\},\quad\alpha\in[0,1].$
The intuition is that $C_{\alpha}^{\sigma}(x)$ can be appropriately inflated or deflated by varying $\sigma$, and this is relevant because the Gaussian approximation $\mathsf{G}_{x}^{\sigma}$ assigns probability at least $1-\alpha$ to the original IM’s $\alpha$-cut $C_{\alpha}(x)$ if $C_{\alpha}^{\sigma}(x)\supseteq C_{\alpha}(x)$ or, equivalently, if
$\sup_{\theta\not\in C_{\alpha}^{\sigma}(x)}\pi_{x}(\theta)\leq\alpha.$
Since the contour $\pi_{x}$ is itself approximately Gaussian (Martin and Williams, 2025) and the maximum likelihood estimator $\hat{\theta}_{x}$, also the mode of $\pi_{x}$, is in $C_{\alpha}^{\sigma}(x)$, the above supremum should be attained on the boundary $\partial C_{\alpha}^{\sigma}(x)$. Moreover, since equality in the above display implies a near-perfect match between the IM’s and the posited Gaussian $\alpha$-cuts, a reasonable goal is to find a root to the function
$g_{\alpha}(\sigma):=\max_{\theta\in\partial C_{\alpha}^{\sigma}(x)}\pi_{x}(\theta)-\alpha.$
Design of an iterative algorithm to find this root requires care, primarily because evaluating $\pi_{x}$ is expensive; so the goal is to evaluate $g_{\alpha}(\sigma)$ with as few $\pi_{x}$ evaluations as possible. Cella and Martin, (2025) propose to represent the boundary of $C_{\alpha}^{\sigma}(x)$ by $2D$-many vectors
$\vartheta_{s}^{\sigma,\pm}:=\hat{\theta}_{x}\pm\bigl\{\chi_{D,1-\alpha}^{2}\,\sigma_{s}\,/\,\lambda_{s}\}^{1/2}\,e_{s},\quad s=1,\ldots,d,$ (25)
where $(\lambda_{s},e_{s})$ is the eigenvalue–eigenvector pair corresponding to the $s^{\text{th}}$ largest eigenvalue in the spectral decomposition of $J_{x}$. Define the vector function $\hat{g}_{\alpha}$ with components
$\hat{g}_{\alpha,s}(\sigma)=\max\{\pi_{x}(\vartheta_{s}^{\sigma,+}),\pi_{x}(\vartheta_{s}^{\sigma,-})\}-\alpha,\quad s=1,\ldots,D.$ (26)
At least intuitively, negative and positive $\hat{g}_{\alpha,s}(\sigma)$ indicate that the $\alpha$-cut $C_{\alpha}^{\sigma}(x)$ is too large and too small, respectively in the $e_{s}$-direction. From here, Cella and Martin, (2025) suggest applying a stochastic approximation algorithm—see Robbins and Monro, (1951) and Kushner and Yin, (2003)—to construct a sequence $(\sigma^{(t)}:t\geq 1)$ of $D$-vectors that converges to a root of $\hat{g}_{\alpha}$ and, hence, an approximate root of $g_{\alpha}$. For an initial guess $\sigma^{(0)}$, the specific sequence is defined as
$\sigma_{s}^{(t+1)}=\sigma_{s}^{(t)}+w_{t+1}\,\hat{g}_{\alpha,s}(\sigma^{(t)}),\quad s=1,\ldots,D,\quad t\geq 0,$
where $(w_{t})$ is a deterministic sequence that satisfies
$\sum_{t=1}^{\infty}w_{t}=\infty\quad\text{and}\quad\sum_{t=1}^{\infty}w_{t}^{2}<\infty.$
These steps are iterated until (practical) convergence is achieved, and the limit is what I called $\sigma(x,\alpha)$ in Section 5. The basic steps are outlined in Algorithm 1, but I refer to Cella and Martin, (2025) for further details and discussion.
requires: data $x$, eigen-pairs $(\lambda_{s},e_{s})$, and ability to evaluate $\pi_{x}$;
initialize: $\alpha$-level, guess $\sigma^{(0)}$, step size sequence $(w_{t})$, and threshold $\varepsilon>0$;
set: stop = FALSE, $t=0$;
while !stop do
   construct the representative points $\{\vartheta_{s}^{\sigma^{(t)},\pm}:s=1,\ldots,D\}$ as in (25);
   evaluate $\hat{g}_{\alpha,s}(\sigma_{s}^{(t)})$ for $s=1,\ldots,D$ as in (26);
   update $\sigma_{s}^{(t+1)}=\sigma_{s}^{(t)}\pm w_{t+1}\,\hat{g}_{\alpha,s}(\sigma_{s}^{(t)})$ for $s=1,\ldots,D$;
 if $\max_{s}|\sigma_{s}^{(t+1)}-\sigma_{s}^{(t)}|<\varepsilon$ then
    $\sigma(z,\alpha)=\sigma^{(t+1)}$;
    stop = TRUE;
 else
    $t\leftarrow t+1$;
   end if
end while
return $\sigma(x,\alpha)$;
Algorithm 1 Determining $\sigma(x,\alpha)$—from Cella and Martin, (2025)
Appendix G Additional examples
G.1 Bivariate normal correlation
A challenging one-parameter inference problem is that where $X=(X_{1},\ldots,X_{n})$ are iid, with $X_{i}=(X_{1i},X_{2i})$ a bivariate normal random vector with zero means, unit standard deviations, and correlation $\Theta\in\mathbb{T}=(-1,1)$ to be inferred. Ironically, the known mean and variance case is more difficult—it’s a curved exponential family so the minimal sufficient statistic is not complete and there are various ancillary statistics available to condition on (e.g., Basu, 1964). How this curvature affects asymptotic inference is detailed in Reid, (2003). A likelihood-based possibilistic IM, which is exactly valid for all sample sizes and asymptotically efficient, is easy to construct in theory (Martin, 2024a , Example 2), but naive computation is relatively expensive because there’s no pivot structure and no closed-form expression for the maximum likelihood estimator. This motivates the faster, approximate solution suggested in Martin, 2025a and Section 5 above, based on the variational approximation strategy developed in Cella and Martin, (2025).
A relevant feature of the exact IM solution is that, when $\Theta$ is relatively close the extremes, i.e., $\Theta\approx\pm 1$, the IM’s possibility contour tends to be very asymmetric. Consequently, approximation suggested in Martin, 2025a and Section 5 above, which are based on stitching together symmetric, Gaussian credible intervals, tend to be a bit conservative. Other asymmetric distributions, e.g., skewed normal (e.g., Azzalini, 1985), could potentially replace the Gaussian recommendation made in Martin, 2025a , but that’s beyond the scope of the present paper. I’m mentioning this point here because the same conservative behavior will be observed below with the inner probabilistic approximation; this isn’t a shortcoming of the inner probabilistic approximation itself, but of the choice to bound the IM’s asymmetric level sets using symmetric Gaussian credible limits.
For illustration, let $X=(X_{1},\ldots,X_{n})$ denote (a centered and scaled version of) the law school admissions data analyzed in Efron, (1982), where $n=15$ and the maximum likelihood estimator of $\Theta$ is $\hat{\theta}_{x}=0.789$. Figure 12(a) shows a plot of the Bayesian posterior density, based on Jeffreys prior, along with samples from the inner probabilistic approximation. (The latter is based on first reparametrizing in terms of Fisher’s z-transformation, i.e., $\Psi=\tanh^{-1}\Theta$, obtaining the inner probabilistic approximation of $\Psi$, and then transforming back to $\Theta=\tanh\Psi$.) Both distributions are concentrated around the maximum likelihood estimator and have similar shapes, but the Bayes posterior is a little tighter than inner probabilistic approximation.
Is the Bayes posterior distribution too concentrated in the above illustration? If so, then is this aggressive behavior systematic or just a fluke in this one example? To answer these questions, I carried out a brief simulation study, involving 3000 data sets of size $n=10$ drawn from a standard bivariate normal distribution with $\Theta=0.7$. For both methods, and each data set $X$, I evaluate
$U_{X}=1-\bigl|2Q_{X}(\Theta)-1\bigr|,$
where $Q_{X}(\Theta)$ is the cumulative distribution function of the derived posterior distribution, based on the Bayes posterior or the inner probabilistic approximation, evaluated at the true $\Theta=0.7$. If the random variable $U_{X}$ is stochastically no smaller than ${\sf Unif}(0,1)$, then the corresponding equal-tailed posterior credible intervals have frequentist coverage probability at or above the nominal level; if $U_{X}$ is stochastically smaller than ${\sf Unif}(0,1)$, the credible intervals undercover. Figure 12(b) plots the cumulative distribution function of $U_{X}$ for the two methods and we find that the Jeffreys-prior Bayes solution and the inner probabilistic approximation solution yield $U_{X}$ that are stochastically smaller and larger than ${\sf Unif}(0,1)$, respectively. This means that the former’s credible intervals generally undercover while the latter’s generally overcover; therefore, the Bayes posterior density displayed in Figure 12(a) might indeed be “too concentrated.”
(a) Posterior results in Efron’s example
(b) Cumulative distribution function of $U_{X}$
Figure 12: Comparison of Jeffreys-prior Bayes and the possibilistic IM’s inner probabilistic approximation. In Panel (a), the solid line is the Bayes posterior density and the bars are based on samples from the inner probabilistic approximation. In Panel (b), the black and gray lines correspond to the Bayes and inner probabilistic approximation solutions, respectively. Definition and explanation of $U_{X}$ is given in the text.
G.2 Gamma shape and scale
Inference on the parameters of a gamma model is a classical problem. When the shape parameter is unknown, this model lacks the group transformation structure that makes specification of default priors having suitable reliability properties non-trivial. The go-to no-prior Bayes solution invokes Jeffreys prior on the shape–scale parameter pair. Alternatively, a possibilistic IM solution is readily available and doesn’t require any prior specification and has reliability guarantees. This IM solution is somewhat expensive to compute, however, at least if a naive approach is taken. Martin, 2025a showed how this naive approach can be drastically improved and it’s on this more sophisticated/efficient approach that my solution is based. Indeed, here I’m proposing the inner probabilistic approximation of that possibilistic IM, and the sampling from that “posterior distribution” is based on the strategy put forward in Martin, 2025a .
As a first illustration, reconsider the data analyzed in Fraser et al., (1997) on the survival time (in weeks) for $n=20$ rats exposed to a certain amount of radiation. Assuming a gamma model with unknown shape and scale parameters, the posterior distributions for the two no-prior Bayes solutions—Jeffreys prior Bayes and the IM’s inner probabilistic approximation—are shown in Figure 13, both in the original and log-transformed parametrization. The two posteriors look generally similar and, after log-transformation, they both have Gaussian-like elliptical shapes. The only notable difference is that the Bayes solution is a bit more tightly concentrated than the IM-based solution.
(a) shape–scale parametrization
(b) log(shape)–log(scale) parametrization
Figure 13: Black points are samples from the Jeffreys-prior Bayes posterior distribution in the gamma model; gray points are samples from the IM’s inner probabilistic approximation.
Like in the bivariate normal correlation example above, it’s natural to ask if the Bayes solution is actually too tightly concentrated. To assess this, I’ll do a brief simulation study. Here I draw 2500 samples of size $n=20$, with true shape and scale parameters set to be $\Theta_{1}=7$ and $\Theta_{2}=3$, respectively. Similar to the bivariate normal example in the previous section, for both no-prior Bayes solutions, define the quadratic form of the form
$\kappa_{x}(\theta)=(\log\theta-\hat{\mu}_{x})^{\top}\hat{\sigma}_{x}^{-1}(\log\theta-\hat{\mu}_{x}),$
where $\hat{\mu}_{x}$ and $\hat{\sigma}_{x}$ are estimates of the posterior mean and posterior covariance matrix of the log-transformed shape and scale parameters. Then I let
$Q_{x}(\theta)=\mathsf{Q}_{x}\{\kappa_{x}(\widetilde{\Theta})\leq\kappa_{x}(\theta)\}$
denote the distribution function of the quadratic form relative to the posterior distribution and, finally, I define
$U_{X}=1-\bigl|2Q_{X}(\Theta)-1\bigr|,$
where $\Theta=(7,3)$ is the true shape and scale parameters. Figure 14 plots the distribution functions of $U_{X}$ for each of the two methods. Being above the diagonal line indicates that the elliptical credible sets derived from the method’s posterior samples have frequentist coverage probability below the nominal level; similarly, being on or below the line indicates that the credible sets have frequentist coverage probability at or above the nominal level. In this case, the Jeffreys-prior Bayes solution shows severe undercoverage while the IM’s inner probabilistic approximation shows signs of very slight undercoverage for 90 and 95% credible regions. This suggests that the Jeffreys-prior Bayes solutions displayed in Figure 13 might indeed be too tightly concentrated.
Figure 14: Comparison of Jeffreys-prior Bayes and the possibilistic IM’s inner probabilistic approximation, in particular, distribution functions of $U_{X}$ corresponding to the Bayes (black) and IM inner probabilistic approximation (gray).